Specifications and Notation for MULTEXT-East Lexicon Encoding

Contributors:

Bulgarian: R.Pavlov, L.Dimitrova, L.Sinapova, K.Simov Croatian: M.Tadić Czech: V.Petkevič Estonian: H.J.Kaalep

English: N.Ide, G.Priest-Dorman, T.Erjavec Hungarian: L.Tihanyi, T.Váradi, C.Oravecz

Romanian: D.Tufiş, A.M.Barbu

Slovene: T.Erjavec, P.Holozan, V.Gorjanc, M.Stabej

Editor Tomaž Erjavec

Edition

Multext-East / Concede

March 21th, 2001

Supported by EU projects Multext-East, Concede and TELRI

Foreword

The purpose of this document is to (i) provide harmonised lexical specifications for eight languages — Bulgarian, Croatian, Czech, Estonian, English, Hungarian, Romanian, Slovene — and (ii) formulate the relevant notation that is used in the lexicons and annotated corpora contributed by the language groups.

This edition of the lexical specifications is the revised and expanded version of the MULTEXT-East report D11F. This report was, in turn, based on two proposals of lexicon specifications presented in the Multext D1-6-1B Deliverable of the Multext Project (Bel, Calzolari and Monachini eds. 1995) and in the Eagles document of the Lexicon sub-group on Morphosyntactic annotation (Monachini and Calzolari 1995).

The current ’Concede’ version offers the addition of a new language, Croatian, and additions (for Romanian and Slovene) to the common tables in terms of attributes and values. The tables for Slovene have also been localised. The L^{A}TEX document has been converted to Latin-2 encoding, so that the language specific characters are correctly displayed in print and in the HTML version of this document.

Pisa, October 1995

Ljubljana, December 1997

Ljubljana, March 2001

Contents

1 Background 1 1.1 Description of the proposal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Lexical lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Organization of the language-specific chapters . . . . . . . . . . . . . . . . . . . . 3

I Common Morphosyntactic Specifications 4

2 Table of categories 5

3 Tables of attribute-values 6 3.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.10 Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.11 Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.13 Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.14 Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.15 List of values with codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.16 List of attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

II Language Specific Applications 32

4 Application to English 33 4.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.10 Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.11 Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.13 Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.14 Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

5 Application to Romanian 48 5.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 5.10 Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.11 Interjections (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 5.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 5.13 Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 5.14 Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6 Application to Slovene 72 6.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 6.2 Table of Categories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 6.3 Noun (N) // Samostalnik (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 6.4 Verb (V) // Glagol (G) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 6.5 Adjective (A) // Pridevnik (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 6.6 Adverb (R) // Prislov (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 6.7 Pronoun (P) // Zaimek (Z) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 6.8 Adposition (S) // Predlog (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 6.9 Conjunction (C) // Veznik (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 6.10 Particle (Q) // Členek (L) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.11 Interjection (I) // Medmet (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.12 Numeral (M) // Števnik (K) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.13 Abbreviation (Y) // Okrajšava (O) . . . . . . . . . . . . . . . . . . . . . . . . . . 105 6.14 Residual (X) // Neuvrščeno (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

7 Application to Czech 106 7.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 7.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 7.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 7.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 7.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 7.10 Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.11 Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 7.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 7.13 Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 7.14 Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

8 Application to Bulgarian 125 8.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 8.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 8.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 8.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 8.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 8.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 8.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 8.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 8.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 8.10 Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 8.11 Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 8.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 8.13 Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 8.14 Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

9 Application to Estonian 138 9.1 Nouns (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 9.2 Verbs (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 9.3 Adjectives (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 9.4 Pronouns (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 9.5 Determiners (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 9.6 Atricles (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 9.7 Adverbs (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 9.8 Adpositions (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 9.9 Conjunctions (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 9.10 Numerals (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 9.11 Interjections (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 9.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 9.13 Abreviations (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 9.14 Particles (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

10 Application to Hungarian 182 10.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 10.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 10.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 10.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 10.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 10.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 10.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 10.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 10.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 10.10Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 10.11Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 10.12Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 10.13Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 10.14Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 10.15Derivation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 10.16Compounding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

11 Application to Croatian 202 11.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 11.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 11.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 11.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 11.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 11.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 11.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 11.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 11.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 11.10Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 11.11Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 11.12Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 11.13Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 11.14Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

Chapter 1

Background

1.1 Description of the proposal

The proposal has been prepared in the Multext table format, which displays the specifications (as sets of attribute-values, see below for further details about the notation), with their respective codes used to mark them in the lexicons. Two types of features are distingished:

(i) the minimal core features These are shared by most of the languages and appear at the top of the list of features — they are deliniated by a row of asterisks (*). We tried to keep this set in common to all the Multext and Multext-East languages. In such a way, the comparability across the information encoded in the lexical lists of Central & Eastern and of the six original Multext languages is to a certain extent ensured. A side-effect is that certain values are present in the tables, although none of the MULTEXT-East languages makes used of them/

(ii) the Multext-East language-specific features These are shared by the Multext-East languages, but differ from the Multext ones. The formulation of this set has been, as already mentioned, highly delicate, due also to the fact that many language-specific values were presented in the applications and sometimes the same (or ’similar enough’) morphosyntactic phenomena were referred to with two different attribute or value names. The phase of recognition and harmonisation of the semantics of some attributes, values and naming conventions has, hence, required much effort.

This representation, with the concrete applications which display and exemplify the attributes and values and provide their internal constraints and relationships, makes the proposal selfexplanatory. Other groups can easily test the specifications on their language, simply by following the method of the applications. The possibility of incorporating idiosyncratic classes and distinctions after the common core features makes the proposal relatively adaptable and flexible, without compromising compatibility.

1.2 Lexical lists

The specifications presented here have been used in the encoding of word-form lexica of the languages. These lexica contain one entry per line, where an entry has the following form:

word-form ⟨TAB⟩ lemma ⟨TAB⟩ morphosyntactic description

This reports describes the morphosyntactic descriptions used in the lexica and also in the annotated multilingual corpus.

1.3 Notation

In Multext the notation has been chosen following current practices for NLP, where information is represented in attribute-value formalisms and following the idea that it should also be selfinformative for human understanding. At the same time, a relativelly compact encoding was maintained. The notation format has the following main characteristics:

• attributes are marked by positions;

• values are represented by a single character;

• a special marker reflects the non applicability of a given attribute.

These characteristics make the proposed lexical notation similar to attribute-value pairs used in unification based formalisms (see the Multext D1-6-1B Deliverable for further details).

The linear strings of characters representing the morphosyntactic descriptions are constructed following the philosophy of the Intermediate Format proposed in the Eagles Corpus proposal (Leech and Wilson, 1994), i.e. of having agreed symbols in predefined and fixed positions: the positions of a string of characters are numbered 0, 1, 2, etc. in the following way:

a. the agreed character at position 0 encodes part-of-speech;

b. each character at position 1, 2, n, encodes the value of one attribute (person, gender, number, etc.);

c. if an attribute does not apply, the corresponding position in the string contains a special marker, the hyphen (’-’).

Example: Ncms- (Noun, common, masculine, singular, nocase)

This notation adopts the Eagles Intermediate Format with a small revision: the Intermediate Format encodes information by means of digits, while in Multext characters of a mnemonic nature are preferred.

The marker ’-’ has a special semantics and it means ’not-applicable’. As stated above, its function is to keep the relationship established between attributes and values. It is used in the following cases:

(a) not relevant to a particular language, e.g. Gender to Estonian.

(b) not applicable to a particular combination of attribute-values, i.e. although the attribute is used by a category in a given language it does not apply to a particular subclass of the category; e.g., Person applies to Pronouns, but not to the Type demonstrative.

(c) not applicable to a particular lexical item, i.e. although the attribute applies to the rest of its paradigm, e.g., Gender in the paradigm of English Personal Pronouns applies only to the 3rd person, I, you vs. she, he).

Finally, it should be noted that in the lexica trailing hyphens have been omitted, as this often leads to a more compact encoding. Hence codes like Ncms- are written as Ncms.

1.4 Organization of the language-specific chapters

The applications of the proposal to the individual eight languages are then given in Chapter II. The language application parts present the lexicon specifications category by category and are structured as follows:

(a) a chapter describing the features and values pertinent to a given category in the form of tables with examples from the language in question; these tables have been mostly automatically generated from the common tables. The language specific tables, however, do often have significant additional information, e.g. examples or localisation of names.

(b) a chapter providing the allowed combinations of values for the particular language, which is, in some cases, supplemented by c) a list of examples from the lexicon or corpus.

Two different strategies of displaying combinations have been followed:

(i) all the admitted combinations are provided together with an example. This has the disadvantage of producing big lists, given the high number of features and values to combine together, but the advantage of providing the only legal combinations with the relevant constraints in the application of some features/values in presence of other features or values or combination of them (see e.g Gender in Personal Pronouns).

(ii) a mathematical expression (essentially a regular expression) describing the combinations is provided, which can subsequently used to generate all combinations, possibly including some which are not valid.

Part I

Common Morphosyntactic Specifications Table of categories

=============== ==== ====
Part-of-Speech  Code Atts
=============== ==== ====
Noun             N    10
Verb             V    14
Adjective        A    12
Pronoun          P    17
Determiner       D    10
Article          T     5
Adverb           R     6
Adposition       S     4
Conjunction      C     7
Numeral          M    12
Interjection     I     2
Residual         X     0
Abbreviation     Y     5
Particle         Q     3
=============== ==== ====

Each category acts a type, i.e. it introduces attribute value pairs. The equivalence of a particular attribute-value at a particular position in a morpho-syntactic description is thus defined only with respect to a given category.

The attributes and the values pertinent to the above categories are presented in the the first 14 sections of the next chapter. The final two sections list the values and attributes defined in the tables.

Tables of attribute-values

The common Multext-East tables of attribute values are given for all categories above and have a rigid structure, which makes them suitable for automatically verifying the conformance of a particular morphosyntactic description with the tables, or for expanding a morphosyntactic description into its more verbose form.

At the beginning of each category the number of attributes introduced by the category is given. This is followed by a string summarizing the valid attributes of the category. The features constituting the minimal core, in common with Multext (and Eagles) are graphically contained between strings of stars (*).

This information is followed by the table proper. The table has the following columns:

1. Position gives the position of the attribute in the string of the morphosyntactic description;

2. Attribute gives the name of the attribute;

3. Value gives the name and one-letter code of the attribute-value;

4. en, ro, sl, cs, bg, et, hu, hr i.e. a column for each of the languages. For easier comparison between them they have been grouped by language family: first English as the ’hub’ language, then Romance (Romanian), followed by Slavic (Slovene, Czech, Bulgarian), and finaly Finno-Ungric (Estonian, Hungarian). Croation has been added later, and not yet grouped.

The language columns define (by marking with ’x’), in the first line, whether the category is used by the language, and in subsequent lines, which attribute-values a particular language uses.

Multext-East / Concede  Morphosyntactic specifications V.2— Common Tables  7

3.1 Noun (N)

11 Positions
**** **** **** **** **** ---- ---- ---- ---- ---- ----
PoS  Type Gend Numb Case Def  Cltc Anim OwnN OwnP OwdN
**** **** **** **** **** ---- ---- ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           common         c  x   x   x   x   x   x   x   x
proper         p  x   x   x   x   x   x   x   x
- -------------- -------------- -
2 Gender         masculine      m  x   x   x   x   x           x
feminine       f  x   x   x   x   x           x
neuter         n  x   x   x   x   x           x
- -------------- -------------- -
3 Number         singular       s  x   x   x   x   x   x   x   x
plural         p  x   x   x   x   x   x   x   x
dual           d          x   x
count          t                  x
- -------------- -------------- -
4 Case           nominative     n          x   x   x   x   x   x
genitive       g          x   x       x   x   x
dative         d          x   x           x   x
accusative     a          x   x           x   x
vocative       v      x       x   x           x
locative       l          x   x               x
instrumental   i          x   x           x   x
direct         r      x
oblique        o      x
partitive      1                      x
illative       x                      x   x
inessive       2                      x   x
elative        e                      x   x
allative       t                      x   x
adessive       3                      x   x
ablative       b                      x   x
translative    4                      x
terminative    9                      x   x
essive         w                      x   x
abessive       5                      x
komitative     k                      x
aditive        7                      x
temporalis     m                          x
causalis       c                          x
sublative      s                          x
delative       h                          x
sociative      q                          x
factive        y                          x
superessive    p                          x

                 distributive   u                          x
* ***************************** *
5 Definiteness   no             n      x           x
                 yes            y      x           x
                 short_art      s                  x
                 full_art       f                  x
- -------------- -------------- -
6 Clitic         no             n      x
                 yes            y      x
- -------------- -------------- -
7 Animate        no             n          x   x           x   x
                 yes            y          x   x           x   x
- -------------- -------------- -
8 Owner_Number   singular       s                          x
                 plural         p                          x
- -------------- -------------- -
9 Owner_Person   first          1                          x
                 second         2                          x
                 third          3                          x
- -------------- -------------- -
10Owned_Number   singular       s                          x
                 plural         p                          x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.2 Verb (V)

15 Positions
**** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ---- ----
PoS  Type VFrm Tens Pers Numb Gend Voic Neg  Def  Cltc Case Anim Clt2 Aspt
**** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           main           m  x   x   x   x   x   x   x   x
auxiliary      a  x   x   x   x   x   x   x   x
modal          o  x   x   x   x       x       x
copula         c      x   x   x               x
base           b  x
- -------------- -------------- -
2 VForm          indicative     i  x   x   x   x   x   x   x   x
subjunctive    s      x
imperative     m      x   x   x   x   x   x   x
conditional    c  x       x   x       x   x   x
infinitive     n  x   x   x   x       x   x   x
participle     p  x   x   x   x   x   x       x
gerund         g      x           x   x
supine         u          x           x
transgressive  t              x
quotative      q                      x
- -------------- -------------- -
3 Tense          present        p  x   x   x   x   x   x   x   x
imperfect      i      x           x   x       x
future         f          x   x               x
past           s  x   x   x   x   x   x   x   x
pluperfect     l      x                       x
aorist         a                  x           x
- -------------- -------------- -
4 Person         first          1  x   x   x   x   x   x   x   x
second         2  x   x   x   x   x   x   x   x
third          3  x   x   x   x   x   x   x   x
- -------------- -------------- -
5 Number         singular       s  x   x   x   x   x   x   x   x
plural         p  x   x   x   x   x   x   x   x
dual           d          x
- -------------- -------------- -
6 Gender         masculine      m      x   x   x   x           x
feminine       f      x   x   x   x           x
neuter         n      x   x   x   x           x
*********************************
7 Voice          active         a          x   x   x   x       x
passive        p          x   x   x   x       x
- -------------- -------------- -
8 Negative       no             n          x   x       x       x
yes            y          x   x       x       x

- -------------- -------------- -
9 Definiteness   no             n                  x       x
yes            y                  x       x
short_art      s                  x
full_art       f                  x
1s2s           2                          x
- -------------- -------------- -
10Clitic         no             n      x
yes            y      x
- -------------- -------------- -
11Case           nominative     n
genitive       g
dative         d
accusative     a
locative       l
instrumental   i
illative       x                      x
inessive       2                      x
elative        e                      x
translative    4                      x
abessive       5                      x
- -------------- -------------- -
12Animate        no             n              x
yes            y              x
- -------------- -------------- -
13Clitic_s       no             n              x
yes            y              x
- -------------- -------------- -
14Aspect         progressive    p          x
perfective     e          x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.3 Adjective (A)

13 Positions
**** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ----
PoS  Type Degr Gend Numb Case Def  Cltc Anim Form OwnN OwnP OwdN
**** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           qualificative  f  x   x   x   x           x   x
indefinite     i
possessive     s          x   x               x
ordinal        o          x
- -------------- -------------- -
2 Degree         positive       p  x   x   x   x       x   x   x
comparative    c  x   x   x   x       x   x   x
superlative    s  x   x   x   x       x   x   x
elative        e          x
- -------------- -------------- -
3 Gender         masculine      m      x   x   x   x           x
feminine       f      x   x   x   x           x
neuter         n      x   x   x   x           x
- -------------- -------------- -
4 Number         singular       s      x   x   x   x   x   x   x
plural         p      x   x   x   x   x   x   x
dual           d          x   x
- -------------- -------------- -
5 Case           nominative     n          x   x       x   x   x
genitive       g          x   x       x   x   x
dative         d          x   x           x   x
accusative     a          x   x           x   x
vocative       v      x       x               x
locative       l          x   x               x
instrumental   i          x   x           x   x
direct         r      x
oblique        o      x
partitive      1                      x
illative       x                      x   x
inessive       2                      x   x
elative        e                      x   x
allative       t                      x   x
adessive       3                      x   x
ablative       b                      x   x
translative    4                      x
terminative    9                      x   x
essive         w                      x   x
abessive       5                      x
komitative     k                      x
aditive        7                      x
temporalis     m                          x

causalis       c                          x
sublative      s                          x
delative       h                          x
sociative      q                          x
factive        y                          x
superessive    p                          x
distributive   u                          x
essive_formal  f                          x
* ***************************** *
6 Definiteness   no             n      x   x       x           x
yes            y      x   x       x           x
short_art      s                  x
full_art       f                  x
- -------------- -------------- -
7 Clitic         no             n      x
yes            y      x
- -------------- -------------- -
8 Animate        no             n          x   x           x   x
yes            y          x   x           x   x
- -------------- -------------- -
9 Formation      nominal        n              x
compound       c              x
- -------------- -------------- -
10Owner_Number   singular       s                          x
plural         p                          x
- -------------- -------------- -
11Owner_Person   first          1                          x
second         2                          x
third          3                          x
---------------- -------------- -
12Owned_Number   singular       s                          x
plural         p                          x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.4 Pronoun (P)

18 Positions
**** **** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ...
PoS  Type Pers Gend Numb Case OwnN OwnG Cltc RefT SynT Def  Anim Clt2
**** **** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ...
... ---- ---- ---- ----
PrFr OwnP OwdN WhT
... ---- ---- ---- ----
=================================  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           personal       p  x   x   x   x   x   x   x   x
demonstrative  d      x   x   x   x   x   x   x
indefinite     i      x   x   x   x   x   x   x
possessive     s  x   x   x   x   x   x   x   x
interrogative  q  x       x   x   x   x   x   x
relative       r  x       x   x   x   x   x   x
exclamative    e
reflexive      x  x   x   x   x   x   x   x   x
reciprocal     y                      x   x
negative       z      x   x   x   x
general        g  x       x   x   x
int_rel        w      x
emphatic       h      x
determinal     m                      x
ex_there       t  x
- -------------- -------------- -
2 Person         first          1  x   x   x   x   x   x   x   x
second         2  x   x   x   x   x   x   x   x
third          3  x   x   x   x   x   x   x   x
- -------------- -------------- -
3 Gender         masculine      m  x   x   x   x   x           x
feminine       f  x   x   x   x   x           x
neuter         n  x   x   x   x   x           x
- -------------- -------------- -
4 Number         singular       s  x   x   x   x   x   x   x   x
plural         p  x   x   x   x   x   x   x   x
dual           d          x   x
- -------------- -------------- -
5 Case           nominative  *  n  x   x1  x   x   x   x   x   x
genitive    *  g      x2  x   x       x   x   x
dative      *  d      x2  x   x   x       x   x
accusative  *  a  x   x1  x   x   x       x   x
vocative       v      x                       x
locative       l          x   x               x
instrumental   i          x   x           x   x
direct      *  r      xR1
oblique     *  o      xR2
partitive      1                      x

illative       x                      x   x
inessive       2                      x   x
elative        e                      x   x
allative       t                      x   x
adessive       3                      x   x
ablative       b                      x   x
translative    4                      x
terminative    9                      x   x
essive         w                      x   x
abessive       5                      x
komitative     k                      x
aditive        7                      x
temporalis     m                          x
causalis       c                          x
sublative      s                          x
delative       h                          x
sociative      q                          x
factive        y                          x
superessive    p                          x
distributive   u                          x
essive_formal  f                          x
- -------------- -------------- -
6 Owner_Number   singular       s  x   x   x   x           x   x
plural         p  x   x   x   x           x   x
dual           d          x
- -------------- -------------- -
7 Owner_Gender   masculine      m  x       x   x               x
feminine       f  x       x   x               x
neuter         n          x   x               x
*********************************  EN  RO  SL  CS  BG  ET  HU  HR
8 Clitic         no             n      x   x   x   x           x
yes            y      x   x   x   x           x
bound          b          x
- -------------- -------------- -
9 Referent_Type  personal       p          x   x   x           x
possessive     s          x   x   x           x
attributive    a                  x
quantitative   q                  x
- -------------- -------------- -
10Syntactic_Type nominal        n          x   x               x
adjectival     a          x   x               x
adverbial      r          x
- -------------- -------------- -
11Definiteness   no             n                  x
yes            y                  x
short_art      s                  x
full_art       f                  x
- -------------- -------------- -
12Animate        no             n          x   x               x
yes            y          x   x               x
- -------------- -------------- -
13Clitic_s       yes            y              x
no             n              x

- -------------- -------------- -
14Pronoun_Form   strong         s      x
                 weak           w      x
- -------------- -------------- -
15Owner_Person   first          1                          x
                 second         2                          x
                 third          3                          x
- -------------- -------------- -
16Owned_Number   singular       s                          x
                 plural         p                          x
- -------------- -------------- -
17Wh_Type        relative       r  x
                 question       q  x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

Notes

* In the Romanian case system the value ’direct’ conflates ’nominative’ and ’accusative’, while the value ’oblique’ conflates ’genitive’ and ’dative’.

If we want to list these values as belonging to the same Case universe, i.e. if we want either ’direct’, ’nominative’ and ’accusative’ or ’oblique’, ’genitive’ and ’dative’ appear all together in the same system, their semantics should be explained and their internal relationships represented. We suggest here to follow the same strategy adopted within the Eagles validation exercise, where different kinds of relationships between values have been formally represented by adding indexes to the values of interest:

- in this example, the kind of relation ’direct’ and ’oblique’ have with other values of the system is of ’replacement’. We mark this labelling ’direct’ and ’oblique’ with the index ’R’ and adding the numerical indexes 1 to ’direct’ and 2 to ’oblique’, which point to the values which are respectively replaced.

3.5 Determiner (D)

11 Positions
**** **** **** **** **** **** **** **** ---- ---- ----
PoS  Type Pers Gend Numb Case OwnN OwnG Cltc Mod  WhT
**** **** **** **** **** **** **** **** ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   -   -   -   -   -   -
= ============== ============== =
1 Type           demonstrative  d  x   x
                 indefinite     i  x   x
                 possessive     s  x   x
                 interrogative  q
                 relative       r
                 exclamative    e
                 article        a
                 general        g  x
                 int_rel        w      x
                 negative       z      x
                 emphatic       h      x
- -------------- -------------- -
2 Person         first          1  x   x
                 second         2  x   x
                 third          3  x   x
- -------------- -------------- -
3 Gender         masculine      m      x
                 feminine       f      x
                 neuter         n      x
- -------------- -------------- -
4 Number         singular       s  x   x
                 plural         p  x   x
- -------------- -------------- -
5 Case           direct         r      x
                 oblique        o      x
- -------------- -------------- -
6 Owner_Number   singular       s  x   x
                 plural         p  x   x
- -------------- -------------- -
7 Owner_Gender   masculine      m  x
                 feminine       f  x
                 neuter         n  x
*********************************
8 Clitic         no             n      x
                 yes            y      x
- -------------- -------------- -
9 Modific_Type   prenomin       e      x
                 postnomin      o      x
- -------------- -------------- -
10Wh_Type        relative       r  x
                 question       q  x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.6 Article (T)

6 Positions
**** **** **** **** **** ----
PoS  Type Gend Numb Case Cltc
**** **** **** **** **** ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  -   x   -   -   -   -   x   -
= ============== ============== =
1 Type           definite       f      x                   x
                 indefinite     i      x                   x
                 possessive     s      x
                 demonstrative  d      x
- -------------- -------------- -
2 Gender         masculine      m      x
                 feminine       f      x
                 neuter         n      x
- -------------- -------------- -
3 Number         singular       s      x
                 plural         p      x
- -------------- -------------- -
4 Case           direct         r      x
                 oblique        o      x
*********************************
5 Clitic         no             n      x
                 yes            y      x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.7 Adverb (R)

7 Positions
**** **** **** ---- ---- ---- ----
PoS  Type Degr Cltc Numb Pers WhT
**** **** **** ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           general        g      x   x   x   x       x   x
                 particle       p      x                   x
                 causal         o                          x
                 negative       z      x
                 adjectival     a                  x
                 verbal         v                          x
                 modifier       m  x   x                   x
                 specifier      s  x
                 int_rel        w      x
                 portmanteau    c      x
                 interrogative  q                          x
- -------------- -------------- -
2 Degree         positive       p  x   x   x   x               x
                 comparative    c  x   x   x   x               x
                 superlative    s  x   x   x   x               x
                 elative        e          x
*********************************
3 Clitic         no             n      x                   x
                 yes            y      x                   x
- -------------- -------------- -
4 Number         singular       s                          x
                 plural         p                          x
- -------------- -------------- -
5 Person         first          1                          x
                 second         2                          x
                 third          3                          x
- -------------- -------------- -
6 Wh_Type        relative       r  x
                 question       q  x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.8 Adposition (S)

5 Positions
**** **** **** ---- ----
PoS  Type Form Case Cltc
**** **** **** ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           preposition    p  x   x   x   x   x   x       x
                 postposition   t  x                   x   x
- -------------- -------------- -
2 Formation      simple         s      x   x   x               x
                 compound       c      x   x   x               x
*********************************
3 Case           nominative     n
  (req.by prep.) genitive       g      x   x   x               x
                 dative         d      x   x   x               x
                 accusative     a      x   x   x               x
                 locative       l          x   x               x
                 instrumental   i          x   x               x
- -------------- -------------- -
4 Clitic         no             n      x
                 yes            y      x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.9 Conjunction (C)

8 Positions
**** **** ---- ---- ---- ---- ---- ----
PoS  Type Form CTyp SubT Cltc Numb Pers
**** **** ---- ---- ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           coordinating   c  x   x   x   x   x   x   x   x
                 subordinating  s  x   x   x   x   x   x   x   x
                 portmanteau    r      x
*********************************
2 Formation      simple         s      x   x       x       x   x
                 compound       c      x   x       x       x   x
- -------------- -------------- -
3 Coord_Type     simple         s      x
                 repetit        r      x
                 correlat       c      x
                 sentence       p                          x
                 words          w                          x
                 initial        i  x
                 non_initial    n  x
- -------------- -------------- -
4 Sub_Type       negative       z      x
                 positive       p      x
- -------------- -------------- -
5 Clitic         no             n      x
                 yes            y      x
- -------------- -------------- -
6 Number         singular       s              x
                 plural         p              x
- -------------- -------------- -
7 Person         first          1              x
                 second         2              x
                 third          3              x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.10 Numeral (M)

13 Positions
**** **** **** **** **** ---- ---- ---- ---- ---- ---- ---- ----
PoS  Type Gend Numb Case Frm  Def  Cltc Clas Anim OwnN OwnP OwdN
**** **** **** **** **** ---- ---- ---- ---- ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           cardinal       c  x   x   x   x   x   x   x   x
ordinal        o  x   x   x   x   x   x   x   x
fractal        f      x                   x
multiple       m      x   x   x               x
collect        l      x                   x
special        s          x   x               x
- -------------- -------------- -
2 Gender         masculine      m      x   x   x   x           x
feminine       f      x   x   x   x           x
neuter         n      x   x   x   x           x
- -------------- -------------- -
3 Number         singular       s      x   x   x   x   x   x   x
plural         p      x   x   x   x   x   x   x
dual           d          x   x
- -------------- -------------- -
4 Case           nominative     n          x   x       x   x   x
genitive       g          x   x       x   x   x
dative         d          x   x           x   x
accusative     a          x   x           x   x
vocative       v                              x
locative       l          x   x               x
instrumental   i          x   x           x   x
direct         r      x
oblique        o      x
partitive      1                      x
illative       x                      x   x
inessive       2                      x   x
elative        e                      x   x
allative       t                      x   x
adessive       3                      x   x
ablative       b                      x   x
translative    4                      x
terminative    9                      x   x
essive         w                      x   x
abessive       5                      x
komitative     k                      x
aditive        7                      x
temporalis     m                          x
causalis       c                          x
sublative      s                          x
delative       h                          x

sociative      q                          x
factive        y                          x
superessive    p                          x
distributive   u                          x
essive_formal  f                          x
multiplicative 6                          x
*********************************
5 Form           digit          d      x   x   x   x   x   x   x
roman          r      x   x   x   x   x   x   x
letter         l      x   x   x   x   x   x   x
both           b      x
m_form         m                  x
approx         a                  x
- -------------- -------------- -
6 Definiteness   no             n      x   x       x
yes            y      x   x       x
short_art      s                  x
full_art       f                  x
- -------------- -------------- -
7 Clitic         no             n      x
yes            y      x
- -------------- -------------- -
8 Class          definite1      1              x
definite2      2              x
definite34     3              x
definite       f              x
demonstrative  d              x
indefinite     i              x
interrogative  q              x
relative       r              x
- -------------- -------------- -
9 Animate        no             n          x   x               x
yes            y          x   x               x
- -------------- -------------- -
10Owner_Number   singular       s                          x
plural         p                          x
- -------------- -------------- -
11Owner_Person   first          1                          x
second         2                          x
third          3                          x
- -------------- -------------- -
12Owned_Number   singular       s                          x
plural         p                          x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.11 Interjection (I)

3 Positions
**** ---- ----
PoS  Type Form
**** ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Type           mood           m                          x
                 other          o                          x
- -------------- -------------- -
2 Formation      simple         s                  x           x
                 compound       c                  x           x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.12 Residual (X)

1 Position
****
PoS
****
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.13 Abbreviation (Y)

6 Positions
**** ---- ---- ---- ---- ----
PoS  SynT Gend Numb Case Def
**** ---- ---- ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  x   x   x   x   x   x   x   x
= ============== ============== =
1 Syntactic_Type nominal        n      x               x       x
verbal         v      x               x
adjectival     a      x               x
adverbial      r      x               x       x
pronominal     p      x
- -------------------------------
2 Gender         masculine      m      x                       x
feminine       f      x                       x
neuter         n      x                       x
- -------------- -------------- -
3 Number         singular       s      x               x       x
plural         p      x               x       x
- -------------- -------------- -
4 Case           nominative     n                      x       x
genitive       g                      x       x
dative         d                              x
accusative     a                              x
locative       l                              x
instrumental   i                              x
direct         r      x
oblique        o      x
vocative       v      x
partitive      1                      x
illative       x                      x
inessive       2                      x
elative        e                      x
allative       t                      x
adessive       3                      x
ablative       b                      x
translative    4                      x
terminative    9                      x
essive         w                      x
abessive       5                      x
komitative     k                      x
aditive        7                      x
- -------------- -------------- -
5 Definiteness   yes            y      x
no             n      x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.14 Particle (Q)

5 Positions
**** ---- ---- ----
PoS  Type Form Cltc
**** ---- ---- ----
= ============== ============== =  EN  RO  SL  CS  BG  ET  HU  HR
P ATT            VAL            C  -   x   x   x   x   -   -   x
= ============== ============== =
1 Type           negative       z      x           x           x
                 infinitive     n      x
                 subjunctive    s      x
                 aspect         a      x
                 future         f      x
                 general        g                  x
                 comparative    c                  x
                 verbal         v                  x
                 interrogative  q                  x           x
                 modal          o                  x           x
                 affirmative    r                              x
- -------------- -------------- -
2 Formation      simple         s                  x
                 compound       c                  x
- -------------- -------------- -
3 Clitic         no             n      x
                 yes            y      x
=================================  EN  RO  SL  CS  BG  ET  HU  HR

3.15 List of values with codes

The values presented within the tables are, in the following, listed in alphabetical order; the first column gives the name of the value, the second column its code and the third lists attributes for which the value is appropriate.

1s2s           2    Definiteness
abessive       5    Case
ablative       b    Case
accusative     a    Case
active         a    Voice
adessive       3    Case
aditive        7    Case
adjectival     a    Syntactic_Type, Type
adverbial      r    Syntactic_Type
allative       t    Case
aorist         a    Tense
approx         a    Form
article        a    Type
aspect         a    Type
attributive    a    Referent_Type
auxiliary      a    Type
base           b    VForm
both           b    Form
bound          b    Clitic
cardinal       c    Type
causal         o    Type
causalis       c    Case
collect        l    Type
common         c    Type
comparative    c    Degree, Type
compound       c    Formation
conditional    c    VForm
coordinating   c    Type
copula         c    Type
correlat       c    Coord_Type
count          t    Number
dative         d    Case
definite       f    Class, Type
definite1      1    Class
definite2      2    Class
definite34     3    Class
delative       h    Case
demonstrative  d    Class, Type
determinal     m    Type
digit          d    Form
direct         r    Case
distributive   u    Case
dual           d    Number, Owner_Number
elative        e    Case, Degree
emphatic       h    Type
essive         w    Case
essive_formal  f    Case
exclamative    e    Type
ex_there       t    Type/P
factive        y    Case
feminine       f    Gender, Owner_Gender
first          1    Owner_Person, Person
fractal        f    Type
full_art       f    Definiteness
future         f    Tense, Type
general        g    Type
genitive       g    Case
gerund         g    VForm

illative       x    Case
imperative     m    VForm
imperfect      i    Tense
indefinite     i    Class, Type
indicative     i    VForm
inessive       2    Case
infinitive     n    Type, VForm
initial        i    Coord_Type
instrumental   i    Case
int_rel        w    Type
interrogative  q    Class, Type
komitative     k    Case
letter         l    Form
locative       l    Case
m_form         m    Form
main           m    Type
masculine      m    Gender, Owner_Gender
modal          o    Type
modifier       m    Type
mood           m    Type
multiple       m    Type
multiplicative 6    Case
negative       z    Sub_Type, Type
neuter         n    Gender, Owner_Gender
no             n    Animate, Clitic, Clitic_s, Definiteness, Negative
nominal        n    Formation, Syntactic_Type
nominative     n    Case
non_initial    n    Coord_Type
oblique        o    Case
ordinal        o    Type
other          o    Type
participle     p    VForm
particle       p    Type
partitive      1    Case
passive        p    Voice
past           s    Tense
perfective     e    Aspect
personal       p    Referent_Type, Type
pluperfect     l    Tense
plural         p    Number, Owned_Number, Owner_Number
portmanteau    c    Type (Adverb)
portmanteau    r    Type (Conjunction)
positive       p    Degree, Sub_Type
possessive     s    Referent_Type, Type
postnomin      o    Modific_Type
postposition   t    Type
prenomin       e    Modific_Type
preposition    p    Type
present        p    Tense
progressive    p    Aspect
pronominal     p    Syntactic_Type
proper         p    Type
qualificative  f    Type
quantitative   q    Referent_Type
question       q    Wh_Type
quotative      q    VForm
reciprocal     y    Type
reflexive      x    Type
relative       r    Class, Type, Wh_Type
repetit        r    Coord_Type
roman          r    Form
second         2    Owner_Person, Person
sentence       p    Coord_Type
short_art      s    Definiteness
simple         s    Coord_Type, Formation
singular       s    Number, Owned_Number, Owner_Number

sociative      q    Case
special        s    Type/M
specifier      s    Type/R
strong         s    Pronoun_Form
subjunctive    s    Type, VForm
sublative      s    Case
subordinating  s    Type
superessive    p    Case
superlative    s    Degree
supine         u    VForm
temporalis     m    Case
terminative    9    Case
third          3    Owner_Person, Person
transgressive  t    VForm
translative    4    Case
verbal         v    Type
vocative       v    Case
weak           w    Pronoun_Form
words          w    Coord_Type
yes            y    Animate, Clitic, Clitic_s, Definiteness, Negative

3.16 List of attributes

In this section, all the attributes presented in the tables are listed in alphabetical order. For some attributes which are not self-explanatory, a brief description of their semantics is provided.

Attribute  Relevant to Category
Animate  Adj Noun Num Pron Verb
Aspect  Verb
Case  Abbr Adj Adpos Art Det Noun Num Pron Verb
Class  Num
Clitic  Adj Adpos Adv Art Conj Det Noun Num Part Pron Verb
Clitic_s  Pron
Coord_Type  Conj
Definiteness  Abbr Adj Noun Num Pron Verb
Degree  Adj Adv
Form  Num
Formation  Adj Adpos Conj Interj Part
Gender  Abbr Adj Art Det Noun Num Pron Verb
Modific_Type  Det
Negative  Verb
Number  Abbr Adj Adv Art Conj Det Noun Num Pron Verb
Owned_Number Adj Noun Num Pron
Owner_Gender Det Pron
Owner_Number Adj Det Noun Num Pron
Owner_Person Adj Noun Num Pron
Person  Adv Conj Det Pron Verb
Pronoun_Form Pron
Referent_Type Pron
Sub_Type  Conj
Syntactic_Type Abbr Pron
Tense  Verb
Type  Adj Adpos Adv Art Conj Det Interj Noun Num Part Pron Verb
VForm  Verb
Voice  Verb
Wh_Form  Adv Det Pron

Class: distinguishes subtypes of Numerals in Czech which have a distinct syntactic distributions: e.g. subclasses for 1, 2, 3&4, etc. are distinguished.

Clitic_s: the ’yes’ value of the Clitic_s attribute denotes Czech pronouns having the clitic morpheme ’s’ appended as a suffix.

Definiteness: corresponds to the definite and indefinite article in English, which is expressed in the Slavic languages by suffixes. For Bulgarian singular masculine there are two forms: full article and short article (full is used when a sing.masc. form is the syntactic subject of the clause, otherwise short article is used). The distinction full vs. short is not made for feminine, neuter and plural forms. Definiteness is also used in Romanian.

Form: used to distinguish different forms of numerals (Roman, digit, ’letter’) and in Bulgarian also for the semantic distinction between numerals that refer to male persons (but not children) or groups of male+female.

Formation: refers to the graphical components: simple, i.e. consisting of one word; compound, i.e. consisting of more than one word.

Modific_Type: refers to the prenominal or postnominal positions of Determiners which distinguish different forms in Romanian.

Negative: the value ’yes’ encodes negative verbal word-forms in Czech, Slovene and in Estonian.

Owned_Number: in the Hungarian system, different word-forms are distinguished for nominals on the basis of so called ’anaphoric possessive’ number, i.e. the number of the thing(s) possessed by the nominal in question.

Owner_Gender: used to encode the Gender of the possessor in Pronouns and (in Romanian) Determiners.

Owner_Number: used to specify the possessor number in Pronouns, as well as (in Romanian) in Determiners, and (in Hungarian) in Adjectives and Nouns.

Owner_Person: used to specify the possessor person in in Hungarian in Adjectives and Nouns.

Pronoun_Form: used to encode weak and strong pronouns in Romanian.

Referent_Type: used to distinguish reflexive personal from reflexive possessive pronouns in the Slavic languages. In Bulgarian, it also describes a subdivision on the basis of semantic features which have effect on the morpho-syntactic paradigm, e.g. quantitative: the pronoun refers to quantity, etc.

Sub_Type: used in Romanian to distinguish negative from positive Conjunctions.

Syntactic_Type: used to distinguish the nominal and adjectival function of Pronouns in Slovene and Czech. Also used in Abbreviations to signal the Part of Speech of the abbreviation; currently used only by Romanian and Estonian.

References

Tomaž Erjavec, Monica Monachini (eds.): Report MTE D1.1 F, Specifications and Notation for Lexicon Encoding. MULTEXT-East Final Report, D1.1, December 1997, Institute Jozef Stefan, Ljubljana. ⟨http://nl.ijs.si/ME/CD/docs/mte-d11f/ ⟩

Tomaž Erjavec, Ann Lawson, Laurent Romary (eds.) (1998): East meets West: A Compendium of Multilingual Resources. CD-ROM, produced and distributed by TELRI Association e.V., ISBN: 3-922641-46-6. Vol II (MULTEXT-East results) is available at ⟨http://nl.ijs.si/ME/CD/ ⟩

Bel N., Calzolari N. and Monachini M. eds. (1995): “Common Specifications and Notation for Lexicon Encoding and Preliminary Proposal for the Tagsets”, MULTEXT Deliverable D1.6.1B, Pisa.

Leech J. and A. Wilson (1994): “Morphosyntactic Annotation”, EAGLES Interim Report, Pisa.

Monachini M. and N. Calzolari (1995): “Synopsis and Comparison of Morphosyntactic Phenomena Encoded in Lexicons and in Corpora and Application to European Languages”, EAGLES document EAG-LSG-T4.6/CSG-T3.2, Pisa.

MULTEXT (1993): “MULTEXT Technical Annex”, Aix-en-Provence.

MULTEXT WP1.6 Report A2 (1994): “Common Specifications and Notation for Lexicon Encoding”, MULTEXT Report Milestone A2.

Part II

Language Specific Applications Application to English

Authors: Nancy Ide, Greg Priest-Dorman, Tomaž Erjavec.

The specification is based on the MULTEXT application to English, detailed in MULTEXT Milestone B Deliverable D1.6.1B, March 1995.

4.1 Noun (N)

1.1. Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           common         c
                 proper         p
- -------------- -------------- -
2 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case                          -
* ***************************** *
5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -
- -------------- -------------- -
7 Animate                       -
- -------------- -------------- -
8 Owner_Number                  -
- -------------- -------------- -
9 Owner_Person                  -
---------------- -------------- -

Multext-East / Concede  Morphosyntactic specifications V.2— English  34

10Owned_Number                  -
=================================

1.2 Combinations
Lexicon   Definition
------------------------------------
Nc        Noun common
Nc-p      Noun common plural
Nc-s      Noun common singular
Ncf       Noun common feminine
Ncfp      Noun common feminine plural
Ncfs      Noun common feminine singular
Ncm       Noun common masculine
Ncmp      Noun common masculine plural
Ncms      Noun common masculine singular
Ncn       Noun common neuter
Ncnp      Noun common neuter plural
Ncns      Noun common neuter singular
Np        Noun proper
Np-p      Noun proper plural
Np-s      Noun proper singular
Npfs      Noun proper feminine singular
Npms      Noun proper masculine singular
Npnp      Noun proper neuter plural
Npns      Noun proper neuter singular

4.2 Verb (V)

2.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           main           m
                 auxiliary      a
                 modal          o
           l.s.  base           b
- -------------- -------------- -
2 VForm          indicative     i
                 conditional    c
                 infinitive     n
                 participle     p
- -------------- -------------- -
3 Tense          present        p
                 past           s
- -------------- -------------- -
4 Person         first          1
                 second         2

                 third          3
- -------------- -------------- -
5 Number         singular       s
                 plural         p
- -------------- -------------- -
6 Gender                        -
*********************************
7 Voice                         -
- -------------- -------------- -
8 Negative                      -
- -------------- -------------- -
9 Definiteness                  -
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
=================================
Notes:
These tags  do  not  attempt to represent  distinctions found  in  the
various compound  verb-forms.  These  are  composed  of  a sequence of
auxiliary and non-finite verb as follows:
        future             will/shall + base
        conditional        would/should + base
        passive            be + past participle
        perfect            have + past participle
        past perfect       have-past + past participle
        present continuous be + present participle
        infinitive         to + base
So there is no  "aspect"  attribute or  "future"  value,  for example.
Examples:
    go     Vmn, Vmip-p, Vmip1s, Vmip2s
    goes   Vmip3s
    going  Vmpp
    gone   Vmps
    went   Vmis
    have   Van, Vmn, Vaip-p, Vaip1s, Vaip2s, Vmip-p, Vmip1s, Vmip2s
    has    Vaip3s, Vmip3s
    had    Vais, Vmis, Vmps
    having Vapp, Vmpp
    be     Van, Vmnp
    am     Vaip1s, Vmip1s
    are    Vaip-p, Vaip2s, Vmip-p, Vmip2s
    is     Vaip3s, Vmip3s
    was    Vais1s, Vais3s, Vmis1s, Vmis3s
    were   Vacs, Vais-p, Vais2s, Vmis-p, Vmis2s,

    been   Vais, Vaps
    being  Vapp
    will   Vmn, Vmip-p, Vmip1s, Vmip2s, Voip
    would  Voip

2.2 Combinations
Lexicon   Definition
------------------------------------
Vacs      Verb auxiliary conditional past
Vaip-p    Verb auxiliary indicative present plural
Vaip1s    Verb auxiliary indicative present first singular
Vaip2s    Verb auxiliary indicative present second singular
Vaip3s    Verb auxiliary indicative present third singular
Vais      Verb auxiliary indicative past
Vais-p    Verb auxiliary indicative past plural
Vais1s    Verb auxiliary indicative past first singular
Vais2s    Verb auxiliary indicative past second singular
Vais3s    Verb auxiliary indicative past third singular
Van       Verb auxiliary infinitive
Vapp      Verb auxiliary participle present
Vaps      Verb auxiliary participle past
Vmcs      Verb main conditional past
Vmip      Verb main indicative present
Vmip-p    Verb main indicative present plural
Vmip1s    Verb main indicative present first singular
Vmip2s    Verb main indicative present second singular
Vmip3s    Verb main indicative present third singular
Vmis      Verb main indicative past
Vmis-p    Verb main indicative past plural
Vmis1s    Verb main indicative past first singular
Vmis2s    Verb main indicative past second singular
Vmis3s    Verb main indicative past third singular
Vmn       Verb main infinitive
Vmnp      Verb main infinitive present
Vmpp      Verb main participle present
Vmps      Verb main participle past
Voip      Verb modal indicative present

4.3 Adjective (A)

3.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           qualificative  f
- -------------- -------------- -
2 Degree         positive       p

                 comparative    c
                 superlative    s
- -------------- -------------- -
3 Gender                        -
- -------------- -------------- -
4 Number                        -
- -------------- -------------- -
5 Case                          -
* ***************************** *
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Animate                       -
- -------------- -------------- -
9 Formation                     -
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
---------------- -------------- -
12Owned_Number                  -
=================================

Notes:
Gender, number and case irrelevant for English.
Attributive/predicative distinction  reflects positional constraints -
some adjectives  ('mere',   'utter',  etc.)  only appear in prenominal
position,  while others  ('awake',   'devoid',  etc.)  only appear  in
predicative position. Most can appear in either.
Since  many  English  comparatives  &  superlatives  are  formed  with
more/most,  "positive"  cannot be interpreted as  "neither comparative
nor superlative". See "Adverbs".
Examples:
    big             Af, Afp
    bigger          Afc
    biggest         Afs
    more peculiar   Rsc + Afp
    most remarkable Rss + Afp
    awake           Afp
    mere            Afp

3.2 Combinations
Lexicon   Definition
------------------------------------
Af        Adjective qualificative

Afc       Adjective qualificative comparative
Afp       Adjective qualificative positive
Afs       Adjective qualificative superlative

4.4 Pronoun (P)

4.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           personal       p
                 possessive     s
                 interrogative  q
                 relative       r
                 reflexive      x
                 general        g
           l.s.  ex_there       t
- -------------- -------------- -
2 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
3 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Case           nominative     n
                 accusative     a
- -------------- -------------- -
6 Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
7 Owner_Gender   masculine      m
                 feminine       f
*********************************
8 Clitic                        -
- -------------- -------------- -
9 Referent_Type                 -
- -------------- -------------- -
10Syntactic_Type                -
- -------------- -------------- -
11Definiteness                  -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -

- -------------- -------------- -
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person                  -
- -------------- -------------- -
16Owned_Number                  -
- -------------- -------------- -
17Wh_Type        relative       r
             question       q
=================================

Notes:
"General"  pronouns  are  those  which are  not personal,  possessive,
demonstrative  or reflexive.  The  choice of these  four categories is
based  on  distributional  facts,  though  at  a  rather high level of
abstraction.   They  enter  into  anaphoric  dependencies   which  are
signalled morphosyntactically  and are therefore  (in principle)  more
amenable  to  automatic  detection.  Most  general  pronouns  do  not,
although they too sometimes encode number information.
"WH"  attribute added to allow for combination of possessive and WH in
"whose".
Examples:
Pg3n           some, all, ...
Pg3ns          each, something, nothing, everything, -body, ...
Pg3np          both, ...
Pg3ns          this, that
Pg3np          these, those
Ps3---s        mine
Ps3            yours
Ps3---sm       his
Ps3---sf       hers
?              its
Ps3---p        ours, theirs
Ps3--------------q Ps3--------------r whose
Pp1-sn         I
Pp2            you
Pp3msn         he
Pp3fsn         she
Pp3ns          it
Pp1-pn         we
Pp--pn         they
Pp--sn         me
Pp3msa         him
Pp3fsa         her
Pp1-pa         us
Pp3-pa         them
Px1-s          myself
Px2-s          yourself
Px3ms          himself

    Px3fs          herself
    Px3ns          itself
    Px1-p          ourselves
    Px2-p          yourselves
    Px3-p          themselves
    Pg3--------------q Pg3--------------r  which, whatever
    Pg3--------------q what

4.2 Combinations
Lexicon             Definition
---------------------------------------------------------------------------
Pg3--------------q  Pronoun general third question
Pg3--------------r  Pronoun general third relative
Pg3-p               Pronoun general third plural
Pg3-s               Pronoun general third singular
Pg3fs               Pronoun general third feminine singular
Pg3ms               Pronoun general third masculine singular
Pg3n                Pronoun general third neuter
Pg3np               Pronoun general third neuter plural
Pg3ns               Pronoun general third neuter singular
Pp---------------q  Pronoun personal question
Pp---------------r  Pronoun personal relative
Pp---a-----------q  Pronoun personal accusative question
Pp---a-----------r  Pronoun personal accusative relative
Pp--pn              Pronoun personal plural nominative
Pp--sn              Pronoun personal singular nominative
Pp1-pa              Pronoun personal first plural accusative
Pp1-pn              Pronoun personal first plural nominative
Pp1-sn              Pronoun personal first singular nominative
Pp2                 Pronoun personal second
Pp2-p               Pronoun personal second plural
Pp3-pa              Pronoun personal third plural accusative
Pp3fs               Pronoun personal third feminine singular
Pp3fsa              Pronoun personal third feminine singular accusative
Pp3fsn              Pronoun personal third feminine singular nominative
Pp3ms               Pronoun personal third masculine singular
Pp3msa              Pronoun personal third masculine singular accusative
Pp3msn              Pronoun personal third masculine singular nominative
Pp3ns               Pronoun personal third neuter singular
Ps3                 Pronoun possessive third
Ps3--------------q  Pronoun possessive third question
Ps3--------------r  Pronoun possessive third relative
Ps3---p             Pronoun possessive third plural
Ps3---s             Pronoun possessive third singular
Ps3---sf            Pronoun possessive third singular feminine
Ps3---sm            Pronoun possessive third singular masculine
Pt3                 Pronoun ex_there third
Px1-p               Pronoun reflexive first plural
Px2-p               Pronoun reflexive second plural
Px2-s               Pronoun reflexive second singular
Px3-p               Pronoun reflexive third plural

Px3-s               Pronoun reflexive third singular
Px3fs               Pronoun reflexive third feminine singular
Px3ms               Pronoun reflexive third masculine singular
Px3ns               Pronoun reflexive third neuter singular

4.5 Determiner (D)

5.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           demonstrative  d
                 indefinite     i
                 possessive     s
           l.s.  general        g
- -------------- -------------- -
2 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
3 Gender                        -
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Case                          -
- -------------- -------------- -
6 Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
7 Owner_Gender   masculine      m
                 feminine       f
                 neuter         n
*********************************
8 Clitic                        -
- -------------- -------------- -
9 Modific_Type                  -
- -------------- -------------- -
10Wh_Type        relative       r
                 question       q
=================================

Notes:
Definite and indefinite articles  represented  as  values  of  "Type".
Examples:
    Dd            the

    Di            a/an
    Dg--s         this, that
    Dg--p         these, those
    Ds1---s       my
    Ds2           your
    Ds3---sm      his
    Ds3---sf      her
    Ds3---sn      its
    Ds1---p       our
    Ds3---p       their
    Ds--------q Ds--------r   whose
    Dg--------q Dg--------r   which
    Dg--------q               what
    Dg            each, some, ...
    Dg--p         all, both, certain, many, ...

5.2 Combinations
Lexicon       Definition
-----------------------------------------------
Dd            Determiner demonstrative
Dg            Determiner general
Dg--------q   Determiner general question
Dg--------r   Determiner general relative
Dg--p         Determiner general plural
Dg--s         Determiner general singular
Di            Determiner indefinite
Ds--------q   Determiner possessive question
Ds--------r   Determiner possessive relative
Ds1---p       Determiner possessive first plural
Ds1---s       Determiner possessive first singular
Ds2           Determiner possessive second
Ds3---p       Determiner possessive third plural
Ds3---sf      Determiner possessive third singular feminine
Ds3---sm      Determiner possessive third singular masculine
Ds3---sn      Determiner possessive third singular neuter

4.6 Article (T)

Not applicable.

4.7 Adverb (R)

7.1 Lexicon
= ============== ============== =
P ATT            VAL            C

= ============== ============== =
1 Type
                 modifier       m
           l.s.  specifier      s
- -------------- -------------- -
2 Degree         positive       p
                 comparative    c
                 superlative    s
*********************************
3 Clitic                        -
- -------------- -------------- -
4 Number                        -
- -------------- -------------- -
5 Person                        -
- -------------- -------------- -
6 Wh_Type        relative       r
                 question       q
=================================

Notes:
No distinction has  been made  between different types  of  "modifier"
adverbs  ("sentence-modifying",   "VP-modifying",  etc.),  since their
distributions overlap considerably.

Examples:
    Rmp Rsp   so, too, very, as
    R-p---q   how
    Rmp Rsc   more
    Rmp Rss   most
    Rmp       quickly, soon, here, then, now
    Rmc       better, worse
    Rms       best, worst
    Rmp---q Rmp---r where, when,
    R-p---q   how
    Rmp---q   why

7.2 Combinations
Lexicon   Definition
------------------------------------
R-p---q   Adverb positive question
Rmc       Adverb modifier comparative
Rmp       Adverb modifier positive
Rmp---q   Adverb modifier positive question
Rmp---r   Adverb modifier positive relative
Rms       Adverb modifier superlative
Rsc       Adverb specifier comparative
Rsp       Adverb specifier positive
Rss       Adverb specifier superlative

4.8 Adposition (S)

8.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           preposition    p
                 postposition   t
- -------------- -------------- -
2 Formation                     -
*********************************
3 Case                          -
- -------------- -------------- -
4 Clitic                        -
=================================

Notes:
Postpositions are rare in English.
"possessive" 's and ' might be considered postpositions, especially if
the alternative  is  to  assign  them to the  unique membership  class
(where by definition they would be unrelated).
Examples:
    Sp       in, near, behind,...
    St       notwithstanding, ago

8.2 Combinations
Lexicon   Definition
------------------------------------
S         Adposition
Sp        Adposition preposition
St        Adposition postposition

4.9 Conjunction (C)

9.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           coordinating   c
                 subordinating  s
*********************************
2 Formation                     -

- -------------- -------------- -
3 Coord_Type l.s
            l.s. initial        i
            l.s. non_initial    n
- -------------- -------------- -
4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- -
6 Number                        -
- -------------- -------------- -
7 Person                        -
=================================

Notes:
Subordinating   conjunctions  are  often  identical   to  prepositions
(before, after, since, ...).
Examples:
    Cc-n  and, or, but
    Cc-i  either, neither, both
    Cs    for, that, because, ...

9.2 Combinations
Lexicon   Definition
------------------------------------
Cc        Conjunction coordinating
Cc-i      Conjunction coordinating initial
Cc-n      Conjunction coordinating non_initial
Cs        Conjunction subordinating

4.10 Numeral (M)

10.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           cardinal       c
                 ordinal        o
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -

*********************************
5 Form                          -
- -------------- -------------- -
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Class                         -
- -------------- -------------- -
9 Animate                       -
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================

Notes:
These have not been subsumed under adjectives,  pronouns, determiners,
etc.   because  the   internal   structure  of  complex   numerals  is
idiosyncratic.
Examples:
    Mc    six
    Mo    sixth

10.2 Combinations
Lexicon   Definition
------------------------------------
Mc        Numeral cardinal
Mo        Numeral ordinal

4.11 Interjection (I)

11.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type                          -
- -------------- -------------- -
2 Formation                     -
=================================

11.2 Combinations
Lexicon   Definition
------------------------------------
I         Interjection

4.12 Residual (X)

Not applicable.

4.13 Abbreviation (Y)

13.1 Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Syntactic_Type                -
- -------------------------------
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
- -------------- -------------- -
5 Definiteness                  -
=================================

10.2 Combinations
Lexicon   Definition
------------------------------------
Y         Abbreviation

4.14 Particle (Q)

Not applicable.

Application to Romanian

Authors: D.Tufiş and A.M.Barbu.

5.1 Noun (N)

In Romanian the following attribute-value pairs are applicable to Nouns:

1.1 Type
Attribute           value                Ro. example
Type                common  (c)          carte
                    proper  (p)          Ion

1.2 Gender
Attribute           value                Ro. example
Gender              masculine  (m)       băiatul
                    feminine   (f)       casa
                    neuter     (n)       fir (m.sg.),fire (f.pl.)
In Romanian the declension of a neuter noun always follows in singular
a masculine paradigm and in plural a feminine one. Specific
implementations could take advantage of this rule and by organizing
the paradigmatic space in partial paradigms (masc-sing, masc-pl,
fem-sing, fem-pl) to get rid of neuter value for the gender attribute.

1.3 Number
Attribute           value                Ro. example
Number              singular  (s)        fată
                    plural    (p)        fete

Multext-East / Concede  Morphosyntactic specifications V.2— Romanian  49

1.4 Case
There are five functional cases in Romanian (nominative, genitive,
dative, accusative and vocative) but they are expressed by at most
three syncretic forms for each number: 1. nominative-accusative;
2. genitive-dative; 3. vocative. The ambiguity of a syncretic case
form may be solved at the syntactic level. Taking into account this
syncretism we considered the following attribute-value combinations.
Attribute           value                Ro. example
Case                direct(d)            omul
                    oblique (o)          omului
                    vocative (v)         omule
At the word level, the distinction between the three syncretic forms
is possible only for the definite nouns. Masculine indefinite nouns
are not differentiated by case (except, sometimes for vocative. The
feminine indefinite nouns have different forms in singular for
nominative-accusative, genitive-dative and vocative, but the
genitive-dative singular forms are also indistinguishable from the
plural forms:
fată       - nominative-accusative, vocative singular, indefinite
fato       - vocative singular, indefinite
fete       - genitive-dative singular, nominative-accusative,
             genitive-dative, vocative plural, indefinite
băiat      - nominative-accusative, genitive-dative, vocative
             singular, indefinite
băiete     - vocative singular, indefinite
băieţi     - nominative-accusative, genitive-dative, vocative,
             plural, indefinite

1.5 Definiteness
In Romanian, nouns can be marked for definiteness with the enclitic
definite article.
Attribute           value                Ro. example
Definiteness        yes (y)              omul
                    no (n)               om

1.6 Clitic
See discussion in 2.10
Attribute           value                Ro. example
Clitic              no (n)               sora
                    yes (y)              soru-mea

Combinations
Here, and in all the sections to follow, the trailing hyphens are deleted.
=========================
Tag          Example
=========================
Ncmsrn       frate
Ncmson       frate
Ncmsvn       frate
Ncmsry       fratele
Ncmsoy       fratelui
Ncmprn       fraţi
Ncmpon       fraţi
Ncmpvn       fraţi
Ncmpry       fraţii
Ncmpoy       fraţilor
Ncmpvy       fraţilor
Ncfsrn       soră
Ncfsvn       soro
Ncfson       surori
Ncfprn       surori
Ncfpon       surori
Ncfpvn       surori
Ncfsoyy      soră(-sii)
Ncfpry       surorile
Ncfpoy       surorilor
Ncfpvy       surorilor
Ncfsryy      soră(-sa)
Ncmsrn       creion
Ncmson       creion
Ncmsryy      creionu-(i)
Ncmsoy       creionului
Ncfprn       creioane
Ncfpon       creioane
Ncfpry       creioanele
Ncfpoy       creioanelor
Npfsr        Ioana
Npfso        Ioanei
Npmsrn       Bucureşti
Npmsry       Bucureştiul

5.2 Verb (V)

The following attribute-value pairs are applicable to verbs in Romanian:

2.1 Type

Type                main (m)             a vedea
                    auxiliary (a)        a avea, a fi, a voi
                    modal (o)            a putea, a trebui
                    copulative (c)       a fi, a deveni

2.2 VForm
Traditionally, Romanian linguistics distinguishes between predicative
and non-predicative moods. This distinction may be easily mapped into
finite/non-finite dichotomy: indicative, subjunctive and imperative
are finite; infinitive, participle and gerund are non-finite (only
synthetic (non-compound) moods were mentioned; we use the opposition
synthetic-analytic to distinguish between concatenative (synthetic)
and compound (analytic) morpho-lexical phenomena).
Attribute           value                Ro. example
VForm               indicative(i)        vine
                    subjunctive (s)      vină
                    imperative (m)       vino
                    infinitive (n)       veni
                    participle (p)       venit
                    gerund (g)           venind
As only synthetic forms were considered, the values conditional and
presumptive for the VForm attribute were left out in Romanian. Another
value for VForm which was left out is the Supine. It appears mostly
with a preposition, except for a few intransitive verbs when they are
subordinated to the impersonal verb a trebui (must). Only the
preposition allows for differentiating a supine from a
participle-masculine-singular.

2.3 Tense
The synthetic values for the attribute Tense, as listed in the table
below, apply to the indicative mood. The value p (present) is used
also for subjunctive and infinitive. All the other moods either have
no tense or have compound form tenses.
Attribute           value                Ro. example
Tense               present (p)          văd
                    imperfect (i)        vedeam
                    past (s)             văzui
                    pluperfect (l)       văzusem

2.4 Person
Attribute           value                Ro. example
Person              first (1)            văd
                    second (2)           vezi
                    third (3)            vede

The following features are pertinent to those moods which permit an
adjectival use, i.e. participle and gerund. However, the adjectival
use of gerund is extremely rare (o mână tremurndă
- a shaking hand) and therefore gender and number apply mainly for the
participle.

2.5 Number
Attribute           value                Ro. example
Number              singular (s)         bătut
                    plural (p)           bătuţi

2.6 Gender
Attribute           value                Ro. example
Gender              masculine(m)         bătut
                    feminine(f)          bătută
                    neuter(n)            bătut-bătute

2.7, 2.8, 2.9 Voice, Negation, Definiteness
As these feature values are either realized as compound forms or
are irrelevant for Romanian they were omitted in our word-level encoding.

2.10 Clitic
The cliticization phenomenon in Romanian is not restricted to
verb-pronoun relationship, but may also be observed with the (main)
verb and the auxiliary, the noun or adjective with pronoun, with noun
or adjective with copula, pronoun with auxiliary, preposition with
(indefinite) article, numeral or (indefinite) pronoun, negative adverb
with verb, auxiliary or pronoun, and some others (mainly created
through the contracted forms of the verb "a fi"-to be). We restrict
ourselves to considering only the graphically marked clicitizations.
In such cases, the two, three or (sometimes) four constituents of a
cliticized word-form are always separated by a hyphen. Omitting the
hyphen in such cases is an unacceptable error in written Romanian.
The examples below illustrate the specific types of cliticization we
will consider:
dă-mi-l !                 (you give +to me+it/him = give it/him to me)
lăsatu-ne-ai              (left + us+ you have = you have left us)
sparge-ţi-s-ar lampa      (break + your + s-part-conj + has lamp =
                           I wish your lamp would break)
soră-mii                  (sister + my = to my sister)
faţa-i                    (his/her face | the face is)
roşu-i                    (his/her red | the red is )
m-au                      (me + they have )

într-o gaură              (in + a hole)
într-o oră                (in + one hour)
într-unele                (into + some of)
n-aud                     (not  + I hear = I do not hear)
n-am                      (not + I/we have = I have not )
nu-mi                     (not + to me )
The order in which different constituents of a cliticized form appear
is governed by precise morphological rules. For instance, the
auxiliaries always appear in the last position. The main verbs, except
those beginning either with the letter "a" or the letter î and
the contracted forms (-s and -i) of the verb "a fi"- to be, always
appear in the first position, nouns and adjectives always precede the
cliticized pronouns, the negative adverbial particles nu- and n- only
appear in the first position, and so on.  However, in order to reduce
spurious ambiguity in morpho-lexical encoding, we considered the
attribute CLITIC as relevant only in those cases where the
clicitization phenomenon resulted in graphemic modification of the
cliticized term (as it is the epenthetic u in the gerund form shown in
the table below).
Attribute           value                Ro. example
Clitic              no (n)               am ridicat
                    yes (y)              ridicându-l

Combinations
========================
Tag           Example
========================
Vmii1s        abandonam
Vmii2s        abandonai
Vmii3s        abandona
Vmii1p        abandonam
Vmii2p        abandonaţi
Vmii3p        abandonau
Vmis1s        abandonai
Vmis2s        abandonaşi
Vmis3s        abandonă
Vmis1p        abandonarăm
Vmis2p        abandonarăţi
Vmis3p        abandonă
Vmil1s        abandonasem
Vmil2s        abandonaseşi
Vmil3s        abandonase
Vmil1p        abandonaserăm
Vmil2p        abandonaserăţi
Vmil3p        abandonaseră
Vmip1s        abandonez
Vmsp1s        abandonez
Vmip2s        abandonezi
Vmsp2s        abandonezi
Vmip3s        abandonează

Vmip3p        abandonează
Vmsp3s        abandoneze
Vmsp3p        abandoneze
Vmsp1p        abandonăm
Vmsp2p        abandonaţi
Vmm-2s        abandonează
Vmm-2p        abandonaţi
Vmnp          abandona
Vmp--sm       abandonat
Vmp--sm---y   abandonatu
Vmp--sf       abandonată
Vmp--pf       abandonate
Vmp--pm       abandonaţi
Vmg           abandonând
Vmg-------y   abandonându
Va--1s        as,
Voip          trebuie
Vcip1s        sunt
Vcip3p        sunt
Vcip1s----y   -s
Vcip3p----y   -s

5.3 Adjective (A)

3.1 Type
Although it is not common practice in Romanian linguistics, one could
make the distinction between qualificative and determinative
adjectives. However, the attribute-value pairs proposed in this
section are appropriate for the qualificative adjectives.
Attribute           value                Ro. example
Type                qualificative (f)    frumos

3.2 Degree
The default value is positive; adjectives have also comparative and
superlative degrees, but in most of the cases they are expressed by
means of analytical forms (e.g. comp. mai bun (better), superl. cel
mai bun (the best)). A few adjectives have intrinsic etymological
comparative or superlative meanings (e.g. comparatives: anterior,
major; superlatives: optim, maxim, extrem etc.). The prefixes super-,
extra-, ultra- etc., are quite productive in forming the
quasi-analytic1 superlatives.  Adjectives are characterized by gender,
number and case.
Attribute           value                Ro. example
Degree              positive (p)         frumos

                    comparative (c)      ulterior
                    superlative (s)      extrem

3.3 Gender
Attribute           value                Ro. example
Gender              masculine (m)        bun
                    feminine (f)         bună
                    neuter (n)           sg.bun/pl.bune

3.4 Number
Attribute           value                Ro. example
Number              singular (s)         bun
                    plural (p)           buni

3.5 Case
The adjectives present the same case syncretism as the nouns, except
for few adjectives that have an additional special form for
Genitive-Dative cases in plural (e.g. G.D.pl.  multor).
Attribute           value                Ro. example
Case                direct(r)            bunul
                    oblique (o)          bunului
                    vocative (v)         bunule

3.6 Definiteness
In noun-adjective construction, the definite article may attach
enclitically to either adjectives or modified nouns (never to both of
them). If present, the definite article attaches to the right of the
first word in the sequence.
Bunul  om  (The kind man)
Omul  bun. (The kind man)
Attribute           value                Ro. example
Definiteness        yes (y)              bunul
                    no (n)               bun

3.7 Clitic
See discussion on clitics in 2.10

Combinations

====================
Tag            Example
====================
Afpmsrn        bun
Afpmson        bun
Afpmsvn        bun
Afpmprn        buni
Afpmpon        buni
Afpmpvn        buni
Afpmsry        bunul
Afpmsoy        bunului
Afpmpry        bunii
Afpmpoy        bunilor
Afpfsrn        bună
Afpfsvn        bună
Afpfson        bune
Afpfprn        bune
Afpfpon        bune
Afpfpvn        bune
Afpfsry        buna
Afpfsoy        bunei
Afpfpry        bunele
Afpfpoy        bunelor
Afcmsrn        ulterior
Afcmson        ulterior
Afsmsrn        extrem
Afp            gri

5.4 Pronoun (P)

4.1 Type
In Romanian it is worth differentiating the negative pronoun from
other indefinite pronouns: a negative pronoun cannot be an argument
for a verb unless the verb itself is negated too (e.g. Nu am
văzut pe nimeni / *Am văzut pe nimeni).
Attribute           value                Ro. example
Type                demonstrative(d)     acesta
                    indefinite (i)       oricine
                    possesive(s)         (al) meu
                    int_rel (w)          ce
                    personal(p)          eu
                    reflexive(x)         se
                    negative (z)         nimeni
                    emphatic (h)         însumi

4.2 Person

Attribute           value                Ro. example
Person              first (1)            eu
                    second (2)           tu
                    third (3)            el

4.3 Gender
Attribute           value                Ro. example
Gender              masculine (m)        el
                    feminine (f)         ea
                    neuter (n)           sg.acesta/pl.acestea

4.4 Number
Attribute           value                Ro. example
Number              singular (s)         eu
                    plural (p)           noi

4.5 Case
For the second person of the personal pronoun in both singular and
plural there is the vocative case too.The direct and oblique values
are needed for the syncretic causal forms of the other pronouns than
the personal ones.
Attribute           value                Ro. example
Case                nominative (n)       el
                    genitive (g)         (al) lui
                    dative (d)           lui
                    accusative (a)       (pe) el
                    vocative (v)         tu, voi!
                    direct(r)            acesta
                    oblique (o)          acestuia

4.6 Owner_Number
This attribute is meaningful for the Possessive pronouns and refers to
the grammatical number of the possessor.
Attribute           value                Ro. example
Owner_Number        singular (s)         meu
                    plural (p)           nostru
The gramatical number of the possesed object(s) is expressed by the
attribute Number (described above).

4.7 Owner_Gender

This attribute is irrelevant for Romanian.

4.8 Clitic
See discussion on clitics in 2.10.

4.9, 4.10, 4.11, 4.12, 4.13 Referent_Type, Syntactic_Type,
Definiteness, Animate, Clitic_s
These attributes are irrelevant for Romanian.

4.14 Pronoun_Form
For Romanian we need an attribute (called Pronoun_Form) to make the
distinction between strong and weak forms of the same pronoun. All the
weak forms can be adjoined to the adjacent words both proclitically or
enclitically. In such cases the junction is always graphically marked
by a hyphen between the pronoun and the neighboring word. The hyphen
also marks possible elisions from either pronoun or the adjacent word.
Although in traditional grammar books the demonstrative, int_rel and
indefinite pronouns are not characterised by person, in our
dictionaries they are recorded (for reasons beyond morpho-lexical
encoding) as 3rd person (the same as nouns). However, for the
automatic tagging this value has been marked as irrelevant.
Attribute           value                Ro. example
Pronoun_Form        strong (s)           lui
                    weak (w)             îi, i-

Combinations
====================================
Tag                    Example
====================================
Pp1msn--------s        eu
Pp1msd--------w        mi
Pp1msd--------s        mie
Pp1msd--y-----s        mi-
Pd-msr                 acesta
Pd-mso                 acestuia
Pi-mpr                 toţi
Ps1fsrs                mea
Pw-mso                 cărui
Pn-msr                 nimeni
Ph1msr                 însumi
Ph1fsr                 însămi
Px3msa--------s        sine
Px3msa--------w        se
Px3msa--y-----w        s-

5.5 Determiner (D)

5.1 Type
The need for a negative value of the determiners' Type attribute is
argued on the same lines as in the section on pronoun' s Type.  In
Romanian the negative determiner is expressed by the unit nici +
indefinite article (e.g. nici un, nici o).  In Romanian, there are
specific forms for the so-called emphatic determiner, which may
accompany both a noun and a personal pronoun: fata
însăşi (the girl herself), also ea
însăşi (she herself).
Attribute           value                Ro. example
Type                demonstrative (d)    acest
                    indefinite (i)       orice
                    possessive (s)       meu
                    int_rel (w)          ce
                    negative (z)         nici un
                    emphatic (h)         însuşi

5.2 Person
This attribute is meaningful for the Possessive determiners and refers
to the grammatical person of the possessor.
Attribute           value                Ro. example
Person              first (1)            meu
                    second (2)           tău
                    third (3)            său

5.3 Gender
Attribute           value                Ro. example
Gender              masculine (m)        meu
                    feminine (f)         mea
                    neuter (n)           sg.meu/pl.mele

5.4 Number
Attribute           value                Ro. example
Number              singular (s)         meu
                    plural (p)           mei

5.5 Case

Case                direct(d)            aceasta
                    oblique (o)          acestei

5.6 Owner_Number
This attribute is meaningful for the Possessive determiners and refers
to the grammatical number of the possessor.
Attribute           value                Ro. example
Owner_Number       singular (s)          meu
                   plural (p)            nostru

5.7 Owner_Gender
This attribute is irrelevant for Romanian.

5.8 Clitic
See discussion on clitics in 2.10.
Attribute           value                Ro. example
Clitic              no (n)               mama mea
                    yes (y)              maică- mea

5.9 Modific_Type
As mentioned in the corresponding section on Pronoun, the Modific_Type
attribute is relevant for some determiners too. The prenominal
determiner always precedes the noun (e.g.acest băiat - this
boy), whereas the postnominal determiner appears only after the noun
(e.g. băiatul acesta - this boy).
Attribute           value                Ro. example
Modific_Type        prenominal (e)       acest
                    postnominal (o)      acesta

Combinations
=======================
Tag           Example
=======================
Dd-mso---e    acestui
Dd-mso---o    acestuia
Di-mpr        toţi
Ds1fsrs       mea
Dw-msr        care
Dw-mso        cărui
Dz-msr        nici_un
Dh1msr        însumi

Dh1fsr        însămi
Although in traditional grammar books the demonstrative, indefinite
and int_rel determiners are not characterised by person, in our
dictionaries they are recorded (for reasons beyond morpho-lexical
encoding) as 3rd person (the same as nouns). However, for the
automatic tagging this value has been marked as irrelevant.

5.6 Article (T)

6.1 Type
Although it presents only a few items, the article in Romanian has
four types, unlike in most of the European languages. Beside the two
recommended types: definite and indefinite which have the generally
known semantic value, Romanian uses two additional types of articles,
which are semantically subordinated to the definite article but which
have special forms and meanings:
- the possessive article (also called genitival article) is an element
  in the structure of the possessive pronoun, of the ordinal numeral
  (e.g. al meu (mine) and al treilea (the third)), and of the indefinite
  genitive forms of the nouns (e.g. capitol al cărţii
  (chapter of the book)).
- the demonstrative article links a definite noun to its determinants,
  links a numeral or an adjective to a noun, and it is a constituent
  part of the relative superlative (e.g.  fata cea mare (the elder
  girl), cel lenes, (the lazy), respectively prietenul cel mai bun
  (the best friend)).
Notice that the definite article has only enclitic forms, except for
one proclitical form (lui + proper noun: lui Ion).
Attribute           value                Ro. example
Type                definite (f)         lui
                    indefinite (i)       un
                    possessive (s)       al
                    demonstrative (d)    cel

6.2 Gender
Attribute           value                Ro. example
Gender              masculine (m)        un
                    feminine (f)         o
                    neuter (n)           sg.cel/pl.cele

6.3 Number
Attribute           value                Ro. example
Number              singular (s)         un
                    plural (p)           nişte

6.4 Case
Attribute           value                Ro. example
Case                direct(r)            cel
                    oblique (o)          celui

6.5 Clitic
The inflected forms of the foreign-origin words (mainly nouns) not
fully assimilated, are usually written with a hyphen between the
base-form and the inflectional ending.  In our encoding, we classified
these endings (which are supposed to be split by the segmenter) as
clitic articles (clitic attribute is always "y") which can be either
definite (type=f, "-istul") or indefinite (type=i, "ist") and are
characterised by gender (gender=m, "ist"; gender=f, "istă"),
number (number=s, "ist"; number=p, "işti") and case (case=r,
"istul"; case=o, "istului").

Combinations
=====================
Tag          Example
=====================
Tfmso        lui
Tffso        lui
Timsr        un
Tsmpr        ai
Tdfso        celei
Timsry       -ist
Timsoy       -ist
Tfmsry       -istul
Tfmsoy       -istului

5.7 Adverb (R)

7.1 Type
The distinction proposed here considers the principal syntactic
properties of the adverbs.  For Romanian, the general type includes
most of the pronominal adverbs (demonstrative: aici (here),

indefinite: oriunde (anywhere)). As argued before for pronouns and
determiners, a distinct negative value is needed for adverbs as well
(nicăieri - nowhere, niciodată - never). The particle
type covers those adverbs which can dislocate verbal compound forms
(ex. Ea a tot cântat -- She has ever sung) or mark degrees
(ex. circa (about), foarte (very), prea (too)). Such adverbs are cam,
mai, prea, şi, tot, foarte etc.  A useful distinction in
Romanian considers the adverbs which can have predicative role, that
is they can govern a subordinate sentence (ex. Fireşte
că o ştiu -- Certainly I know it). Here (for uniformity
within a multilingual environment), they are squeezed into the
modifier class.  No formal distinction is made between the
interrogative adverbs and the relative ones.
Attribute           value                Ro. example
Type                general (g)          bine, acolo
                    particle (p)         mai, cam
                    negative (z)         nicăieri
                    modifier (m)         fireşte, poate
                    int_rel (w)          cum
                    portmanteau (c)      ca
The "portmanteau" type of adverb was introduced to cover some few
words which can be both adverbs and conjunctions (with adverbial
reading more frequent). This was necessary for tagging purposes.
7.2 Degree
In Romanian, the comparative and superlative of adverbs is formed
analytically with mai (puţin), cel mai (puţin), foarte:
ex. mai repede (faster), cel mai devreme (at the
earliest). Nonetheless there are some adverbs with comparative or
superlative meaning (ex. optim, ulterior,. definitiv). These adverbs
can be used for expressing the absolute superlative of other adverbs
or adjectives: ex. extrem de bine, formidabil de frumos.
Attribute           value                Ro. example
Degree              positive (p)         bine
                    comparative (c)      ulterior
                    superlative (s)      extrem

7.3 Clitic
See discussion on clitics in 2.10.

Combinations
=====================
Tag        Example
=====================
Rgp        repede
Rgs        extraordinar

Rgc        ulterior
Rp         mai
Rz         nicăieri
Rm         probabil
Rw         cum

5.8 Adposition (S)

8.1 Type
It is the preposition type which is only pertinent to Romanian,
although some intercalating adpositions may be seen as a sort of
circumposition, for instance între...şi...
(between..and...).
Attribute           value                Ro. example
Type                preposition (p)      la, pe, în

8.2 Formation
In Romanian there is a distinct class of compound prepositions. Each
of them forms a formal and semantic unit, although graphically they
stay unfused, e.g. de la, pe la, de pe, etc.
Attribute           value                Ro. example
Formation           simple (s)           la, pe, în
                    compound (c)         de la

8.3 Case
This attribute marks the subcategorisation properties of the adpositions.
Attribute           value                Ro. example
Case                genitive (g)         înaintea
                    dative (d)           datorită
                    accusative (a)       la

8.4 Clitic
See discussion on clitics in 2.10.

Combinations
======================
Tag         Example
======================

Spsa        în
Spsay       într-
Spsd        datorită
Spca        de_la

5.9 Conjunction (C)

9.1 Type
The distinction between coordinating and subordinating conjunctions is
pertinent to Romanian as well.
Attribute           value                Ro. example
Type                coordinating (c)     dar
                    subordinating (s)    că, dacă
                    portmanteau (r)        şi
The "portmanteau" type of conjunction applies only to the word "şi"
which can be both a coordonating conjunction and an adverb. The
distinctionamong these interpretations is rather tricky for the
average native speaker and was a constant source of noise in automatic
tagging. Therefore, for the sake of automatic processing we defined this
"portmanteau" type value.
9.2 Formation
Likewise prepositions, we can distinguish  two kinds of conjunctions in
Romanian:
- simple conjunctions: e.g. şi,dar,deşi  etc.
- conjunctions formed periphrastically, with some word/phrase combined
  by a conjunction: din moment ce, fără să,
  fat,ă de cum etc.
Attribute           value                Ro. example
Formation           simple (s)           deoarece
                    compound (c)         de_vreme_ce

9.3 Coord_Type
In Romanian, there are three kinds of conjunctions depending on their
usage: as such or together with other conjunctions or adverbs:
- simple, between conjuncts: Ion ori Maria (John or Mary);
- repetitive, before each conjunct: fie  Ion fie  Maria fie...
  (either John or Mary or...)

- correlative, before a conjoined phrase, it requires specific
  coordinators between conjuncts: atât mama cât şi
  tata (both mother and father).
With respect to the place of the conjunctions, most of them stay
before the conjunct, except for: aşadar, deci (so), dar,
însă (but), dacă (if), which also appear with
expressive value inside the conjoined sentence.
Attribute           value                Ro. example
Coord-Type          simple (s)           dar,deoarece
                    repetit (r)          fie...fie...
                    correlat (c)         atît...cât şi

9.4 Sub_Type
In Romanian, each conjunction requires another mood, so that the
diversity may be controlled by subcategorisation rules.
Attribute           value                Ro. example
Sub-Type            negative (z)         nici
                    positive (p)         dar
This attribute distinguishes among the positive and negative
conjunctions, providing means to control verbal double negation, (as
in case of the negative pronouns, determiners and adverbs): nici NU am
venit, nimeni NU vorbeşte, nici_un tren N-a trecut,
nicăieri N-am văzut

9.5 Clitic
See discussion on clitics in 2.10.
Attribute           value                Ro. example
Clitic              no (n)               că aşa
                    yes (y)              c-aşa

Combinations
======================
Tag          Example
======================
Crssp        şi
Ccssp        dar
Ccrsp        fie...fie
Csrsz        nici...nici
Csscp        de_vreme_ce

5.10 Numeral (M)

10.1 Type
Traditional Romanian grammars usually distinguish seven numeral types,
where five of them have specific forms and the other two are obtained
by composition.
The first group is made up by the following numeral types: cardinal
(trei-three), ordinal (al treilea-the third), fractional (treime-one
third), multiple (întreit-trine), collective
(amândoi-both).
The second group contains the numeral types which are composed by
means of other parts of speech: distributive (câte trei-...each
three...), adverbial (de trei ori-thrice) and again the collective
numeral which also has compound forms (toţi trei-all three).
Nonetheless, as the numerals of the second group have a weak syntactic
cohesion, namely each composition element may be regarded as an
element of the sentence, with its own grammatical function, these last
numeral types are irrelevant for the morphosyntactic annotation.
Attribute           value                Ro. example
Type                cardinal (c)         trei
                    ordinal (o)          (al) treilea
                    fractal (f)          treime
                    multiple (m)         întreit
                    collect(l)           tustrei
In Romanian (as in many other languages) several numerals have noun
behaviour (some grammarians classify such numerals as nouns) with
gender and declension of their own, which they preserve even in the
composition of the superior order numerals; these are, for instance,
sută (hundred), mie (thousand), milion (million) and miliard
(billion). In a sentence most numerals may fulfill the function of
other parts of speech like noun, determiner or adverb.

10.2 Gender
Attribute           value                Ro. example
Gender              masculine (m)        doi, primul
                    feminine  (f)        două, prima
                    neuter (n)           (un) milion, (două) milioane

10.3 Number
Attribute           value                Ro. example
Number              singular (s)         primul
                    plural (p)           primii

10.4 Case
Attribute           value                Ro. example
Case                direct(r)            primul
                    oblique (o)          primului

10.5 Form
Attribute           value                Ro. example
Form                digit (d)            1690
                    letter (l)           unsprezece
                    both (b)             5 mii
                    roman (r)            XIV

10.6 Definiteness
By virtue of their noun or adjective value, some numerals may take the
enclitic article (prim/primul - first/the first). Consequently for the
Romanian, definiteness attribute helps distinguish the enclitic forms
from the other forms.
Attribute           value                Ro. example
Definiteness        yes (y)              primul
                    no (n)               prim

10.7 Clitic
See discussion on clitics in 2.10.

Combinations
=======================
Tag           Example
=======================
Mcmprl        doi
Mcmpol        doi
Momsrl        doilea
Momsol        doilea
Mlmpr         amândoi
Momsrlyy      primu-i
Mffpoly       treimilor

5.11 Interjections (I)

In Romanian there are no relevant  subcategories of interjections.

====================
Tag        Example
====================
I        oh,ah,au
====================

5.12 Residual (X)

No attributes are defined for this category.
==========================
Tag        Example
==========================
X        show, a+b, retro-
==========================

5.13 Abbreviation (Y)

The Syntactic_Type attribute is useful for specifying the grammatical
category of an abbreviation. Although the values for this attribute
could range over the part of speech categories in the language, in
Romanian most of the abbreviations falls into noun class.

13.1 Syntactic_Type
Attribute           value                Ro. example
Syntactic_Type      nominal(n)           d-na (doamna)
                    verbal (v)           v. (vezi)
                    adjectival (a)       ant. (anterior)
                    adverbial(r)         f. (foarte)
                    pronominal(p)        d-ei (dumneaei)

13.2 Gender
Attribute           value                Ro. example
Gender              masculine (m)        d-lui
                    feminine (f)         d-na
                    neuter(n)            apt.

13.3 Number
Attribute           value                Ro. example
Number              singular (s)         d-na
                    plural (p)           d-nele

13.4 Case
Attribute           value                Ro. example
Case                direct (r)           d-na
                    oblique (o)          d-nei

13.5 Definiteness
Attribute           value                Ro. example
Definiteness        yes (y)              d-nele
                    no (n)               d-ne

Combinations
======================
Tag           Example
======================
Ynmsry        d-ul
Ynfsoy        d-nei
Ynnsry        apt.

5.14 Particle (Q)

14.1 Type
Attribute           value                Ro. example
Type                negation(z)          nu,n-
                    infinitive (n)       a
                    subjunctive (s)      să
                    aspect               fi
                    future               o

14.2 Formation
This attribute is not relevant for Romanian.

14.3 Clitic
See  discussion on clitics in 2.10.
Attribute           value                Ro. example
Clitic              yes (y)              n-am
                    no (n)               nu am

Combinations
=================
Tag       Example
=================
Qz        nu
Qz-y      n-
Qn        a
Qs        să
Qa        fi
Qf        o

Application to Slovene

Authors: Tomaž Erjavec, Peter Holozan, Vojko Gorjanc, Marko Stabej

Acknowledgements: The authors thank Velimir Gjurin, France Žagar, Vladimír Petkevič, Lydia Sinapova and David Stermole for their much appreciated comments and suggestions. All errors of course remain our own.

6.1 Introduction

The first version of these specifications was produced in the scope of the MULTEXT-East project. The tables and lexicon are described in T. Erjavec: The MULTEXT-East Slovene Lexicon. ERK’98, ⟨http://nl.ijs.si/et/Bib/ERK98/erk/ ⟩.

The second version of the guidelines was produced for the FIDA Slovene reference corpus, ⟨http://www.fida.net/ ⟩. Here the specifications were revised and localised to Slovene. Finally, the harmonisation of the FIDA version with the common Multext-East one was accomplished in the context of the Concede project.

The rest of this report is structured as follows: we first give the table of categories and then a section for all the Parts of Speech distinguished in Slovene.

For each PoS we first give the table of feature values, based on the common MULTEXT-East tables. In addition to giving the English terms each attribute-value table gives also the Slovene names of the attributes, their values and their MSD codes. This localised specification enables not only automatic expansion of e.g. the MSD Ncmsa–n to Noun common masculine singular

accusative -animate but also the translation into Somet–n, meaning Samostalnik občno_ime
moški ednina tožilnik -živost.

Each PoS section contains, in addition to the attribute-value tables, also a section with ’feature co-occurrence restrictions’ for the allowed combinations and a listing with enumerated allowed (lexical) MSDs. The listing accompanies each MSD with an example word-form and lemma.

Because the Slovene tables are based on the common MULTEXT-East ones, new attributes introduced in FIDA had to be added at the tail. This leads in some cases to ungainly MSDs; for example, the new Verb attribute of Aspect has position 14, leading to MSDs such as Von–––––-e. In the future, it might be worth investigating a ’minimum length method of localisation’ where attributes are allowed to be locally permuted, or change the original MULTEXT tables, to allow East languages to place lexical attributes towards the beginning of the MSDs.

6.2 Table of Categories

=============== ======= =============== =======
PoS-en          Code-en PoS-sl          Code-sl
=============== ======= =============== =======
Noun            N       Samostalnik     S
Verb            V       Glagol          G
Adjective       A       Pridevnik       P
Adverb          R       Prislov         R
Pronoun         P       Zaimek          Z
Adposition      S       Predlog         D
Conjunction     C       Veznik          V
Particle        Q       Členek          L
Interjection    I       Medmet          M
Numeral         M       Števnik         K
Abbreviation    Y       Okrajšava       O
Residual        X       Neuvrščeno      N
Article         -       Člen            -
Determiner      -                       -
=============== ======= =============== =======

Notes:

• The order of the categories above is different from the one in the common tables for MULTEXT-East languages.

• The category Article (Slovene: ’člen’) exists for Slovene, but it encompasses only two words (’ta’ / this, ’en’ / one) and even these are used only colloquially. Furthermore, these two words can also function as a pronoun or numeral respectively. As it would be difficult to disambiguate between the categories, Article is not used for Slovene.

6.3 Noun (N) // Samostalnik (S)

= ============== ============== ==== ============== ============== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl         C-sl
= ============== ============== ==== ============== ============== ====
1 Type           common         c    vrsta          občno_ime      o
                 proper         p                   lastno_ime     l
- -------------- -------------- ---- -------------- -------------- ----
2 Gender         masculine      m    spol           moški          m
                 feminine       f                   ženski         z
                 neuter         n                   srednji        s
- -------------- -------------- ---- -------------- -------------- ----
3 Number         singular       s    število        ednina         e
                 plural         p                   množina        m
                 dual           d                   dvojina        d
- -------------- -------------- ---- -------------- -------------- ----
4 Case           nominative     n    sklon          imenovalnik    i
                 genitive       g                   rodilnik       r
                 dative         d                   dajalnik       d
                 accusative     a                   tožilnik       t
                 locative       l                   mestnik        m
                 instrumental   i                   orodnik        o
* ***************************** **** ************** ************** ****
5 Definiteness                  -                                 -
- -------------- -------------- ---- -------------- -------------- ----
6 Clitic                        -                                 -
- -------------- -------------- ---- -------------- -------------- ----
7 Animate        no             n    živost         ne             n
                 yes            y                   da             d
- -------------- -------------- ---- -------------- -------------- ----
8 Owner_Number                  -
- -------------- -------------- -
9 Owner_Person                  -
---------------- -------------- -
10Owned_Number                  -
=================================

Notes:

• The ’deverbal noun’ or ’gerund’ is marked as a Noun, e.g. from the Verb ’iskati’ we get ’iskanje’, Ncnsn.

• Animacy is marked only on the Noun masculine singular accusative (N*msa).

• Feminine declension masculine nouns have two Ncmsa forms, both are marked as animate, e.g. vodjo / vodja are both Ncmsa–y.

Combinations

*** **** **** **** **** ---- ---- ---- ========================================
PoS Type Gend Numb Case(Def)(Cltc)Anim Examples
*** **** **** **** **** ---- ---- ---- ========================================
N   p   [mfn]  s    n                 Janez,      Micka,    Kršo
N   p   [mfn]  s    g                 Janeza,     Micke,    Krša
N   p   [mfn]  s    d                 Janezu,     Micki,    Kršu
N   p    m     s    a    -    -    n  Kranj
N   p    m     s    a    -    -    y  Janeza
N   p   [fn]   s    a                             Micko,    Kršo
N   p   [mfn]  s    l                 Janezu,     Micki,    Kršu
N   p   [mfn]  s    i                 Janezom,    Micko,    Kršom
N   p   [mf]   p    n                 Slovenci,   Slovenke
N   p   [mf]   p    g                 Slovencov,  Slovenk
N   p   [mf]   p    d                 Slovencom,  Slovenkam
N   p   [mf]   p    a                 Slovenci,   Slovenke
N   p   [mf]   p    l                 Slovencih,  Slovenkah
N   p   [mf]   p    i                 Slovenci,   Slovenkami
N   p   [mf]   d    n                 Slovenca,   Slovenki
N   p   [mf]   d    g                 Slovencov,  Slovenk
N   p   [mf]   d    d                 Slovencoma, Slovenkama
N   p   [mf]   d    a                 Slovenca,   Slovenki
N   p   [mf]   d    l                 Slovencih,  Slovenkah
N   p   [mf]   d    i                 Slovenci,   Slovenkama
N   c   [mfn]  s    n                 stol,    lipa,   sonce
N   c   [mfn]  s    g                 stola,   lipe,   sonca
N   c   [mfn]  s    d                 stolu,   lipi,   soncu
N   c    m     s    a    -    -    n  stol
N   c    m     s    a    -    -    y  gospoda
N   c   [fn]   s    a                          lipo,   sonce
N   c   [mfn]  s    l                 stolu,   lipi,   soncu
N   c   [mfn]  s    i                 stolom,  lipo,   soncem
N   c   [mfn]  p    n                 stoli,   lipe,   sonca
N   c   [mfn]  p    g                 stolov,  lip,    sonc
N   c   [mfn]  p    d                 stolom,  lipam,  soncem
N   c   [mfn]  p    a                 stole,   lipe,   sonca
N   c   [mfn]  p    l                 stolih,  lipah,  soncih
N   c   [mfn]  p    i                 stoli,   lipami, sonci
N   c   [mfn]  d    n                 stola,   lipi,   sonci
N   c   [mfn]  d    g                 stolov,  lip,    sonc
N   c   [mfn]  d    d                 stoloma, lipama, soncema
N   c   [mfn]  d    a                 stola,   lipi,   sonci
N   c   [mfn]  d    l                 stolih,  lipah,  soncih
N   c   [mfn]  d    i                 stoloma, lipama, soncem
*** **** **** **** **** ---- ---- ---- ========================================

Multext-East / FIDA / Concede  Morphosyntactic specifications V.2— Slovene  76

Lexicon

98 MSDs:

Ncfda(čački / čačka) Ncfdd(čačkama / čačka) Ncfdg(čačk / čačka) Ncfdi(čačkama / čačka) Ncfdl(čačkah / čačka) Ncfdn(čački / čačka) Ncfpa(čačke / čačka) Ncfpd(čačkam / čačka) Ncfpg(čačk / čačka) Ncfpi(čačkami / čačka) Ncfpl(čačkah / čačka) Ncfpn(čačke / čačka) Ncfsa(čačko / čačka) Ncfsd(čački / čačka) Ncfsg(čačke / čačka) Ncfsi(čačko / čačka) Ncfsl(čački / čačka) Ncfsn(čačka / =) Ncmda(čaja / čaj) Ncmdd(čajema / čaj) Ncmdg(čajev / čaj) Ncmdi(čajema / čaj) Ncmdl(čajih / čaj) Ncmdn(čaja / čaj) Ncmpa(čaje / čaj) Ncmpd(čajem / čaj) Ncmpg(čajev / čaj) Ncmpi(čaji / čaj) Ncmpl(čajih / čaj) Ncmpn(čaji / čaj) Ncmsa– n(čaj / čaj) Ncmsa–y(čarodeja / čarodej) Ncmsd(čajniku / čajnik) Ncmsg(čaja / čaj) Ncmsi(čajem / čaj) Ncmsl(čajniku / čajnik) Ncmsn(čaj / =) Ncnda(čaščenji / čaščenje) Ncndd(čaščenjema / čaščenje) Ncndg(čaščenj / čaščenje) Ncndi(čaščenjema / čaščenje) Ncndl(čaščenjih / čaščenje) Ncndn(čaščenji / čaščenje) Ncnpa(čaščenja / čaščenje) Ncnpd(čaščenjem / čaščenje) Ncnpg(čaščenj / čaščenje) Ncnpi(čaščenji / čaščenje) Ncnpl(čaščenjih / čaščenje) Ncnpn(čaščenja / čaščenje) Ncnsa(čaščenje / čaščenje) Ncnsd(čaščenju / čaščenje) Ncnsg(čaščenja / čaščenje) Ncnsi(čaščenjem / čaščenje) Ncnsl(čaščenju / čaščenje) Ncnsn(čaščenje / =) Npfda(Čehinji / Čehinja) Npfdd(Čehinjama / Čehinja) Npfdg(Čehinj / Čehinja) Npfdi(Čehinjama / Čehinja) Npfdl(Čehinjah / Čehinja) Npfdn(Čehinji / Čehinja) Npfpa(Čehinje / Čehinja) Npfpd(Čehinjam / Čehinja) Npfpg(Čehinj / Čehinja) Npfpi(Čehinjami / Čehinja) Npfpl(Čehinjah / Čehinja) Npfpn(Čehinje / Čehinja) Npfsa(Češko / Češka) Npfsd(Češki / Češka) Npfsg(Češke / Češka) Npfsi(Češko / Češka) Npfsl(Češki / Češka) Npfsn(Češka / =) Npmda(Čajkovska / Čajkovski) Npmdd(Čajkovskima / Čajkovski) Npmdg(Čajkovskih / Čajkovski) Npmdi(Čajkovskima / Čajkovski) Npmdl(Čajkovskih / Čajkovski) Npmdn(Čajkovska / Čajkovski) Npmpa(Čajkovske / Čajkovski) Npmpd(Čajkovskim / Čajkovski) Npmpg(Čajkovskih / Čajkovski) Npmpi(Čajkovskimi / Čajkovski) Npmpl(Čajkovskih / Čajkovski) Npmpn(Čajkovski / =) Npmsa–n(Črnomelj / Črnomelj) Npmsa–y(Čajkovskega / Čajkovski) Npmsd(Čajkovskemu / Čajkovski) Npmsg(Čajkovskega / Čajkovski) Npmsi(Čajkovskim / Čajkovski) Npmsl(Čajkovskem / Čajkovski) Npmsn(Čajkovski / =) Npnsa(Škotsko / Škotsko) Npnsd(Škotskemu / Škotsko) Npnsg(Škotskega / Škotsko) Npnsi(Škotskim / Škotsko) Npnsl(Škotskem / Škotsko) Npnsn(Škotsko / =)

6.4 Verb (V) // Glagol (G)

= ============== ============== ==== ============== ============== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl         C-sl
= ============== ============== ==== ============== ============== ====
1 Type           main           m    vrsta          polnopomenski  p
                 modal          o                   naklonski      n
                 copula         c                   vezni          v
- -------------- -------------- ---- -------------- -------------- ----
2 VForm          indicative     i    gl. oblika     povednik       p
                 imperative     m                   velelnik       v
                 conditional    c                   pogojnik       g
                 infinitive     n                   nedoločnik     n
                 participle     p                   deležnik       d
                 supine         u                   namenilnik     m
- -------------- -------------- ---- -------------- -------------- ----
3 Tense          present        p    čas            sedanjik       s
                 future         f                   prihodnjik     p
                 past           s                   nesedanjik     r
- -------------- -------------- ---- -------------- -------------- ----
4 Person         first          1    oseba          prva           1
                 second         2                   druga          2
                 third          3                   tretja         3
- -------------- -------------- ---- -------------- -------------- ----
5 Number         singular       s    število        ednina         e
                 plural         p                   množina        m
                 dual           d                   dvojina        d
- -------------- -------------- ---- -------------- -------------- ----
6 Gender         masculine      m    spol           moški          m
                 feminine       f                   ženski         z
                 neuter         n                   srednji        s
******************************* **** ************** ************** ****
7 Voice          active         a    način          tvornik        t
                 passive        p                   trpni_deležnik r
- -------------- -------------- ---- -------------- -------------- ----
8 Negative       no             n    nikalnost      nezanikani     n
                 yes            y                   zanikani       d
- -------------- -------------- ---- -------------- -------------- ----
9 Definiteness                  -
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -

13Clitic_s                      -
- -------------- -------------- ---- -------------- -------------- ----
14Aspect  !NEW!  progressive    p    vid            nedovršni      n
                 perfective     e                   dovršni        d
= ============================= ==== ============== ============== ====

Notes:

• The verb ’to be’ / ’biti’ in all its functions is Type=c;

• The "past participle" is used for making all the compound active tenses (future, past, pulperfect) and is encoded analytically as Type=participle, Tense=past, Voice=Active.

• The passive participles is encoded analytically as Type=participle, Tense=’-’, Voice=Passive. This encoding will only be used for the passive participle in the predicative position, e.g. ’on je bil tepen’ / he was beaten. The case marked passive participles in the attributive position (e.g. ’tepen pes’ / beaten dog) or those in non-nominative case (e.g. ’hrano imam skuhano’ / I have the food cooked[acc]) will be classified as (qualitative) adjectives.

• The adjectival (e.g. ’stokajoč’ / moaning) and adverbial (e.g. ’leže’ / lying down) participles are classified as adjectives and adverbs respectively.

• Negative is always marked as ’n’ except for two verbs: ’nočem’ / not to want, ’nisem’ / not to be

Combinations

*** **** **** **** **** **** **** ---- ---- =======================
PoS Type VFrm Tens Pers Numb Gend Voic Neg  Examples
*** **** **** **** **** **** **** ---- ---- =======================
V   c     n    -    -    -    -    -   -   biti
V   c     c    -    -    -    -    -   -   bi
V   c     m    p    1   [pd]  -    -   -   bodimo, bodiva
V   c     m    p    2   [spd] -    -   -   bodi, bodite, bodita
V   c     i    f  [123] [spd] -    -   -   bom, boš, bo, bomo, bova,...
V   c     i    p  [123] [spd] -    - [ny]  nisem, ..., sem, si, je, smo, ...
V   c     p    s    -   [spd][mfn] a   -   bil, bila, bili, ...
V   o     n    -    -     -    -   -   -   hoteti
V   o     m    p    1   [pd]  -    -   -   dovolimo, dovoliva
V   o     m    p    2   [spd] -    -   -   dovoli, dovolite, dovolita
V   o     i    p  [123] [spd]  -   - [ny]  nočem, ..., hočem, hočeš
V   o     p    s    -   [spd][mfn] a   -   smel, smela, smeli,...
V   o     p    -    -   [spd][mfn] p   -   dovoljen, dovoljeni, ...
V   m     n    -    -    -    -    -   -   delati
V   m     u    -    -    -    -    -   -   delat
V   m     m    p    1   [pd]  -    -   -   delajmo, delajva
V   m     m    p    2   [spd] -    -   -   delaj, delajte, delajta
V   m     i    p  [123] [spd] -    - [ny]  nimam, delam, delava, delamo
V   m     p    -    -   [spd][mfn] p   -   dovoljen, dovoljeni, ...
V   m     p    s    -   [spd][mfn] p   -   dovolil, dovolili, ...
*** **** **** **** **** **** **** ---- ---- =======================

Lexicon

128 MSDs:

Vcc(bi / biti) Vcif1d(bova / biti) Vcif1p(bomo / biti) Vcif1s(bom / biti) Vcif2d(bosta / biti) Vcif2p(boste / biti) Vcif2s(boš / biti) Vcif3d(bosta / biti) Vcif3p(bodo / biti) Vcif3s(bo / biti) Vcip1d–n(sva / biti) Vcip1d–y(nisva / biti) Vcip1p–n(smo / biti) Vcip1p–y(nismo / biti) Vcip1s–n(sem / biti) Vcip1s–y(nisem / biti) Vcip2d–n(sta / biti) Vcip2d–y(nista / biti) Vcip2p–n(ste / biti) Vcip2p–y(niste / biti) Vcip2s–n(si / biti) Vcip2s–y(nisi / biti) Vcip3d–n(sta / biti) Vcip3d–y(nista / biti) Vcip3p–n(so / biti) Vcip3p– y(niso / biti) Vcip3s–n(je / biti) Vcip3s–y(ni / biti) Vcmp1d(bodiva / biti) Vcmp1p(bodimo / biti) Vcmp2d(bodita / biti) Vcmp2p(bodite / biti) Vcmp2s(bodi / biti) Vcn(biti / =) Vcps-dfa(bili / biti) Vcps-dma(bila / biti) Vcps-dna(bili / biti) Vcps-pfa(bile / biti) Vcps-pma(bili / biti) Vcps-pna(bila / biti) Vcps-sfa(bila / biti) Vcps-sma(bil / biti) Vcps-sna(bilo / biti) Vmip1d–n(čakava / čakati) Vmip1d– y(nimava / imeti) Vmip1p–n(čakamo / čakati) Vmip1p–y(nimamo / imeti) Vmip1s–n(čakam / čakati) Vmip1s–y(nimam / imeti) Vmip2d–n(čakata / čakati) Vmip2d–y(nimata / imeti) Vmip2p–n(čakate / čakati) Vmip2p–y(nimate / imeti) Vmip2s–n(čakaš / čakati) Vmip2s–y(nimaš / imeti) Vmip3d–n(čakata / čakati) Vmip3d–y(nimata / imeti) Vmip3p–n(čakajo / čakati) Vmip3p–y(nimajo / imeti) Vmip3s–n(čaka / čakati) Vmip3s–y(nima / imeti) Vmmp1d(čakajva / čakati) Vmmp1p(čakajmo / čakati) Vmmp2d(čakajta / čakati) Vmmp2p(čakajte / čakati) Vmmp2s(čakaj / čakati) Vmn(čakati / =) Vmp–dfp(čaščeni / častiti) Vmp–dmp(čaščena / častiti) Vmp–dnp(čaščeni / častiti) Vmp–pfp(čaščene / častiti) Vmp–pmp(čaščeni / častiti) Vmp–pnp(čaščena / častiti) Vmp–sfp(čaščena / častiti) Vmp–smp(čaščen / častiti) Vmp– snp(čaščeno / častiti) Vmps-dfa(čakali / čakati) Vmps-dma(čakala / čakati) Vmps-dna(čakali / čakati) Vmps-pfa(čakale / čakati) Vmps-pma(čakali / čakati) Vmps-pna(čakala / čakati) Vmps-sfa(čakala / čakati) Vmps-sma(čakal / čakati) Vmps-sna(čakalo / čakati) Vmu(čakat / čakati) Voip1d–n(želiva / želeti) Voip1d–y(nočeva / hoteti) Voip1p–n(želimo / želeti) Voip1p–y(nočemo / hoteti) Voip1s–n(želim / želeti) Voip1s–y(nočem / hoteti) Voip2d–n(želita / želeti) Voip2d–y(nočeta / hoteti) Voip2p–n(želite / želeti) Voip2p–y(nočete / hoteti) Voip2s–n(želiš / želeti) Voip2s–y(nočeš / hoteti) Voip3d–n(želita / želeti) Voip3d–y(nočeta / hoteti) Voip3p–n(želijo / želeti) Voip3p–y(nočejo / hoteti) Voip3s–n(želi / želeti) Voip3s–y(noče / hoteti) Vomp1d(želiva / želeti) Vomp1p(želimo / želeti) Vomp2d(želita / želeti) Vomp2p(želite / želeti) Vomp2s(želi / želeti) Von(želeti / =) Vop–dfp(želeni / želeti) Vop–dmp(želena / želeti) Vop–dnp(želeni / želeti) Vop–pfp(želene / želeti) Vop–pmp(želeni / želeti) Vop–pnp(želena / želeti) Vop–sfp(želena / želeti) Vop–smp(želen / želeti) Vop–snp(želeno / želeti) Vops-dfa(želeli / želeti) Vops-dma(želela / želeti) Vops-dna(želeli / želeti) Vops-pfa(želele / želeti) Vops-pma(želeli / želeti) Vopspna(želela / želeti) Vops-sfa(želela / želeti) Vops-sma(želel / želeti) Vops-sna(želelo / želeti)

6.5 Adjective (A) // Pridevnik (P)

= ============== ============== ==== ============== ============== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl         C-sl
= ============== ============== ==== ============== ============== ====
1 Type           qualificative  f    vrsta          kakovostni     k
                 possessive     s                   svojilni       s
           l.s.  ordinal        o                   vrstni         v
- -------------- -------------- ---- -------------- -------------- ----
2 Degree         positive       p    stopnja        osnovnik       o
                 comparative    c                   primernik      p
                 superlative    s                   presežnik      s
 !NEW!           elative        e                   elativ         e
- -------------- -------------- ---- -------------- -------------- ----
3 Gender         masculine      m    spol           moški          m
                 feminine       f                   ženski         z
                 neuter         n                   srednji        s
- -------------- -------------- ---- -------------- -------------- ----
4 Number         singular       s    število        ednina         e
                 plural         p                   množina        m
                 dual           d                   dvojina        d
- -------------- -------------- ---- -------------- -------------- ----
5 Case           nominative     n    sklon          imenovalnik    i
                 genitive       g                   rodilnik       r
                 dative         d                   dajalnik       d
                 accusative     a                   tožilnik       t
                 locative       l                   mestnik        m
                 instrumental   i                   orodnik        o
* ***************************** **** ************** ************** ****
6 Definiteness   no             n    določnost      ne             n
                 yes            y                   da             d
- -------------- -------------- ---- -------------- -------------- ----
7 Clitic                        -                                 -
- -------------- -------------- ---- -------------- -------------- ----
8 Animate        no             n    živost         ne             n
                 yes            y                   da             d
- -------------- -------------- ---- -------------- -------------- ----
9 Formation                     -
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
---------------- -------------- -
12Owned_Number                  -
=============================== =

Notes:

• The three deverbative adjectival participles (part. -l, part -n / t, part. č) are also classified as qualificative adjectives. e.g. ’otekel (obraz), premagan, (sovrag) govoreč (pes)’ are marked Afpmsn.

• Definiteness (-i ending = yes, -0 = no) is marked only on the Adjective qualificative positive masculine singular nominative (Afpmsn) and syncretic accusative (Afpmsa). The uninflecting ’poceni’ has no definiteness.

• Animacy is marked only (-ega ending = yes, else = no) on the Adjective masculine singular accusative (A**msa*-[ny]). Afpmsa has three forms: no-Definiteness, +Animate; -Definiteness, -Animate; +Definiteness, -Animate. The uninflecting ’poceni’ has no animacy or definiteness.

• Some (26) adjectives have the additional ’archaic’ indefinite form also (-a- and -e-, both weak and thus absent from all oblique forms): ’hladan’ in addition to ’hladen’ / cold, both with Afpmsnn (non-definite) and syncretic Afpmsan-n (non-definite, non-animate). This could be encoded (but isn’t) in the Formation attribute.

These are the archaic Afpmsnn: šibak dolžan droban gnojan grenak hladan krasan krepak krotak lahak mehak močan motan nemočan plehak prešibak premočan pretežak rosan sladak strašan svetal težak teman tesan voljan

Combinations

*** **** **** **** **** **** ---- ---- ---- ================================
PoS Type Degr Gend Numb Case Def (Cltc)Anim Examples
*** **** **** **** **** **** ---- ---- ---- ================================
A   s     p [mfs][spd][ngdali] -    -   -   stričev, galjotove, ...
A   o     p [mfs][spd][ngdali] -    -   -   dlakocepski, slovenskega, ...
A   f     p   m   s      n   [ny]   -   -   lep, lepi
A   f     p   m   s      a   [ny]   -   n   lep, lepi
A   f     p   m   s      a     -    -   y   lepega
A   f     p   m   s    [gdli]  -    -   -   lepega, lepemu, lepem, lepim
A   f     p   m  [pd] [ngdali] -    -   -   lepi, lepih, ..., lepa, ...
A   f     p  [fs][spd][ngdali] -    -   -   lepa, lepe, ..., lepo, ...
A   f   [cs][mfs][spd][ngdali] -    -   -   lepši, lepšega,...,najlepši,...
*** **** **** **** **** **** ================================================

Lexicon

279 MSDs:

Afcfda(čistejši / čist) Afcfdd(čistejšima / čist) Afcfdg(čistejših / čist) Afcfdi(čistejšima / čist) Afcfdl(čistejših / čist) Afcfdn(čistejši / čist) Afcfpa(čistejše / čist) Afcfpd(čistejšim / čist) Afcfpg(čistejših / čist) Afcfpi(čistejšimi / čist) Afcfpl(čistejših / čist) Afcfpn(čistejše / čist) Afcfsa(čistejšo / čist) Afcfsd(čistejši / čist) Afcfsg(čistejše / čist) Afcfsi(čistejšo / čist) Afcfsl(čistejši / čist) Afcfsn(čistejša / čist) Afcmda(čistejša / čist) Afcmdd(čistejšima / čist) Afcmdg(čistejših / čist) Afcmdi(čistejšima / čist) Afcmdl(čistejših / čist) Afcmdn(čistejša / čist) Afcmpa(čistejše / čist) Afcmpd(čistejšim / čist) Afcmpg(čistejših / čist) Afcmpi(čistejšimi / čist)

Afcmpl(čistejših / čist) Afcmpn(čistejši / čist) Afcmsa–n(čistejši / čist) Afcmsa–y(čistejšega / čist) Afcmsd(čistejšemu / čist) Afcmsg(čistejšega / čist) Afcmsi(čistejšim / čist) Afcmsl(čistejšem / čist) Afcmsn(čistejši / čist) Afcnda(čistejši / čist) Afcndd(čistejšima / čist) Afcndg(čistejših / čist) Afcndi(čistejšima / čist) Afcndl(čistejših / čist) Afcndn(čistejši / čist) Afcnpa(čistejša / čist) Afcnpd(čistejšim / čist) Afcnpg(čistejših / čist) Afcnpi(čistejšimi / čist) Afcnpl(čistejših / čist) Afcnpn(čistejša / čist) Afcnsa(čistejše / čist) Afcnsd(čistejšemu / čist) Afcnsg(čistejšega / čist) Afcnsi(čistejšim / čist) Afcnsl(čistejšem / čist) Afcnsn(čistejše / čist) Afpfda(čajni / čajen) Afpfdd(čajnima / čajen) Afpfdg(čajnih / čajen) Afpfdi(čajnima / čajen) Afpfdl(čajnih / čajen) Afpfdn(čajni / čajen) Afpfpa(čajne / čajen) Afpfpd(čajnim / čajen) Afpfpg(čajnih / čajen) Afpfpi(čajnimi / čajen) Afpfpl(čajnih / čajen) Afpfpn(čajne / čajen) Afpfsa(čajno / čajen) Afpfsd(čajni / čajen) Afpfsg(čajne / čajen) Afpfsi(čajno / čajen) Afpfsl(čajni / čajen) Afpfsn(čajna / čajen) Afpmda(čajna / čajen) Afpmdd(čajnima / čajen) Afpmdg(čajnih / čajen) Afpmdi(čajnima / čajen) Afpmdl(čajnih / čajen) Afpmdn(čajna / čajen) Afpmpa(čajne / čajen) Afpmpd(čajnim / čajen) Afpmpg(čajnih / čajen) Afpmpi(čajnimi / čajen) Afpmpl(čajnih / čajen) Afpmpn(čajni / čajen) Afpmsa(poceni / poceni) Afpmsa–y(čajnega / čajen) Afpmsan-n(čajen / čajen) Afpmsay-n(čajni / čajen) Afpmsd(čajnemu / čajen) Afpmsg(čajnega / čajen) Afpmsi(čajnim / čajen) Afpmsl(čajnem / čajen) Afpmsn(poceni / =) Afpmsnn(čajen / =) Afpmsny(čajni / čajen) Afpnda(čajni / čajen) Afpndd(čajnima / čajen) Afpndg(čajnih / čajen) Afpndi(čajnima / čajen) Afpndl(čajnih / čajen) Afpndn(čajni / čajen) Afpnpa(čajna / čajen) Afpnpd(čajnim / čajen) Afpnpg(čajnih / čajen) Afpnpi(čajnimi / čajen) Afpnpl(čajnih / čajen) Afpnpn(čajna / čajen) Afpnsa(čajno / čajen) Afpnsd(čajnemu / čajen) Afpnsg(čajnega / čajen) Afpnsi(čajnim / čajen) Afpnsl(čajnem / čajen) Afpnsn(čajno / čajen) Afsfda(najčistejši / čist) Afsfdd(najčistejšima / čist) Afsfdg(najčistejših / čist) Afsfdi(najčistejšima / čist) Afsfdl(najčistejših / čist) Afsfdn(najčistejši / čist) Afsfpa(najčistejše / čist) Afsfpd(najčistejšim / čist) Afsfpg(najčistejših / čist) Afsfpi(najčistejšimi / čist) Afsfpl(najčistejših / čist) Afsfpn(najčistejše / čist) Afsfsa(najčistejšo / čist) Afsfsd(najčistejši / čist) Afsfsg(najčistejše / čist) Afsfsi(najčistejšo / čist) Afsfsl(najčistejši / čist) Afsfsn(najčistejša / čist) Afsmda(najčistejša / čist) Afsmdd(najčistejšima / čist) Afsmdg(najčistejših / čist) Afsmdi(najčistejšima / čist) Afsmdl(najčistejših / čist) Afsmdn(najčistejša / čist) Afsmpa(najčistejše / čist) Afsmpd(najčistejšim / čist) Afsmpg(najčistejših / čist) Afsmpi(najčistejšimi / čist) Afsmpl(najčistejših / čist) Afsmpn(najčistejši / čist) Afsmsa–n(najčistejši / čist) Afsmsa–y(najčistejšega / čist) Afsmsd(najčistejšemu / čist) Afsmsg(najčistejšega / čist) Afsmsi(najčistejšim / čist) Afsmsl(najčistejšem / čist) Afsmsn(najčistejši / čist) Afsnda(najčistejši / čist) Afsndd(najčistejšima / čist) Afsndg(najčistejših / čist) Afsndi(najčistejšima / čist) Afsndl(najčistejših / čist) Afsndn(najčistejši / čist) Afsnpa(najčistejša / čist) Afsnpd(najčistejšim / čist) Afsnpg(najčistejših / čist) Afsnpi(najčistejšimi / čist) Afsnpl(najčistejših / čist) Afsnpn(najčistejša / čist) Afsnsa(najčistejše / čist) Afsnsd(najčistejšemu / čist) Afsnsg(najčistejšega / čist) Afsnsi(najčistejšim / čist) Afsnsl(najčistejšem / čist) Afsnsn(najčistejše / čist) Aopfda(čarovniški / čarovniški) Aopfdd(čarovniškima / čarovniški) Aopfdg(čarovniških / čarovniški) Aopfdi(čarovniškima / čarovniški) Aopfdl(čarovniških / čarovniški) Aopfdn(čarovniški / čarovniški) Aopfpa(čarovniške / čarovniški) Aopfpd(čarovniškim / čarovniški) Aopfpg(čarovniških / čarovniški) Aopfpi(čarovniškimi / čarovniški) Aopfpl(čarovniških / čarovniški) Aopfpn(čarovniške / čarovniški) Aopfsa(čarovniško / čarovniški) Aopfsd(čarovniški / čarovniški) Aopfsg(čarovniške / čarovniški) Aopfsi(čarovniško / čarovniški) Aopfsl(čarovniški / čarovniški) Aopfsn(čarovniška / čarovniški) Aopmda(čarovniška / čarovniški) Aopmdd(čarovniškima / čarovniški) Aopmdg(čarovniških / čarovniški) Aopmdi(čarovniškima / čarovniški) Aopmdl(čarovniških / čarovniški) Aopmdn(čarovniška / čarovniški) Aopmpa(čarovniške / čarovniški) Aopmpd(čarovniškim / čarovniški) Aopmpg(čarovniških / čarovniški) Aopmpi(čarovniškimi / čarovniški) Aopmpl(čarovniških / čarovniški) Aopmpn(čarovniški / čarovniški) Aopmsa–n(čarovniški / čarovniški) Aopmsa–y(čarovniškega / čarovniški) Aopmsd(čarovniškemu / čarovniški) Aopmsg(čarovniškega / čarovniški) Aopmsi(čarovniškim / čarovniški) Aopmsl(čarovniškem / čarovniški)

Aopmsn(čarovniški / =) Aopnda(čarovniški / čarovniški) Aopndd(čarovniškima / čarovniški) Aopndg(čarovniških / čarovniški) Aopndi(čarovniškima / čarovniški) Aopndl(čarovniških / čarovniški) Aopndn(čarovniški / čarovniški) Aopnpa(čarovniška / čarovniški) Aopnpd(čarovniškim / čarovniški) Aopnpg(čarovniških / čarovniški) Aopnpi(čarovniškimi / čarovniški) Aopnpl(čarovniških / čarovniški) Aopnpn(čarovniška / čarovniški) Aopnsa(čarovniško / čarovniški) Aopnsd(čarovniškemu / čarovniški) Aopnsg(čarovniškega / čarovniški) Aopnsi(čarovniškim / čarovniški) Aopnsl(čarovniškem / čarovniški) Aopnsn(čarovniško / čarovniški) Aspfda(Šeligovi / Šeligov) Aspfdd(Šeligovima / Šeligov) Aspfdg(Šeligovih / Šeligov) Aspfdi(Šeligovima / Šeligov) Aspfdl(Šeligovih / Šeligov) Aspfdn(Šeligovi / Šeligov) Aspfpa(Šeligove / Šeligov) Aspfpd(Šeligovim / Šeligov) Aspfpg(Šeligovih / Šeligov) Aspfpi(Šeligovimi / Šeligov) Aspfpl(Šeligovih / Šeligov) Aspfpn(Šeligove / Šeligov) Aspfsa(Šeligovo / Šeligov) Aspfsd(Šeligovi / Šeligov) Aspfsg(Šeligove / Šeligov) Aspfsi(Šeligovo / Šeligov) Aspfsl(Šeligovi / Šeligov) Aspfsn(Šeligova / Šeligov) Aspmda(Šeligova / Šeligov) Aspmdd(Šeligovima / Šeligov) Aspmdg(Šeligovih / Šeligov) Aspmdi(Šeligovima / Šeligov) Aspmdl(Šeligovih / Šeligov) Aspmdn(Šeligova / Šeligov) Aspmpa(Šeligove / Šeligov) Aspmpd(Šeligovim / Šeligov) Aspmpg(Šeligovih / Šeligov) Aspmpi(Šeligovimi / Šeligov) Aspmpl(Šeligovih / Šeligov) Aspmpn(Šeligovi / Šeligov) Aspmsa– n(Šeligov / Šeligov) Aspmsa–y(Šeligovega / Šeligov) Aspmsd(Šeligovemu / Šeligov) Aspmsg(Šeligovega / Šeligov) Aspmsi(Šeligovim / Šeligov) Aspmsl(Šeligovem / Šeligov) Aspmsn(Šeligov / =) Aspnda(Šeligovi / Šeligov) Aspndd(Šeligovima / Šeligov) Aspndg(Šeligovih / Šeligov) Aspndi(Šeligovima / Šeligov) Aspndl(Šeligovih / Šeligov) Aspndn(Šeligovi / Šeligov) Aspnpa(Šeligova / Šeligov) Aspnpd(Šeligovim / Šeligov) Aspnpg(Šeligovih / Šeligov) Aspnpi(Šeligovimi / Šeligov) Aspnpl(Šeligovih / Šeligov) Aspnpn(Šeligova / Šeligov) Aspnsa(Šeligovo / Šeligov) Aspnsd(Šeligovemu / Šeligov) Aspnsg(Šeligovega / Šeligov) Aspnsi(Šeligovim / Šeligov) Aspnsl(Šeligovem / Šeligov) Aspnsn(Šeligovo / Šeligov)

6.6 Adverb (R) // Prislov (R)

= ============== ============== ==== ============== ============== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl         C-sl
= ============== ============== ==== ============== ============== ====
1 Type           general        g    vrsta          splošni        s
- -------------- -------------- ---- -------------- -------------- ----
2 Degree         positive       p    stopnja        osnovnik       o
                 comparative    c                   primernik      p
                 superlative    s                   presežnik      r
 !NEW!           elative        e                   elativ         e
******************************* **** ************** ************** ****
3 Clitic                        -
- -------------- -------------- -
4 Number                        -
- -------------- -------------- -
5 Person                        -
=================================

Notes:

• The adverbial (e.g. part -e: ’leže’ / lying down) participles are classified as adverbs.

• Some pronouns can also have an adverbial function. This issue has not been considered here.

Combinations

*** **** **** ==============================================================
PoS Type Degr Examples
*** **** **** ==============================================================
 R    g    p  bljuvaje
 R    g    c  kasneje
 R    g    s  najprijazneje
 R    g    e  ???
*** **** **** ==============================================================

Lexicon

3 MSDs:

Rgc(čistejše / čist) Rgp(Šeligovo / Šeligov) Rgs(najčistejše / čist)

6.7 Pronoun (P) // Zaimek (Z)

= ============== ============== ==== ============== ================== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl             C-sl
= ============== ============== ==== ============== ================== ====
1 Type           personal       p    vrsta          osebni             o
demonstrative  d                   kazalni            k
indefinite     i                   nedoločnostni      n
possessive     s                   svojilni           s
interrogative  q                   vprašalni          v
relative       r                   oziralnostni       z
reflexive      x                   povratni           p
negative       z                   nikalni            l
general        g                   celostni           c
- -------------- -------------- ---- -------------- ------------------ ----
2 Person         first          1    oseba          prva               1
second         2                   druga              2
third          3                   tretja             3
- -------------- -------------- ---- -------------- ------------------ ----
3 Gender         masculine      m    spol           moški              m
feminine       f                   ženski             z
neuter         n                   srednji            s
- -------------- -------------- ---- -------------- ------------------ ----
4 Number         singular       s    število        ednina             e
plural         p                   množina            m
dual           d                   dvojina            d
- -------------- -------------- ---- -------------- ------------------ ----
5 Case           nominative     n    sklon          imenovalnik        i
genitive       g                   rodilnik           r
dative         d                   dajalnik           d
accusative     a                   tožilnik           t
locative       l                   mestnik            m
instrumental   i                   orodnik            o
- -------------- -------------- ---- -------------- ------------------ ----
6 Owner_Number   singular       s    št_svojine     ednina             e
plural         p                   množina            m
dual           d                   dvojina            d
- -------------- -------------- ---- -------------- ------------------ ----
7 Owner_Gender   masculine      m    sp_svojine     moški              m
feminine       f                   ženski             z
neuter         n                   srednji            s
******************************* **** ************** ****************** ****
8 Clitic         no             n    naslonka       ne                 n
yes            y                   da                 d
!NEW!           bound          b                   navezna            z
- -------------- -------------- ---- -------------- ------------------ ----

9 Referent_Type  personal       p    nanašanje      osebni             o
                 possessive     s                   svojilni           s
- -------------- -------------- ---- -------------- ------------------ ----
10Syntactic_Type nominal        n    sklad_vloga    samostalniški      s
                 adjectival     a                   pridevniški        p
 !NEW!           adverbial      r                   prislovni          r
- -------------- -------------- ---- -------------- ------------------ ----
11Definiteness                  -                                     -
- -------------- -------------- ---- -------------- ------------------ ----
12Animate        no             n    živost         ne                 n
                 yes            y                   da                 d
- -------------- -------------- ---- -------------- ------------------ ----
13Clitic_s                      -
- -------------- -------------- -
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person                  -
- -------------- -------------- -
16Owned_Number                  -
=================================

Notes:

• The Type taxonomy for pronouns is different from the one of Slovene grammars. Mostly Sl. grammar classes were merged. In particular, the indefinite and relative types contain many of pronoun types from Slovene grammars: (nedoločnostni = nedoločni + poljubnostni + mnogostni + istostni + drugostni) (oziralnostni = oziralni + oziralno poljubnostni)

• Type=relative includes so called ’relative general’ (oziralno poljubnostni) pronouns. They can consist of two words, where the second is ’koli’. Variant spelling exists: ’kolikorkoli’, ’kolikor koli’. Possible solution: they can be treated as compounds.

• Type=reflexive ecompasses all reflexive pronouns (sebe, se, sebi, si, svoj) as well as ’se’ and ’si’ in all their roles. Personal and possessive reflexives are further distinguished via the Referent_Type attribute. ’se’ and ’si’ in their roles as an obligatory but dislocated constituent’ of reflexive verbs have the unique MSD Px——y. ’se’ can also appear as a suffix: e.g. ’zase’ / for_oneself with variant spelling ’za se’.

• Owner_Gender is relevant only for third person singular possessive pronouns, and distinguishes only masculine and feminine forms. It will be ’-’ in all other cases.

• As in Slovene grammars, pronouns are distinguished between having a nominal and adjectival function. All pronoun types except demonstrative and possessive can be nominal, and all except personal can be adjectival. However, the distinction is often hard to make in practice.

• Referent_Type is used to distinguish personal reflexives (which include ’se’ in all its functions) from the possesive reflexives (’svoj’).

• Animacy is marked only on the masculine singular accusative (P..msa) of non-personal Pronouns. The P..msa ending with -ega ending is +Animate, the other -Animate, cf. Nouns and Adjectives.

Combinations

*** **** **** **** **** **** **** **** --- --- --- --- --- ================
PoS Type Pers Gend Numb Case PosN PosG Clc Ref Syn Def Ani Examples
*** **** **** **** **** **** **** **** --- --- --- --- --- ================
P    x    -    -    -    -    -    -    y   -   -   -   -  se, si
P    x    -    -    -  [gdali]-    -    y   p   n   -   -  se, si,...
P    x    -    -    -  [gdali]-    -    n   p   n   -   -  sebe,sebi,...
P    x    -    m    s    a    -    -    -   s   a   - [ny] svoj, svojega
P    x    -    m    s  [ngdli]-    -    -   s   a   -   -  svoj, svojega,...
P    x    -    m  [pd][ngdali]-    -    -   s   a   -   -  svoji, svoja,...
P    x    -  [fn][spd][ngdali]-    -    -   s   a   -   -  svoja, svoje,...
P    p  [123]  - [spd][ngdali]-    -  [ny]  -   a   -   -  njega, ga, jih, ...
P    s  [12][mfn][spd][ngdali][spd]-    -   -   a   - [ny] moj, vajinima, ...
P    s   [3][mfn][spd][ngdali]s  [mfn]  -   -   a   - [ny] njegov, njen, njegova,...
P    s   [3][mfn][spd][ngdali][pd] -    -   -   a   - [ny] njihova, njuna, ...
P    d    -    -    -    -    -    -    -   -   a   -   -  le-tega, onega, ...
P    d    - [mfn][spd][ngdali]-    -    -   - [an]  - [ny] le-ta, le-tega, onega, ...
P    i    -    -    -    -    -    -    -   -   a   -   -  koliko, čigar
P    i    -    -    - [ngdali]-    -    -   -   n   -   -  kaj, marsikaj, ...
P    i    - [mfn][spd][ngdali]-    -    -   - [an]  - [ny] en, tegale, drug, ...
P    q    -    -    -    -    -    -    -   -   a   -   -  koliko
P    q    -    -    - [ngdali]-    -    -   -   n   -   -  kaj, česa, kdo, ...
P    q    - [mfn][spd][ngdali]-    -    -   -   a   - [ny] kakšen, kateri, ...
P    r    -    -    -    -    -    -    -   -   a   -   -  kolikor, čigar
P    r    -    -    - [ngdali]-    -    -   -   n   -   -  kar, česar, kdorkoli, ...
P    r    - [mfn][spd][ngdali]-    -    -   -   a   - [ny] kakršen, kateri
P    z    -    -    -    -    -    -    -   -   a   -   -  nič, nikogar, noben
P    z    -    -    - [gdali] -    -    -   -   n   -   -  nikogar, nikomur, ...
P    z    - [mfn][spd][ngdali]-    -    -   -   a   - [ny] nikakršen, noben
P    g    -    -   -     -     -   -    -   -   a   -   -  vsakogar
P    g    -    -   -  [ngdali] -   -    -   -   n   -   -  vsakdo, vsakogar, ...
P    g    - [mfn][spd][ngdali] -   -    -   -  [an] - [ny] ves, vsega, vsa, vsak, ...
*** **** **** **** **** **** **** **** --- --- --- --- --- ================

Lexicon

1335 MSDs:

Pd——–a(le-onega / =) Pd-fda—-a(le-oni / le-oni) Pd-fda—-n(le-te / le-ta) Pd-fdd—-a(le-onima / le-oni) Pdfdd—-n(le-tema / le-ta) Pd-fdg—-a(le-onih / le-oni) Pd-fdg—-n(le-teh / le-ta) Pd-fdi—-a(le-onima / le-oni) Pdfdi—-n(le-tema / le-ta) Pd-fdl—-a(le-onih / le-oni) Pd-fdl—-n(le-teh / le-ta) Pd-fdn—-a(le-oni / le-oni) Pd-fdn—- n(le-te / le-ta) Pd-fpa—-a(le-one / le-oni) Pd-fpa—-n(le-te / le-ta) Pd-fpd—-a(le-onim / le-oni) Pd-fpd—-n(le-tem / le-ta) Pd-fpg—-a(le-onih / le-oni) Pd-fpg—-n(le-teh / le-ta) Pd-fpi—-a(le-onimi / le-oni) Pd-fpi—-n(le-temi / le-ta) Pd-fpl—-a(le-onih / le-oni) Pd-fpl—-n(le-teh / le-ta) Pd-fpn—-a(le-one / le-oni) Pd-fpn—-n(le-te / le-ta) Pd-fsa—-a(le-ono / le-oni) Pd-fsa—-n(le-to / le-ta) Pd-fsd—-a(le-oni / le-oni) Pd-fsd—-n(le-tej / le-ta) Pd-fsg— -a(le-one / le-oni) Pd-fsg—-n(le-te / le-ta) Pd-fsi—-a(le-ono / le-oni) Pd-fsi—-n(le-to / le-ta) Pd-fsl—-a(le-oni / le-oni) Pd-fsl—-n(le-tej / le-ta) Pd-fsn—-a(le-ona / le-oni) Pd-fsn—-n(le-ta / le-ta) Pd-mda—-a(le-ona / le-oni) Pd-mda—-n(le-ta / le-ta) Pd-mdd—-a(le-onima / le-oni) Pd-mdd—-n(le-tema / le-ta) Pd-mdg—-a(le-onih / leoni) Pd-mdg—-n(le-teh / le-ta) Pd-mdi—-a(le-onima / le-oni) Pd-mdi—-n(le-tema / le-ta) Pd-mdl—-a(le-onih / le-oni) Pd-mdl—-n(le-teh / le-ta) Pd-mdn—-a(le-ona / le-oni) Pd-mdn—-n(le-ta / le-ta) Pd-mpa—-a(le-one / le-oni) Pd-mpa—-n(le-te / le-ta) Pd-mpd—-a(le-onim / le-oni) Pd-mpd—-n(le-tem / le-ta) Pd-mpg—-a(le-onih / le-oni) Pd-mpg—-n(le-teh / le-ta) Pd-mpi—-a(le-onimi / le-oni) Pd-mpi—-n(le-temi / le-ta) Pd-mpl—-a(le-onih / le-oni) Pd-mpl—-n(le-teh / le-ta) Pd-mpn—-a(le-oni / le-oni) Pd-mpn—-n(le-ti / le-ta) Pd-msa—-a-n(le-oni / le-oni) Pd-msa—-a-y(le-onega / le-oni) Pd-msa—-n-n(le-ta / le-ta) Pd-msa—-n-y(le-tega / le-ta) Pd-msd—-a(le-onemu / le-oni) Pd-msd—-n(le-temu / le-ta) Pd-msg—-a(le-onega / le-oni) Pd-msg—-n(le-tega / le-ta) Pd-msi— -a(le-onim / le-oni) Pd-msi—-n(le-tem / le-ta) Pd-msl—-a(le-onem / le-oni) Pd-msl—-n(le-tem / le-ta) Pd-msn— -a(le-oni / =) Pd-msn—-n(le-ta / =) Pd-nda—-a(le-oni / le-oni) Pd-nda—-n(le-te / le-ta) Pd-ndd—-a(le-onima / le-oni) Pd-ndd—-n(le-tema / le-ta) Pd-ndg—-a(le-onih / le-oni) Pd-ndg—-n(le-teh / le-ta) Pd-ndi—-a(le-onima / le-oni) Pd-ndi—-n(le-tema / le-ta) Pd-ndl—-a(le-onih / le-oni) Pd-ndl—-n(le-teh / le-ta) Pd-ndn—-a(le-oni / le-oni) Pd-ndn—-n(le-te / le-ta) Pd-npa—-a(le-ona / le-oni) Pd-npa—-n(le-ta / le-ta) Pd-npd—-a(le-onim / le-oni) Pd-npd—-n(le-tem / le-ta) Pd-npg—-a(le-onih / le-oni) Pd-npg—-n(le-teh / le-ta) Pd-npi—-a(le-onimi / le-oni) Pd-npi—-n(le-temi / le-ta) Pd-npl—-a(le-onih / le-oni) Pd-npl—-n(le-teh / le-ta) Pd-npn—-a(le-ona / leoni) Pd-npn—-n(le-ta / le-ta) Pd-nsa—-a(le-ono / le-oni) Pd-nsa—-n(le-to / le-ta) Pd-nsd—-a(le-onemu / le-oni) Pd-nsd—-n(le-temu / le-ta) Pd-nsg—-a(le-onega / le-oni) Pd-nsg—-n(le-tega / le-ta) Pd-nsi—-a(le-onim / le-oni) Pd-nsi—-n(le-tem / le-ta) Pd-nsl—-a(le-onem / le-oni) Pd-nsl—-n(le-tem / le-ta) Pd-nsn—-a(le-ono / le-oni) Pd-nsn—-n(le-to / le-ta) Pg——–a(vsakogar / =) Pg—a—-n(vsakogar / vsakdo) Pg—d—-n(vsakomur / vsakdo) Pg—g—-n(vsakogar / vsakdo) Pg—i—-n(vsakomer / vsakdo) Pg—l—-n(vsakomer / vsakdo) Pg—n—-n(vsakdo / =) Pg-fda—-a(obe / oba) Pg-fda—-n(obe / oba) Pg-fdd—-a(obema / oba) Pg-fdd—-n(obema / oba) Pg-fdg— -a(obeh / oba) Pg-fdg—-n(obeh / oba) Pg-fdi—-a(obema / oba) Pg-fdi—-n(obema / oba) Pg-fdl—-a(obeh / oba) Pg-fdl—-n(obeh / oba) Pg-fdn—-a(obe / oba) Pg-fdn—-n(obe / oba) Pg-fpa—-a(vsake / vsak) Pg-fpa—- n(vse / ves) Pg-fpd—-a(vsakim / vsak) Pg-fpd—-n(vsem / ves) Pg-fpg—-a(vsakih / vsak) Pg-fpg—-n(vseh / ves) Pg-fpi—-a(vsakimi / vsak) Pg-fpi—-n(vsemi / ves) Pg-fpl—-a(vsakih / vsak) Pg-fpl—-n(vseh / ves) Pg-fpn—- a(vsake / vsak) Pg-fpn—-n(vse / ves) Pg-fsa—-a(vsako / vsak) Pg-fsa—-n(vso / ves) Pg-fsd—-a(vsaki / vsak) Pg-fsd—-n(vsej / ves) Pg-fsg—-a(vsake / vsak) Pg-fsg—-n(vse / ves) Pg-fsi—-a(vsako / vsak) Pg-fsi—-n(vso / ves) Pg-fsl—-a(vsaki / vsak) Pg-fsl—-n(vsej / ves) Pg-fsn—-a(vsa / ves) Pg-fsn—-n(vsa / ves) Pg-mda—-a(oba / oba) Pg-mda—-n(oba / oba) Pg-mdd—-a(obema / oba) Pg-mdd—-n(obema / oba) Pg-mdg—-a(obeh / oba) Pg-mdg—-n(obeh / oba) Pg-mdi—-a(obema / oba) Pg-mdi—-n(obema / oba) Pg-mdl—-a(obeh / oba) Pg-mdl— -n(obeh / oba) Pg-mdn—-a(oba / oba) Pg-mdn—-n(oba / oba) Pg-mpa—-a(vsake / vsak) Pg-mpa—-n(vse / ves) Pg-mpd—-a(vsakim / vsak) Pg-mpd—-n(vsem / ves) Pg-mpg—-a(vsakih / vsak) Pg-mpg—-n(vseh / ves) Pgmpi—-a(vsakimi / vsak) Pg-mpi—-n(vsemi / ves) Pg-mpl—-a(vsakih / vsak) Pg-mpl—-n(vseh / ves) Pg-mpn—- a(vsaki / vsak) Pg-mpn—-n(vsi / ves) Pg-msa—-a-n(ves / ves) Pg-msa—-a-y(vsakega / vsak) Pg-msa—-n-n(ves / ves) Pg-msa—-n-y(vsega / ves) Pg-msd—-a(vsakemu / vsak) Pg-msd—-n(vsemu / ves) Pg-msg—-a(vsakega / vsak) Pg-msg—-n(vsega / ves) Pg-msi—-a(vsakim / vsak) Pg-msi—-n(vsem / ves) Pg-msl—-a(vsakem / vsak) Pg-msl—-n(vsem / ves) Pg-msn—-a(ves / =) Pg-msn—-n(ves / =) Pg-nda—-a(obe / oba) Pg-nda—-n(obe / oba) Pg-ndd—-a(obema / oba) Pg-ndd—-n(obema / oba) Pg-ndg—-a(obeh / oba) Pg-ndg—-n(obeh / oba) Pg-ndi— -a(obema / oba) Pg-ndi—-n(obema / oba) Pg-ndl—-a(obeh / oba) Pg-ndl—-n(obeh / oba) Pg-ndn—-a(obe / oba) Pg-ndn—-n(obe / oba) Pg-npa—-a(vsa / ves) Pg-npa—-n(vsa / ves) Pg-npd—-a(vsakim / vsak) Pg-npd— -n(vsem / ves) Pg-npg—-a(vsakih / vsak) Pg-npg—-n(vseh / ves) Pg-npi—-a(vsakimi / vsak) Pg-npi—-n(vsemi / ves) Pg-npl—-a(vsakih / vsak) Pg-npl—-n(vseh / ves) Pg-npn—-a(vsa / ves) Pg-npn—-n(vsa / ves) Pg-nsa—- a(vsako / vsak) Pg-nsa—-n(vse / ves) Pg-nsd—-a(vsakemu / vsak) Pg-nsd—-n(vsemu / ves) Pg-nsg—-a(vsakega / vsak) Pg-nsg—-n(vsega / ves) Pg-nsi—-a(vsakim / vsak) Pg-nsi—-n(vsem / ves) Pg-nsl—-a(vsakem / vsak) Pg-nsl—-n(vsem / ves) Pg-nsn—-a(vsako / vsak) Pg-nsn—-n(vse / ves) Pi——–a(drugega / =) Pi—a—-n(kaj / kaj) Pi—d—-n(čemu / kaj) Pi—g—-n(česa / kaj) Pi—i—-n(čim / kaj) Pi—l—-n(čem / kaj) Pi—n—-n(kaj / =) Pi-fda—-a(čigavi / čigav) Pi-fda—-n(drugi / drug) Pi-fdd—-a(čigavima / čigav) Pi-fdd—-n(drugima / drug) Pi-fdg—-a(čigavih / čigav) Pi-fdg—-n(drugih / drug) Pi-fdi—-a(čigavima / čigav) Pi-fdi—-n(drugima / drug) Pi-fdl—-a(čigavih / čigav) Pi-fdl—-n(drugih / drug) Pi-fdn—-a(čigavi / čigav) Pi-fdn—-n(drugi / drug) Pi-fpa—-a(čigave / čigav) Pi-fpa—-n(druge / drug) Pi-fpd—-a(čigavim / čigav) Pi-fpd—-n(drugim / drug) Pifpg—-a(čigavih / čigav) Pi-fpg—-n(drugih / drug) Pi-fpi—-a(čigavimi / čigav) Pi-fpi—-n(drugimi / drug) Pifpl—-a(čigavih / čigav) Pi-fpl—-n(drugih / drug) Pi-fpn—-a(čigave / čigav) Pi-fpn—-n(druge / drug) Pi-fsa—- a(čigavo / čigav) Pi-fsa—-n(drugo / drug) Pi-fsd—-a(čigavi / čigav) Pi-fsd—-n(drugi / drug) Pi-fsg—-a(čigave / čigav) Pi-fsg—-n(druge / drug) Pi-fsi—-a(čigavo / čigav) Pi-fsi—-n(drugo / drug) Pi-fsl—-a(čigavi / čigav) Pi-fsl—-n(drugi / drug) Pi-fsn—-a(čigava / čigav) Pi-fsn—-n(druga / drug) Pi-mda—-a(čigava / čigav) Pi-mda— -n(druga / drug) Pi-mdd—-a(čigavima / čigav) Pi-mdd—-n(drugima / drug) Pi-mdg—-a(čigavih / čigav) Pimdg—-n(drugih / drug) Pi-mdi—-a(čigavima / čigav) Pi-mdi—-n(drugima / drug) Pi-mdl—-a(čigavih / čigav) Pi-mdl—-n(drugih / drug) Pi-mdn—-a(čigava / čigav) Pi-mdn—-n(druga / drug) Pi-mpa—-a(čigave / čigav) Pi-mpa—-n(druge / drug) Pi-mpd—-a(čigavim / čigav) Pi-mpd—-n(drugim / drug) Pi-mpg—-a(čigavih / čigav) Pi-mpg—-n(drugih / drug) Pi-mpi—-a(čigavimi / čigav) Pi-mpi—-n(drugimi / drug) Pi-mpl—-a(čigavih / čigav) Pi-mpl—-n(drugih / drug) Pi-mpn—-a(čigavi / čigav) Pi-mpn—-n(drugi / drug) Pi-msa—-a-n(čigav / čigav) Pi-msa—-a-y(čigavega / čigav) Pi-msa—-n-n(drug / drug) Pi-msa—-n-y(drugega / drug) Pi-msd—-a(čigavemu / čigav) Pi-msd—-n(drugemu / drug) Pi-msg—-a(čigavega / čigav) Pi-msg—-n(drugega / drug) Pi-msi—-a(čigavim / čigav) Pi-msi—-n(drugim / drug) Pi-msl—-a(čigavem / čigav) Pi-msl—-n(drugem / drug) Pi-msn—-a(čigav / =) Pi-msn—-n(drug / =) Pi-nda—-a(čigavi / čigav) Pi-nda—-n(drugi / drug) Pi-ndd—-a(čigavima / čigav) Pindd—-n(drugima / drug) Pi-ndg—-a(čigavih / čigav) Pi-ndg—-n(drugih / drug) Pi-ndi—-a(čigavima / čigav) Pi-ndi—-n(drugima / drug) Pi-ndl—-a(čigavih / čigav) Pi-ndl—-n(drugih / drug) Pi-ndn—-a(čigavi / čigav) Pi-ndn—-n(drugi / drug) Pi-npa—-a(čigava / čigav) Pi-npa—-n(druga / drug) Pi-npd—-a(čigavim / čigav) Pinpd—-n(drugim / drug) Pi-npg—-a(čigavih / čigav) Pi-npg—-n(drugih / drug) Pi-npi—-a(čigavimi / čigav) Pi-npi—-n(drugimi / drug) Pi-npl—-a(čigavih / čigav) Pi-npl—-n(drugih / drug) Pi-npn—-a(čigava / čigav) Pi-npn—-n(druga / drug) Pi-nsa—-a(čigavo / čigav) Pi-nsa—-n(drugo / drug) Pi-nsd—-a(čigavemu / čigav) Pinsd—-n(drugemu / drug) Pi-nsg—-a(čigavega / čigav) Pi-nsg—-n(drugega / drug) Pi-nsi—-a(čigavim / čigav) Pi-nsi—-n(drugim / drug) Pi-nsl—-a(čigavem / čigav) Pi-nsl—-n(drugem / drug) Pi-nsn—-a(čigavo / čigav) Pinsn—-n(drugo / drug) Pp1-da–n-n(naju / jaz) Pp1-da–y-n(naju / jaz) Pp1-dd–n-n(nama / jaz) Pp1-dd–y-n(nama / jaz) Pp1-dg–n-n(naju / jaz) Pp1-dg–y-n(naju / jaz) Pp1-di–n-n(nama / jaz) Pp1-dl–n-n(naju / jaz) Pp1-dn– n-n(midva / jaz) Pp1-pa–n-n(nas / jaz) Pp1-pa–y-n(nas / jaz) Pp1-pd–n-n(nam / jaz) Pp1-pd–y-n(nam / jaz) Pp1-pg–n-n(nas / jaz) Pp1-pg–y-n(nas / jaz) Pp1-pi–n-n(nami / jaz) Pp1-pl–n-n(nas / jaz) Pp1-pn–n-n(mi / jaz) Pp1-sa–n-n(mene / jaz) Pp1-sa–y-n(me / jaz) Pp1-sd–n-n(meni / jaz) Pp1-sd–y-n(mi / jaz) Pp1-sg–n-n(mene / jaz) Pp1-sg–y-n(me / jaz) Pp1-si–n-n(mano / jaz) Pp1-sl–n-n(meni / jaz) Pp1-sn–n-n(jaz / =) Pp2-da–n-n(vaju / ti) Pp2-da–y-n(vaju / ti) Pp2-dd–n-n(vama / ti) Pp2-dd–y-n(vama / ti) Pp2-dg–n-n(vaju / ti) Pp2-dg–y-n(vaju / ti) Pp2-di–n-n(vama / ti) Pp2-dl–n-n(vaju / ti) Pp2-dn–n-n(vidva / ti) Pp2-pa–n-n(vas / ti) Pp2-pa–y-n(vas / ti) Pp2-pd–n-n(vam / ti) Pp2-pd–y-n(vam / ti) Pp2-pg–n-n(vas / ti) Pp2-pg–y-n(vas / ti) Pp2-pi–n-n(vami / ti) Pp2-pl–n-n(vas / ti) Pp2-pn–n-n(vi / ti) Pp2-sa–n-n(tebe / ti) Pp2-sa–y-n(te / ti) Pp2-sd–n-n(tebi / ti) Pp2-sd–y-n(ti / ti) Pp2-sg–n-n(tebe / ti) Pp2-sg–y-n(te / ti) Pp2-si–n-n(tabo / ti) Pp2-sl–n-n(tebi / ti) Pp2-sn–n-n(ti / =) Pp3fda–n-n(njiju / on) Pp3fda–y-n(jih / on) Pp3fdd–n-n(njima / on) Pp3fdd–y-n(jima / on) Pp3fdg–n-n(njiju / on) Pp3fdg–y-n(jih / on) Pp3fdi–n-n(njima / on) Pp3fdl–n-n(njiju / on) Pp3fdn–n-n(onedve / on) Pp3fpa–n-n(nje / on) Pp3fpa–y-n(jih / on) Pp3fpd–n-n(njim / on) Pp3fpd–y-n(jim / on) Pp3fpg–n-n(njih / on) Pp3fpg–y-n(jih / on) Pp3fpi–n-n(njimi / on) Pp3fpl–n-n(njih / on) Pp3fpn–n-n(one / on) Pp3fsa–n-n(njo / on) Pp3fsa–y-n(jo / on) Pp3fsd–n-n(njej / on) Pp3fsd–y-n(ji / on) Pp3fsg–n-n(nje / on) Pp3fsg–y-n(je / on) Pp3fsi–n-n(njo / on) Pp3fsl–n-n(njej / on) Pp3fsn–n-n(ona / on) Pp3mda–n-n(njiju / on) Pp3mda–y-n(jih / on) Pp3mdd–n-n(njima / on) Pp3mdd–y-n(jima / on) Pp3mdg–n-n(njih / on) Pp3mdg–y-n(jih / on) Pp3mdi–n-n(njima / on) Pp3mdl– n-n(njiju / on) Pp3mdn–n-n(onadva / on) Pp3mpa–n-n(nje / on) Pp3mpa–y-n(jih / on) Pp3mpd–n-n(njim / on) Pp3mpd–y-n(jim / on) Pp3mpg–n-n(njih / on) Pp3mpg–y-n(jih / on) Pp3mpi–n-n(njimi / on) Pp3mpl–n-n(njih / on) Pp3mpn–n-n(oni / on) Pp3msa–n-n(njega / on) Pp3msa–y-n(ga / on) Pp3msd–n-n(njemu / on) Pp3msd– y-n(mu / on) Pp3msg–n-n(njega / on) Pp3msg–y-n(ga / on) Pp3msi–n-n(njim / on) Pp3msl–n-n(njem / on) Pp3msn–n-n(on / =) Pp3nda–n-n(njiju / on) Pp3nda–y-n(jih / on) Pp3ndd–n-n(njima / on) Pp3ndd–y-n(jima / on) Pp3ndg–n-n(njiju / on) Pp3ndg–y-n(jih / on) Pp3ndi–n-n(njima / on) Pp3ndl–n-n(njiju / on) Pp3ndn– n-n(onedve / on) Pp3npa–n-n(nje / on) Pp3npa–y-n(jih / on) Pp3npd–n-n(njim / on) Pp3npd–y-n(jim / on) Pp3npg–n-n(njih / on) Pp3npg–y-n(jih / on) Pp3npi–n-n(njimi / on) Pp3npl–n-n(njih / on) Pp3npn–n-n(ona / on) Pp3nsa–n-n(ono / on) Pp3nsa–y-n(ga / on) Pp3nsd–n-n(njemu / on) Pp3nsd–y-n(mu / on) Pp3nsg–n-n(njega / on) Pp3nsg–y-n(ga / on) Pp3nsi–n-n(njim / on) Pp3nsl–n-n(njem / on) Pp3nsn–n-n(ono / on) Pq——–a(koliko / =) Pq—a—-n(kaj / kaj) Pq—d—-n(čemu / kaj) Pq—g—-n(česa / kaj) Pq—i—-n(čim / kaj) Pq—l—-n(čem / kaj) Pq—n—-n(kaj / =) Pq-fda—-a(čigavi / čigav) Pq-fdd—-a(čigavima / čigav) Pq-fdg—-a(čigavih / čigav) Pq-fdi—-a(čigavima / čigav) Pq-fdl—-a(čigavih / čigav) Pq-fdn—-a(čigavi / čigav) Pq-fpa—-a(čigave / čigav) Pq-fpd—-a(čigavim / čigav) Pq-fpg—-a(čigavih / čigav) Pq-fpi—-a(čigavimi / čigav) Pq-fpl—-a(čigavih / čigav) Pq-fpn—-a(čigave / čigav) Pq-fsa—-a(čigavo / čigav) Pq-fsd—-a(čigavi / čigav) Pq-fsg—-a(čigave / čigav) Pqfsi—-a(čigavo / čigav) Pq-fsl—-a(čigavi / čigav) Pq-fsn—-a(čigava / čigav) Pq-mda—-a(čigava / čigav) Pq-mdd— -a(čigavima / čigav) Pq-mdg—-a(čigavih / čigav) Pq-mdi—-a(čigavima / čigav) Pq-mdl—-a(čigavih / čigav) Pqmdn—-a(čigava / čigav) Pq-mpa—-a(čigave / čigav) Pq-mpd—-a(čigavim / čigav) Pq-mpg—-a(čigavih / čigav) Pq-mpi—-a(čigavimi / čigav) Pq-mpl—-a(čigavih / čigav) Pq-mpn—-a(čigavi / čigav) Pq-msa—-a-n(čigav / čigav) Pq-msa—-a-y(čigavega / čigav) Pq-msd—-a(čigavemu / čigav) Pq-msg—-a(čigavega / čigav) Pq-msi—-a(čigavim / čigav) Pq-msl—-a(čigavem / čigav) Pq-msn—-a(čigav / =) Pq-nda—-a(čigavi / čigav) Pq-ndd—-a(čigavima / čigav) Pq-ndg—-a(čigavih / čigav) Pq-ndi—-a(čigavima / čigav) Pq-ndl—-a(čigavih / čigav) Pq-ndn—-a(čigavi / čigav) Pq-npa—-a(čigava / čigav) Pq-npd—-a(čigavim / čigav) Pq-npg—-a(čigavih / čigav) Pq-npi—-a(čigavimi / čigav) Pq-npl—-a(čigavih / čigav) Pq-npn—-a(čigava / čigav) Pq-nsa—-a(čigavo / čigav) Pq-nsd—-a(čigavemu / čigav) Pq-nsg—-a(čigavega / čigav) Pq-nsi—-a(čigavim / čigav) Pq-nsl—-a(čigavem / čigav) Pq-nsn—-a(čigavo / čigav) Pr——–a(čigar / =) Pr—a—-n(kar / kar) Pr—d—-n(čemur / kar) Pr—g—-n(česar / kar) Pr—i—- n(čimer / kar) Pr—l—-n(čemer / kar) Pr—n—-n(kar / =) Pr-fda—-a(kakršni / kakršen) Pr-fdd—-a(kakršnima / kakršen) Pr-fdg—-a(kakršnih / kakršen) Pr-fdi—-a(kakršnima / kakršen) Pr-fdl—-a(kakršnih / kakršen) Prfdn—-a(kakršni / kakršen) Pr-fpa—-a(kakršne / kakršen) Pr-fpd—-a(kakršnim / kakršen) Pr-fpg—-a(kakršnih / kakršen) Pr-fpi—-a(kakršnimi / kakršen) Pr-fpl—-a(kakršnih / kakršen) Pr-fpn—-a(kakršne / kakršen) Prfsa—-a(kakršno / kakršen) Pr-fsd—-a(kakršni / kakršen) Pr-fsg—-a(kakršne / kakršen) Pr-fsi—-a(kakršno / kakršen) Pr-fsl—-a(kakršni / kakršen) Pr-fsn—-a(kakršna / kakršen) Pr-mda—-a(kakršna / kakršen) Pr-mdd—- a(kakršnima / kakršen) Pr-mdg—-a(kakršnih / kakršen) Pr-mdi—-a(kakršnima / kakršen) Pr-mdl—-a(kakršnih / kakršen) Pr-mdn—-a(kakršna / kakršen) Pr-mpa—-a(kakršne / kakršen) Pr-mpd—-a(kakršnim / kakršen) Pr-mpg—-a(kakršnih / kakršen) Pr-mpi—-a(kakršnimi / kakršen) Pr-mpl—-a(kakršnih / kakršen) Pr-mpn—- a(kakršni / kakršen) Pr-msa—-a-n(kakršen / kakršen) Pr-msa—-a-y(kakršnega / kakršen) Pr-msd—-a(kakršnemu / kakršen) Pr-msg—-a(kakršnega / kakršen) Pr-msi—-a(kakršnim / kakršen) Pr-msl—-a(kakršnem / kakršen) Pr-msn—-a(kakršen / =) Pr-nda—-a(kakršni / kakršen) Pr-ndd—-a(kakršnima / kakršen) Pr-ndg—-a(kakršnih / kakršen) Pr-ndi—-a(kakršnima / kakršen) Pr-ndl—-a(kakršnih / kakršen) Pr-ndn—-a(kakršni / kakršen) Prnpa—-a(kakršna / kakršen) Pr-npd—-a(kakršnim / kakršen) Pr-npg—-a(kakršnih / kakršen) Pr-npi—-a(kakršnimi / kakršen) Pr-npl—-a(kakršnih / kakršen) Pr-npn—-a(kakršna / kakršen) Pr-nsa—-a(kakršno / kakršen) Pr-nsd— -a(kakršnemu / kakršen) Pr-nsg—-a(kakršnega / kakršen) Pr-nsi—-a(kakršnim / kakršen) Pr-nsl—-a(kakršnem / kakršen) Pr-nsn—-a(kakršno / kakršen) Ps1fdad—a(najini / najin) Ps1fdap—a(naši / naš) Ps1fdas—a(moji / moj) Ps1fddd—a(najinima / najin) Ps1fddp—a(našima / naš) Ps1fdds—a(mojima / moj) Ps1fdgd—a(najinih / najin) Ps1fdgp—a(naših / naš) Ps1fdgs—a(mojih / moj) Ps1fdid—a(najinima / najin) Ps1fdip—a(našima / naš) Ps1fdis—a(mojima / moj) Ps1fdld—a(najinih / najin) Ps1fdlp—a(naših / naš) Ps1fdls—a(mojih / moj) Ps1fdnd—a(najini / najin) Ps1fdnp—a(naši / naš) Ps1fdns—a(moji / moj) Ps1fpad—a(najine / najin) Ps1fpap— a(naše / naš) Ps1fpas—a(moje / moj) Ps1fpdd—a(najinim / najin) Ps1fpdp—a(našim / naš) Ps1fpds—a(mojim / moj) Ps1fpgd—a(najinih / najin) Ps1fpgp—a(naših / naš) Ps1fpgs—a(mojih / moj) Ps1fpid—a(najinimi / najin) Ps1fpip—a(našimi / naš) Ps1fpis—a(mojimi / moj) Ps1fpld—a(najinih / najin) Ps1fplp—a(naših / naš) Ps1fpls—a(mojih / moj) Ps1fpnd—a(najine / najin) Ps1fpnp—a(naše / naš) Ps1fpns—a(moje / moj) Ps1fsad— a(najino / najin) Ps1fsap—a(našo / naš) Ps1fsas—a(mojo / moj) Ps1fsdd—a(najini / najin) Ps1fsdp—a(naši / naš) Ps1fsds—a(moji / moj) Ps1fsgd—a(najine / najin) Ps1fsgp—a(naše / naš) Ps1fsgs—a(moje / moj) Ps1fsid— a(najino / najin) Ps1fsip—a(našo / naš) Ps1fsis—a(mojo / moj) Ps1fsld—a(najini / najin) Ps1fslp—a(naši / naš) Ps1fsls—a(moji / moj) Ps1fsnd—a(najina / najin) Ps1fsnp—a(naša / naš) Ps1fsns—a(moja / moj) Ps1mdad— a(najina / najin) Ps1mdap—a(naša / naš) Ps1mdas—a(moja / moj) Ps1mddd—a(najinima / najin) Ps1mddp— a(našima / naš) Ps1mdds—a(mojima / moj) Ps1mdgd—a(najinih / najin) Ps1mdgp—a(naših / naš) Ps1mdgs— a(mojih / moj) Ps1mdid—a(najinima / najin) Ps1mdip—a(našima / naš) Ps1mdis—a(mojima / moj) Ps1mdld— a(najinih / najin) Ps1mdlp—a(naših / naš) Ps1mdls—a(mojih / moj) Ps1mdnd—a(najina / najin) Ps1mdnp— a(naša / naš) Ps1mdns—a(moja / moj) Ps1mpad—a(najine / najin) Ps1mpap—a(naše / naš) Ps1mpas—a(moje / moj) Ps1mpdd—a(najinim / najin) Ps1mpdp—a(našim / naš) Ps1mpds—a(mojim / moj) Ps1mpgd—a(najinih / najin) Ps1mpgp—a(naših / naš) Ps1mpgs—a(mojih / moj) Ps1mpid—a(najinimi / najin) Ps1mpip—a(našimi / naš) Ps1mpis—a(mojimi / moj) Ps1mpld—a(najinih / najin) Ps1mplp—a(naših / naš) Ps1mpls—a(mojih / moj) Ps1mpnd—a(najini / najin) Ps1mpnp—a(naši / naš) Ps1mpns—a(moji / moj) Ps1msad—a-n(najin / najin) Ps1msad—a-y(najinega / najin) Ps1msap—a-n(naš / naš) Ps1msap—a-y(našega / naš) Ps1msas—a-n(moj / moj) Ps1msas—a-y(mojega / moj) Ps1msdd—a(najinemu / najin) Ps1msdp—a(našemu / naš) Ps1msds—a(mojemu / moj) Ps1msgd—a(najinega / najin) Ps1msgp—a(našega / naš) Ps1msgs—a(mojega / moj) Ps1msid—a(najinim / najin) Ps1msip—a(našim / naš) Ps1msis—a(mojim / moj) Ps1msld—a(najinem / najin) Ps1mslp—a(našem / naš) Ps1msls—a(mojem / moj) Ps1msnd—a(najin / =) Ps1msnp—a(naš / =) Ps1msns—a(moj / =) Ps1ndad— a(najini / najin) Ps1ndap—a(naši / naš) Ps1ndas—a(moji / moj) Ps1nddd—a(najinima / najin) Ps1nddp— a(našima / naš) Ps1ndds—a(mojima / moj) Ps1ndgd—a(najinih / najin) Ps1ndgp—a(naših / naš) Ps1ndgs— a(mojih / moj) Ps1ndid—a(najinima / najin) Ps1ndip—a(našima / naš) Ps1ndis—a(mojima / moj) Ps1ndld— a(najinih / najin) Ps1ndlp—a(naših / naš) Ps1ndls—a(mojih / moj) Ps1ndnd—a(najini / najin) Ps1ndnp— a(naši / naš) Ps1ndns—a(moji / moj) Ps1npad—a(najina / najin) Ps1npap—a(naša / naš) Ps1npas—a(moja / moj) Ps1npdd—a(najinim / najin) Ps1npdp—a(našim / naš) Ps1npds—a(mojim / moj) Ps1npgd—a(najinih / najin) Ps1npgp—a(naših / naš) Ps1npgs—a(mojih / moj) Ps1npid—a(najinimi / najin) Ps1npip—a(našimi / naš) Ps1npis—a(mojimi / moj) Ps1npld—a(najinih / najin) Ps1nplp—a(naših / naš) Ps1npls—a(mojih / moj) Ps1npnd—a(najina / najin) Ps1npnp—a(naša / naš) Ps1npns—a(moja / moj) Ps1nsad—a(najino / najin) Ps1nsap—a(naše / naš) Ps1nsas—a(moje / moj) Ps1nsdd—a(najinemu / najin) Ps1nsdp—a(našemu / naš) Ps1nsds—a(mojemu / moj) Ps1nsgd—a(najinega / najin) Ps1nsgp—a(našega / naš) Ps1nsgs—a(mojega / moj) Ps1nsid—a(najinim / najin) Ps1nsip—a(našim / naš) Ps1nsis—a(mojim / moj) Ps1nsld—a(najinem / najin) Ps1nslp—a(našem / naš) Ps1nsls—a(mojem / moj) Ps1nsnd—a(najino / najin) Ps1nsnp—a(naše / naš) Ps1nsns—a(moje / moj) Ps2fdad—a(vajini / vajin) Ps2fdap—a(vaši / vaš) Ps2fdas—a(tvoji / tvoj) Ps2fddd— a(vajinima / vajin) Ps2fddp—a(vašima / vaš) Ps2fdds—a(tvojima / tvoj) Ps2fdgd—a(vajinih / vajin) Ps2fdgp— a(vaših / vaš) Ps2fdgs—a(tvojih / tvoj) Ps2fdid—a(vajinima / vajin) Ps2fdip—a(vašima / vaš) Ps2fdis—a(tvojima / tvoj) Ps2fdld—a(vajinih / vajin) Ps2fdlp—a(vaših / vaš) Ps2fdls—a(tvojih / tvoj) Ps2fdnd—a(vajini / vajin) Ps2fdnp—a(vaši / vaš) Ps2fdns—a(tvoji / tvoj) Ps2fpad—a(vajine / vajin) Ps2fpap—a(vaše / vaš) Ps2fpas— a(tvoje / tvoj) Ps2fpdd—a(vajinim / vajin) Ps2fpdp—a(vašim / vaš) Ps2fpds—a(tvojim / tvoj) Ps2fpgd— a(vajinih / vajin) Ps2fpgp—a(vaših / vaš) Ps2fpgs—a(tvojih / tvoj) Ps2fpid—a(vajinimi / vajin) Ps2fpip— a(vašimi / vaš) Ps2fpis—a(tvojimi / tvoj) Ps2fpld—a(vajinih / vajin) Ps2fplp—a(vaših / vaš) Ps2fpls—a(tvojih / tvoj) Ps2fpnd—a(vajine / vajin) Ps2fpnp—a(vaše / vaš) Ps2fpns—a(tvoje / tvoj) Ps2fsad—a(vajino / vajin) Ps2fsap—a(vašo / vaš) Ps2fsas—a(tvojo / tvoj) Ps2fsdd—a(vajini / vajin) Ps2fsdp—a(vaši / vaš) Ps2fsds— a(tvoji / tvoj) Ps2fsgd—a(vajine / vajin) Ps2fsgp—a(vaše / vaš) Ps2fsgs—a(tvoje / tvoj) Ps2fsid—a(vajino / vajin) Ps2fsip—a(vašo / vaš) Ps2fsis—a(tvojo / tvoj) Ps2fsld—a(vajini / vajin) Ps2fslp—a(vaši / vaš) Ps2fsls— a(tvoji / tvoj) Ps2fsnd—a(vajina / vajin) Ps2fsnp—a(vaša / vaš) Ps2fsns—a(tvoja / tvoj) Ps2mdad—a(vajina / vajin) Ps2mdap—a(vaša / vaš) Ps2mdas—a(tvoja / tvoj) Ps2mddd—a(vajinima / vajin) Ps2mddp—a(vašima / vaš) Ps2mdds—a(tvojima / tvoj) Ps2mdgd—a(vajinih / vajin) Ps2mdgp—a(vaših / vaš) Ps2mdgs—a(tvojih / tvoj) Ps2mdid—a(vajinima / vajin) Ps2mdip—a(vašima / vaš) Ps2mdis—a(tvojima / tvoj) Ps2mdld—a(vajinih / vajin) Ps2mdlp—a(vaših / vaš) Ps2mdls—a(tvojih / tvoj) Ps2mdnd—a(vajina / vajin) Ps2mdnp—a(vaša / vaš) Ps2mdns—a(tvoja / tvoj) Ps2mpad—a(vajine / vajin) Ps2mpap—a(vaše / vaš) Ps2mpas—a(tvoje / tvoj) Ps2mpdd—a(vajinim / vajin) Ps2mpdp—a(vašim / vaš) Ps2mpds—a(tvojim / tvoj) Ps2mpgd—a(vajinih / vajin) Ps2mpgp—a(vaših / vaš) Ps2mpgs—a(tvojih / tvoj) Ps2mpid—a(vajinimi / vajin) Ps2mpip—a(vašimi / vaš) Ps2mpis—a(tvojimi / tvoj) Ps2mpld—a(vajinih / vajin) Ps2mplp—a(vaših / vaš) Ps2mpls—a(tvojih / tvoj) Ps2mpnd—a(vajini / vajin) Ps2mpnp—a(vaši / vaš) Ps2mpns—a(tvoji / tvoj) Ps2msad—a-n(vajin / vajin) Ps2msad—a-y(vajinega / vajin) Ps2msap—a-n(vaš / vaš) Ps2msap—a-y(vašega / vaš) Ps2msas—a-n(tvoj / tvoj) Ps2msas—a-y(tvojega / tvoj) Ps2msdd—a(vajinemu / vajin) Ps2msdp—a(vašemu / vaš) Ps2msds—a(tvojemu / tvoj) Ps2msgd—a(vajinega / vajin) Ps2msgp—a(vašega / vaš) Ps2msgs—a(tvojega / tvoj) Ps2msid—a(vajinim / vajin) Ps2msip—a(vašim / vaš) Ps2msis—a(tvojim / tvoj) Ps2msld—a(vajinem / vajin) Ps2mslp—a(vašem / vaš) Ps2msls—a(tvojem / tvoj) Ps2msnd—a(vajin / =) Ps2msnp—a(vaš / =) Ps2msns—a(tvoj / =) Ps2ndad— a(vajini / vajin) Ps2ndap—a(vaši / vaš) Ps2ndas—a(tvoji / tvoj) Ps2nddd—a(vajinima / vajin) Ps2nddp— a(vašima / vaš) Ps2ndds—a(tvojima / tvoj) Ps2ndgd—a(vajinih / vajin) Ps2ndgp—a(vaših / vaš) Ps2ndgs— a(tvojih / tvoj) Ps2ndid—a(vajinima / vajin) Ps2ndip—a(vašima / vaš) Ps2ndis—a(tvojima / tvoj) Ps2ndld— a(vajinih / vajin) Ps2ndlp—a(vaših / vaš) Ps2ndls—a(tvojih / tvoj) Ps2ndnd—a(vajini / vajin) Ps2ndnp— a(vaši / vaš) Ps2ndns—a(tvoji / tvoj) Ps2npad—a(vajina / vajin) Ps2npap—a(vaša / vaš) Ps2npas—a(tvoja / tvoj) Ps2npdd—a(vajinim / vajin) Ps2npdp—a(vašim / vaš) Ps2npds—a(tvojim / tvoj) Ps2npgd—a(vajinih / vajin) Ps2npgp—a(vaših / vaš) Ps2npgs—a(tvojih / tvoj) Ps2npid—a(vajinimi / vajin) Ps2npip—a(vašimi / vaš) Ps2npis—a(tvojimi / tvoj) Ps2npld—a(vajinih / vajin) Ps2nplp—a(vaših / vaš) Ps2npls—a(tvojih / tvoj) Ps2npnd—a(vajina / vajin) Ps2npnp—a(vaša / vaš) Ps2npns—a(tvoja / tvoj) Ps2nsad—a(vajino / vajin) Ps2nsap—a(vaše / vaš) Ps2nsas—a(tvoje / tvoj) Ps2nsdd—a(vajinemu / vajin) Ps2nsdp—a(vašemu / vaš) Ps2nsds—a(tvojemu / tvoj) Ps2nsgd—a(vajinega / vajin) Ps2nsgp—a(vašega / vaš) Ps2nsgs—a(tvojega / tvoj) Ps2nsid—a(vajinim / vajin) Ps2nsip—a(vašim / vaš) Ps2nsis—a(tvojim / tvoj) Ps2nsld—a(vajinem / vajin) Ps2nslp—a(vašem / vaš) Ps2nsls—a(tvojem / tvoj) Ps2nsnd—a(vajino / vajin) Ps2nsnp—a(vaše / vaš) Ps2nsns— a(tvoje / tvoj) Ps3fdad—a(njuni / njun) Ps3fdap—a(njihovi / njihov) Ps3fdasf–a(njeni / njen) Ps3fdasm– a(njegovi / njegov) Ps3fdasn–a(njegovi / njegov) Ps3fddd—a(njunima / njun) Ps3fddp—a(njihovima / njihov) Ps3fddsf–a(njenima / njen) Ps3fddsm–a(njegovima / njegov) Ps3fddsn–a(njegovima / njegov) Ps3fdgd—a(njunih / njun) Ps3fdgp—a(njihovih / njihov) Ps3fdgsf–a(njenih / njen) Ps3fdgsm–a(njegovih / njegov) Ps3fdgsn–a(njegovih / njegov) Ps3fdid—a(njunima / njun) Ps3fdip—a(njihovima / njihov) Ps3fdisf–a(njenima / njen) Ps3fdism– a(njegovima / njegov) Ps3fdisn–a(njegovima / njegov) Ps3fdld—a(njunih / njun) Ps3fdlp—a(njihovih / njihov) Ps3fdlsf–a(njenih / njen) Ps3fdlsm–a(njegovih / njegov) Ps3fdlsn–a(njegovih / njegov) Ps3fdnd—a(njuni / njun) Ps3fdnp—a(njihovi / njihov) Ps3fdnsf–a(njeni / njen) Ps3fdnsm–a(njegovi / njegov) Ps3fdnsn–a(njegovi / njegov) Ps3fpad—a(njune / njun) Ps3fpap—a(njihove / njihov) Ps3fpasf–a(njene / njen) Ps3fpasm–a(njegove / njegov) Ps3fpasn–a(njegove / njegov) Ps3fpdd—a(njunim / njun) Ps3fpdp—a(njihovim / njihov) Ps3fpdsf–a(njenim / njen) Ps3fpdsm–a(njegovim / njegov) Ps3fpdsn–a(njegovim / njegov) Ps3fpgd—a(njunih / njun) Ps3fpgp— a(njihovih / njihov) Ps3fpgsf–a(njenih / njen) Ps3fpgsm–a(njegovih / njegov) Ps3fpgsn–a(njegovih / njegov) Ps3fpid—a(njunimi / njun) Ps3fpip—a(njihovimi / njihov) Ps3fpisf–a(njenimi / njen) Ps3fpism–a(njegovimi / njegov) Ps3fpisn–a(njegovimi / njegov) Ps3fpld—a(njunih / njun) Ps3fplp—a(njihovih / njihov) Ps3fplsf–a(njenih / njen) Ps3fplsm–a(njegovih / njegov) Ps3fplsn–a(njegovih / njegov) Ps3fpnd—a(njune / njun) Ps3fpnp—a(njihove / njihov) Ps3fpnsf–a(njene / njen) Ps3fpnsm–a(njegove / njegov) Ps3fpnsn–a(njegove / njegov) Ps3fsad—a(njuno / njun) Ps3fsap—a(njihovo / njihov) Ps3fsasf–a(njeno / njen) Ps3fsasm–a(njegovo / njegov) Ps3fsasn–a(njegovo / njegov) Ps3fsdd—a(njuni / njun) Ps3fsdp—a(njihovi / njihov) Ps3fsdsf–a(njeni / njen) Ps3fsdsm–a(njegovi / njegov) Ps3fsdsn–a(njegovi / njegov) Ps3fsgd—a(njune / njun) Ps3fsgp—a(njihove / njihov) Ps3fsgsf–a(njene / njen) Ps3fsgsm–a(njegove / njegov) Ps3fsgsn–a(njegove / njegov) Ps3fsid—a(njuno / njun) Ps3fsip—a(njihovo / njihov) Ps3fsisf–a(njeno / njen) Ps3fsism–a(njegovo / njegov) Ps3fsisn–a(njegovo / njegov) Ps3fsld—a(njuni / njun) Ps3fslp—a(njihovi / njihov) Ps3fslsf–a(njeni / njen) Ps3fslsm–a(njegovi / njegov) Ps3fslsn–a(njegovi / njegov) Ps3fsnd—a(njuna / njun) Ps3fsnp—a(njihova / njihov) Ps3fsnsf–a(njena / njen) Ps3fsnsm–a(njegova / njegov) Ps3fsnsn–a(njegova / njegov) Ps3mdad—a(njuna / njun) Ps3mdap—a(njihova / njihov) Ps3mdasf–a(njena / njen) Ps3mdasm–a(njegova / njegov) Ps3mdasn–a(njegova / njegov) Ps3mddd—a(njunima / njun) Ps3mddp— a(njihovima / njihov) Ps3mddsf–a(njenima / njen) Ps3mddsm–a(njegovima / njegov) Ps3mddsn–a(njegovima / njegov) Ps3mdgd—a(njunih / njun) Ps3mdgp—a(njihovih / njihov) Ps3mdgsf–a(njenih / njen) Ps3mdgsm– a(njegovih / njegov) Ps3mdgsn–a(njegovih / njegov) Ps3mdid—a(njunima / njun) Ps3mdip—a(njihovima / njihov) Ps3mdisf–a(njenima / njen) Ps3mdism–a(njegovima / njegov) Ps3mdisn–a(njegovima / njegov) Ps3mdld— a(njunih / njun) Ps3mdlp—a(njihovih / njihov) Ps3mdlsf–a(njenih / njen) Ps3mdlsm–a(njegovih / njegov) Ps3mdlsn–a(njegovih / njegov) Ps3mdnd—a(njuna / njun) Ps3mdnp—a(njihova / njihov) Ps3mdnsf–a(njena / njen) Ps3mdnsm–a(njegova / njegov) Ps3mdnsn–a(njegova / njegov) Ps3mpad—a(njune / njun) Ps3mpap— a(njihove / njihov) Ps3mpasf–a(njene / njen) Ps3mpasm–a(njegove / njegov) Ps3mpasn–a(njegove / njegov) Ps3mpdd—a(njunim / njun) Ps3mpdp—a(njihovim / njihov) Ps3mpdsf–a(njenim / njen) Ps3mpdsm–a(njegovim / njegov) Ps3mpdsn–a(njegovim / njegov) Ps3mpgd—a(njunih / njun) Ps3mpgp—a(njihovih / njihov) Ps3mpgsf– a(njenih / njen) Ps3mpgsm–a(njegovih / njegov) Ps3mpgsn–a(njegovih / njegov) Ps3mpid—a(njunimi / njun) Ps3mpip—a(njihovimi / njihov) Ps3mpisf–a(njenimi / njen) Ps3mpism–a(njegovimi / njegov) Ps3mpisn–a(njegovimi / njegov) Ps3mpld—a(njunih / njun) Ps3mplp—a(njihovih / njihov) Ps3mplsf–a(njenih / njen) Ps3mplsm– a(njegovih / njegov) Ps3mplsn–a(njegovih / njegov) Ps3mpnd—a(njuni / njun) Ps3mpnp—a(njihovi / njihov) Ps3mpnsf–a(njeni / njen) Ps3mpnsm–a(njegovi / njegov) Ps3mpnsn–a(njegovi / njegov) Ps3msad—a-n(njun / njun) Ps3msad—a-y(njunega / njun) Ps3msap—a-n(njihov / njihov) Ps3msap—a-y(njihovega / njihov) Ps3msasf– a-n(njen / njen) Ps3msasf–a-y(njenega / njen) Ps3msasm–a-n(njegov / njegov) Ps3msasm–a-y(njegovega / njegov) Ps3msasn–a-n(njegov / njegov) Ps3msasn–a-y(njegovega / njegov) Ps3msdd—a(njunemu / njun) Ps3msdp— a(njihovemu / njihov) Ps3msdsf–a(njenemu / njen) Ps3msdsm–a(njegovemu / njegov) Ps3msdsn–a(njegovemu / njegov) Ps3msgd—a(njunega / njun) Ps3msgp—a(njihovega / njihov) Ps3msgsf–a(njenega / njen) Ps3msgsm– a(njegovega / njegov) Ps3msgsn–a(njegovega / njegov) Ps3msid—a(njunim / njun) Ps3msip—a(njihovim / njihov) Ps3msisf–a(njenim / njen) Ps3msism–a(njegovim / njegov) Ps3msisn–a(njegovim / njegov) Ps3msld—a(njunem / njun) Ps3mslp—a(njihovem / njihov) Ps3mslsf–a(njenem / njen) Ps3mslsm–a(njegovem / njegov) Ps3mslsn– a(njegovem / njegov) Ps3msnd—a(njun / =) Ps3msnp—a(njihov / =) Ps3msnsf–a(njen / =) Ps3msnsm–a(njegov / =) Ps3msnsn–a(njegov / =) Ps3ndad—a(njuni / njun) Ps3ndap—a(njihovi / njihov) Ps3ndasf–a(njeni / njen) Ps3ndasm–a(njegovi / njegov) Ps3ndasn–a(njegovi / njegov) Ps3nddd—a(njunima / njun) Ps3nddp—a(njihovima / njihov) Ps3nddsf–a(njenima / njen) Ps3nddsm–a(njegovima / njegov) Ps3nddsn–a(njegovima / njegov) Ps3ndgd— a(njunih / njun) Ps3ndgp—a(njihovih / njihov) Ps3ndgsf–a(njenih / njen) Ps3ndgsm–a(njegovih / njegov) Ps3ndgsn– a(njegovih / njegov) Ps3ndid—a(njunima / njun) Ps3ndip—a(njihovima / njihov) Ps3ndisf–a(njenima / njen) Ps3ndism–a(njegovima / njegov) Ps3ndisn–a(njegovima / njegov) Ps3ndld—a(njunih / njun) Ps3ndlp—a(njihovih / njihov) Ps3ndlsf–a(njenih / njen) Ps3ndlsm–a(njegovih / njegov) Ps3ndlsn–a(njegovih / njegov) Ps3ndnd— a(njuni / njun) Ps3ndnp—a(njihovi / njihov) Ps3ndnsf–a(njeni / njen) Ps3ndnsm–a(njegovi / njegov) Ps3ndnsn– a(njegovi / njegov) Ps3npad—a(njuna / njun) Ps3npap—a(njihova / njihov) Ps3npasf–a(njena / njen) Ps3npasm– a(njegova / njegov) Ps3npasn–a(njegova / njegov) Ps3npdd—a(njunim / njun) Ps3npdp—a(njihovim / njihov) Ps3npdsf–a(njenim / njen) Ps3npdsm–a(njegovim / njegov) Ps3npdsn–a(njegovim / njegov) Ps3npgd—a(njunih / njun) Ps3npgp—a(njihovih / njihov) Ps3npgsf–a(njenih / njen) Ps3npgsm–a(njegovih / njegov) Ps3npgsn– a(njegovih / njegov) Ps3npid—a(njunimi / njun) Ps3npip—a(njihovimi / njihov) Ps3npisf–a(njenimi / njen) Ps3npism–a(njegovimi / njegov) Ps3npisn–a(njegovimi / njegov) Ps3npld—a(njunih / njun) Ps3nplp—a(njihovih / njihov) Ps3nplsf–a(njenih / njen) Ps3nplsm–a(njegovih / njegov) Ps3nplsn–a(njegovih / njegov) Ps3npnd— a(njuna / njun) Ps3npnp—a(njihova / njihov) Ps3npnsf–a(njena / njen) Ps3npnsm–a(njegova / njegov) Ps3npnsn– a(njegova / njegov) Ps3nsad—a(njuno / njun) Ps3nsap—a(njihovo / njihov) Ps3nsasf–a(njeno / njen) Ps3nsasm– a(njegovo / njegov) Ps3nsasn–a(njegovo / njegov) Ps3nsdd—a(njunemu / njun) Ps3nsdp—a(njihovemu / njihov) Ps3nsdsf–a(njenemu / njen) Ps3nsdsm–a(njegovemu / njegov) Ps3nsdsn–a(njegovemu / njegov) Ps3nsgd— a(njunega / njun) Ps3nsgp—a(njihovega / njihov) Ps3nsgsf–a(njenega / njen) Ps3nsgsm–a(njegovega / njegov) Ps3nsgsn–a(njegovega / njegov) Ps3nsid—a(njunim / njun) Ps3nsip—a(njihovim / njihov) Ps3nsisf–a(njenim / njen) Ps3nsism–a(njegovim / njegov) Ps3nsisn–a(njegovim / njegov) Ps3nsld—a(njunem / njun) Ps3nslp— a(njihovem / njihov) Ps3nslsf–a(njenem / njen) Ps3nslsm–a(njegovem / njegov) Ps3nslsn–a(njegovem / njegov) Ps3nsnd—a(njuno / njun) Ps3nsnp—a(njihovo / njihov) Ps3nsnsf–a(njeno / njen) Ps3nsnsm–a(njegovo / njegov) Ps3nsnsn–a(njegovo / njegov) Px——y(se / =) Px—a–npn(sebe / se) Px—a–ypn(se / se) Px—d–npn(sebi / se) Px—d–ypn(si / se) Px—g–npn(sebe / se) Px—g–ypn(se / se) Px—i–npn(sabo / se) Px—l–npn(sebi / se) Px-fda—sa(svoji / svoj) Px-fdd—sa(svojima / svoj) Px-fdg—sa(svojih / svoj) Px-fdi—sa(svojima / svoj) Px-fdl—sa(svojih / svoj) Px-fdn—sa(svoji / svoj) Px-fpa—sa(svoje / svoj) Px-fpd—sa(svojim / svoj) Px-fpg— sa(svojih / svoj) Px-fpi—sa(svojimi / svoj) Px-fpl—sa(svojih / svoj) Px-fpn—sa(svoje / svoj) Px-fsa—sa(svojo / svoj) Px-fsd—sa(svoji / svoj) Px-fsg—sa(svoje / svoj) Px-fsi—sa(svojo / svoj) Px-fsl—sa(svoji / svoj) Pxfsn—sa(svoja / svoj) Px-mda—sa(svoja / svoj) Px-mdd—sa(svojima / svoj) Px-mdg—sa(svojih / svoj) Pxmdi—sa(svojima / svoj) Px-mdl—sa(svojih / svoj) Px-mdn—sa(svoja / svoj) Px-mpa—sa(svoje / svoj) Pxmpd—sa(svojim / svoj) Px-mpg—sa(svojih / svoj) Px-mpi—sa(svojimi / svoj) Px-mpl—sa(svojih / svoj) Pxmpn—sa(svoji / svoj) Px-msa—sa-n(svoj / svoj) Px-msa—sa-y(svojega / svoj) Px-msd—sa(svojemu / svoj) Px-msg—sa(svojega / svoj) Px-msi—sa(svojim / svoj) Px-msl—sa(svojem / svoj) Px-msn—sa(svoj / =) Pxnda—sa(svoji / svoj) Px-ndd—sa(svojima / svoj) Px-ndg—sa(svojih / svoj) Px-ndi—sa(svojima / svoj) Pxndl—sa(svojih / svoj) Px-ndn—sa(svoji / svoj) Px-npa—sa(svoja / svoj) Px-npd—sa(svojim / svoj) Px-npg— sa(svojih / svoj) Px-npi—sa(svojimi / svoj) Px-npl—sa(svojih / svoj) Px-npn—sa(svoja / svoj) Px-nsa—sa(svoje / svoj) Px-nsd—sa(svojemu / svoj) Px-nsg—sa(svojega / svoj) Px-nsi—sa(svojim / svoj) Px-nsl—sa(svojem / svoj) Px-nsn—sa(svoje / svoj) Pz——–a(nič / =) Pz—a—-n(nič / nič) Pz—d—-n(ničemur / nič) Pz—g— -n(ničesar / nič) Pz—i—-n(ničimer / nič) Pz—l—-n(ničemer / nič) Pz—n—-n(nič / =) Pz-fda—-a(nikakršni / nikakršen) Pz-fdd—-a(nikakršnima / nikakršen) Pz-fdg—-a(nikakršnih / nikakršen) Pz-fdi—-a(nikakršnima / nikakršen) Pz-fdl—-a(nikakršnih / nikakršen) Pz-fdn—-a(nikakršni / nikakršen) Pz-fpa—-a(nikakršne / nikakršen) Pz-fpd—-a(nikakršnim / nikakršen) Pz-fpg—-a(nikakršnih / nikakršen) Pz-fpi—-a(nikakršnimi / nikakršen) Pzfpl—-a(nikakršnih / nikakršen) Pz-fpn—-a(nikakršne / nikakršen) Pz-fsa—-a(nikakršno / nikakršen) Pz-fsd—- a(nikakršni / nikakršen) Pz-fsg—-a(nikakršne / nikakršen) Pz-fsi—-a(nikakršno / nikakršen) Pz-fsl—-a(nikakršni / nikakršen) Pz-fsn—-a(nikakršna / nikakršen) Pz-mda—-a(nikakršna / nikakršen) Pz-mdd—-a(nikakršnima / nikakršen) Pz-mdg—-a(nikakršnih / nikakršen) Pz-mdi—-a(nikakršnima / nikakršen) Pz-mdl—-a(nikakršnih / nikakršen) Pz-mdn—-a(nikakršna / nikakršen) Pz-mpa—-a(nikakršne / nikakršen) Pz-mpd—-a(nikakršnim / nikakršen) Pz-mpg—-a(nikakršnih / nikakršen) Pz-mpi—-a(nikakršnimi / nikakršen) Pz-mpl—-a(nikakršnih / nikakršen) Pz-mpn—-a(nikakršni / nikakršen) Pz-msa—-a-n(nikakršen / nikakršen) Pz-msa—-a-y(nikakršnega / nikakršen) Pz-msd—-a(nikakršnemu / nikakršen) Pz-msg—-a(nikakršnega / nikakršen) Pz-msi—-a(nikakršnim / nikakršen) Pz-msl—-a(nikakršnem / nikakršen) Pz-msn—-a(nikakršen / =) Pz-nda—-a(nikakršni / nikakršen) Pz-ndd—-a(nikakršnima / nikakršen) Pz-ndg—-a(nikakršnih / nikakršen) Pz-ndi—-a(nikakršnima / nikakršen) Pz-ndl—-a(nikakršnih / nikakršen) Pz-ndn—-a(nikakršni / nikakršen) Pz-npa—-a(nikakršna / nikakršen) Pznpd—-a(nikakršnim / nikakršen) Pz-npg—-a(nikakršnih / nikakršen) Pz-npi—-a(nikakršnimi / nikakršen) Pznpl—-a(nikakršnih / nikakršen) Pz-npn—-a(nikakršna / nikakršen) Pz-nsa—-a(nikakršno / nikakršen) Pz-nsd— -a(nikakršnemu / nikakršen) Pz-nsg—-a(nikakršnega / nikakršen) Pz-nsi—-a(nikakršnim / nikakršen) Pz-nsl—- a(nikakršnem / nikakršen) Pz-nsn—-a(nikakršno / nikakršen)

6.8 Adposition (S) // Predlog (D)

= ============== ============== ==== ============== ================== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl             C-sl
= ============== ============== ==== ============== ================== ====
1 Type           preposition    p    vrsta          -                  p
- -------------- -------------- ---- -------------- ------------------ ----
2 Formation      simple         s    sestavljenost  enostaven          e
                 compound       c                   priponski          p
******************************* **** ************** ****************** ****
3 Case                               sklon
  (req.by prep.) genitive       g   (ki ga zahteva) rodilnik           r
                 dative         d                   dajalnik           d
                 accusative     a                   tožilnik           t
                 locative       l                   mestnik            m
                 instrumental   i                   orodnik            o
- -------------- -------------- ---- -------------- ------------------ ----
4 Clitic                        -
=================================

Notes:

• Slovene has prepositions only, with some arguable exceptions (e.g. ’navkljub’ / in spite of), which can be pre- or postpositions. All Adpositions in the lexicon are ’prepositions’, so the name of the category is directly traslated as ’Preposition’, and the name of the Type as ’-’.

• Although the MSDs distinguish the category ’Formation’, all the prepositions in the lexicon are ’simple’.

Combinations

*** **** **** ---- =========================================================
PoS Type Form  Case Examples
*** **** **** ---- =========================================================
 S    p    c    -   zase, nase, nanj, nanjo, nanju, ...
 S    p    s    g   brez, izpred, od, ...
 S    p    s    d   k, navzlic, ...
 S    p    s    a   s, po, skozi, ...
 S    p    s    l   po, na, v, ...
 S    p    s    i   pod, s, med, ...
*** **** **** ---- =========================================================

Lexicon

6 MSDs:

Spc(čezenj / čezenj) Spsa(čez / =) Spsd(h / =) Spsg(ž / =) Spsi(ž / =) Spsl(na / =)

6.9 Conjunction (C) // Veznik (V)

= ============== ============== ==== ============== ============== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl         C-sl
= ============== ============== ==== ============== ============== ====
1 Type           coordinating   c    vrsta          priredni       p
                 subordinating  s                   podredni       d
******************************* **** ************** ************** ****
2 Formation      simple         s    oblika         enobesedni     e
                 compound       c                   večbesedni     v
- -------------- -------------- ---- -------------- -------------- ----
3 Coord_Type                    -
- -------------- -------------- -
4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- -
6 Number                        -
- -------------- -------------- -
7 Person                        -
=================================

Notes:

• According to Slov. grammar, some conjunctions are ’two-part’, but these can often be either single or two-part conjunctions with identical first and second conjunct, e.g. ’ali ... ali’; such conunctions are treated as ordinary, ’one-part’ conjunctions, i.e. as Formation=simple.

• Conjunctions are also classified into ’multi-word’ conjunctions, e.g. ’kljub temu da’ / in spite of. These conjuctions constitute one lexical ’word’; should be merged by the text tokeniser e.g.: kljub_temu_da and marked as Formation=compound.

•

Combinations

*** **** ---- ===============================================================
PoS Type Form Examples
*** **** ---- ===============================================================
 C    c    s   in, ali
 C    s    s   da, ki, ali
 C    s    c   medtem_ko
*** **** ====================================================================

Lexicon

3 MSDs:

Ccs(a/=) Csc(kljub_temu_da/=) Css(če/=)

6.10 Particle (Q) // Členek (L)

= ============== ============== ====
P ATT-en         VAL-en         C-en
= ============== ============== ====
1 Type                          -
- -------------- -------------- ----
2 Formation                     -
- -------------- -------------- ----
3 Clitic                        -
====================================

Notes:

• nôtes sem izgubil.

Lexicon

1 MSD: Q(še / =)

6.11 Interjection (I) // Medmet (M)

= ============== ============== ====
P ATT-en         VAL-en         C-en
= ============== ============== ====
1 Type                          -
- -------------- -------------- ----
2 Formation                     -
====================================

Lexicon

1 MSD: I(ah / =)

6.12 Numeral (M) // Števnik (K)

= ============== ============== ==== ============== ================== ====
P ATT-en         VAL-en         C-en ATT-sl         VAL-sl             C-sl
= ============== ============== ==== ============== ================== ====
1 Type           cardinal       c    vrsta          glavni             g
                 ordinal        o                   vrstilni           v
                 multiple       m                   množilni           m
                 special        s                   drugi              d
- -------------- -------------- ---- -------------- ------------------ ----
2 Gender         masculine      m    spol           moški              m
                 feminine       f                   ženski             z
                 neuter         n                   srednji            s
- -------------- -------------- ---- -------------- ------------------ ----
3 Number         singular       s    število        ednina             e
                 plural         p                   množina            m
                 dual           d                   dvojina            d
- -------------- -------------- ---- -------------- ------------------ ----
4 Case           nominative     n    sklon          imenovalnik        i
                 genitive       g                   rodilnik           r
                 dative         d                   dajalnik           d
                 accusative     a                   tožilnik           t
                 locative       l                   mestnik            m
                 instrumental   i                   orodnik            o
******************************* **** ************** ****************** ****
5 Form           digit          d    zapis          arabski            a
                 roman          r                   rimski             b
                 letter         l                   besedni            r
- -------------- -------------- ---- -------------- ------------------ ----
6 Definiteness   no             n    določnost      ne                 n
                 yes            y                   da                 d
- -------------- -------------- ---- -------------- ------------------ ----
7 Clitic                        -                                     -
- -------------- -------------- ---- -------------- ------------------ ----
8 Class                         -                                     -
- -------------- -------------- ---- -------------- ------------------ ----
9 Animate        no             n    živost         ne                 n
                 yes            y                   da                 d
- -------------- -------------- ---- -------------- ------------------ ----
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================

Notes:

• Numerals in Slovene can function as nouns, adjectives or adverbs, and are in grammars described as subtypes of these categories. The above classification runs counter to the established practice and is missing an important syntactic distinction.

• Definiteness (-i ending = yes, -0 = no) is marked only on the multiple and special Numerals in the nominative and accusative masculine singular. See also Def note on Adjective.

• Animacy is marked only (-ega ending = yes, else = no) on the Numerals masculine singular accusative (M*msal*–[ny]). M[ms]msal have three msa forms: no-Definiteness, +Animate; -Definiteness, -Animate; +Definiteness, -Animate. See also Anim note on Adjective.

Combinations

*** **** **** **** **** ---- --- --- --- ---- ===================================
PoS Type Gend Numb Case Form Def Clt Cls Anm  Examples
*** **** **** **** **** ---- --- --- --- ---- ===================================
M    -   -    -     -    d   -    -   -  -    1984, ...
M    -   -    -     -    r   -    -   -  -    MCMXXCIV, ...
M    c   m    s     a    l   -    -   -  n    en
M    c   m    s     a    l   -    -   -  y    enega
M    c   m    s  [ngdli] l   -    -   -  -    eden, en, enemu, enemu, enim
M    c   m   [pd][ngdali]l   -    -   -  -    trije, treh, ..., dva,...
M    c  [fs][spd][ngdali]l   -    -   -  -    ena, eno,...,tri, treh,...,dve,..
M    o   m    s     a    l   -    -   -  n    prvi
M    o   m    s     a    l   -    -   -  y    prvega
M    o   m    s  [ngdli] l   -    -   -  -    prvi, prvemu, ...
M    o   m   [pd][ngdali]l   -    -   -  -    tretji, tretjih,...,drugi,...
M    o  [fs][spd][ngdali]l   -    -   -  -    prva, prvo,...,treh,...
M    m   m    s     n    l  [ny]  -   -  n    dvojen, dvojni
M    m   m    s     a    l  [ny]  -   -  n    dvojen, dvojni
M    m   m    s     a    l   -    -   -  y    dvojnega
M    m   m    s   [gdli] l   -    -   -  -    dvojen, dvejnega, dvojnemu,...
M    m   m   [pd][ngdali]l   -    -   -  -    dvojni, dvojna, dvojnih, ...
M    m  [fs][spd][ngdali]l   -    -   -  -    dvojna, dvojno,..., dvojne,...
M    s   m    s     n    l  [ny]  -   -  n    stoter, stoteri
M    s   m    s     a    l  [ny]  -   -  n    stoter, stoteri
M    s   m    s     a    l   -    -   -  y    stoterega
M    s   m    s   [gdli] l   -    -   -  -    stoterega, stoteremu,...
M    s   m   [pd][ngdali]l   -    -   -  -    stoteri, stotera,...
M    s  [fs][spd][ngdali]l   -    -   -  -    stotera, stotero, ..., stoterih
*** **** **** **** **** ---- --- --- --- ---- ===================================

Lexicon

225 MSDs:

M—-r(C / =) Mcfdal(dve / dva) Mcfddl(dvema / dva) Mcfdgl(dveh / dva) Mcfdil(dvema / dva) Mcfdll(dveh / dva) Mcfdnl(dve / dva) Mcfpal(šest / šest) Mcfpdl(šestdesetim / šestdeset) Mcfpgl(šestdesetih / šestdeset) Mcfpil(šestdesetimi / šestdeset) Mcfpll(šestdesetih / šestdeset) Mcfpnl(šest / šest) Mcfsal(eno / en) Mcfsdl(eni / en) Mcfsgl(ene / en) Mcfsil(eno / en) Mcfsll(eni / en) Mcfsnl(ena / en) Mcmdal(dva / dva) Mcmddl(dvema / dva) Mcmdgl(dveh / dva) Mcmdil(dvema / dva) Mcmdll(dveh / dva) Mcmdnl(dva / =) Mcmpal(šest / šest) Mcmpdl(šestdesetim / šestdeset) Mcmpgl(šestdesetih / šestdeset) Mcmpil(šestdesetimi / šestdeset) Mcmpll(šestdesetih / šestdeset) Mcmpnl(šest / =) Mcmsal—n(en / en) Mcmsal—y(enega / en) Mcmsdl(enemu / en) Mcmsgl(enega / en) Mcmsil(enim / en) Mcmsll(enem / en) Mcmsnl(eden / =) Mcndal(dve / dva) Mcnddl(dvema / dva) Mcndgl(dveh / dva) Mcndil(dvema / dva) Mcndll(dveh / dva) Mcndnl(dve / dva) Mcnpal(šest / šest) Mcnpdl(šestdesetim / šestdeset) Mcnpgl(šestdesetih / šestdeset) Mcnpil(šestdesetimi / šestdeset) Mcnpll(šestdesetih / šestdeset) Mcnpnl(šest / šest) Mcnsal(eno / en) Mcnsdl(enemu / en) Mcnsgl(enega / en) Mcnsil(enim / en) Mcnsll(enem / en) Mcnsnl(eno / en) Mmfdal(dvojni / dvojen) Mmfddl(dvojnima / dvojen) Mmfdgl(dvojnih / dvojen) Mmfdil(dvojnima / dvojen) Mmfdll(dvojnih / dvojen) Mmfdnl(dvojni / dvojen) Mmfpal(dvojne / dvojen) Mmfpdl(dvojnim / dvojen) Mmfpgl(dvojnih / dvojen) Mmfpil(dvojnimi / dvojen) Mmfpll(dvojnih / dvojen) Mmfpnl(dvojne / dvojen) Mmfsal(dvojno / dvojen) Mmfsdl(dvojni / dvojen) Mmfsgl(dvojne / dvojen) Mmfsil(dvojno / dvojen) Mmfsll(dvojni / dvojen) Mmfsnl(dvojna / dvojen) Mmmdal(dvojna / dvojen) Mmmddl(dvojnima / dvojen) Mmmdgl(dvojnih / dvojen) Mmmdil(dvojnima / dvojen) Mmmdll(dvojnih / dvojen) Mmmdnl(dvojna / dvojen) Mmmpal(dvojne / dvojen) Mmmpdl(dvojnim / dvojen) Mmmpgl(dvojnih / dvojen) Mmmpil(dvojnimi / dvojen) Mmmpll(dvojnih / dvojen) Mmmpnl(dvojni / dvojen) Mmmsal—y(dvojnega / dvojen) Mmmsaln–n(dvojen / dvojen) Mmmsaly–n(dvojni / dvojen) Mmmsdl(dvojnemu / dvojen) Mmmsgl(dvojnega / dvojen) Mmmsil(dvojnim / dvojen) Mmmsll(dvojnem / dvojen) Mmmsnln(dvojen / =) Mmmsnly(dvojni / dvojen) Mmndal(dvojni / dvojen) Mmnddl(dvojnima / dvojen) Mmndgl(dvojnih / dvojen) Mmndil(dvojnima / dvojen) Mmndll(dvojnih / dvojen) Mmndnl(dvojni / dvojen) Mmnpal(dvojna / dvojen) Mmnpdl(dvojnim / dvojen) Mmnpgl(dvojnih / dvojen) Mmnpil(dvojnimi / dvojen) Mmnpll(dvojnih / dvojen) Mmnpnl(dvojna / dvojen) Mmnsal(dvojno / dvojen) Mmnsdl(dvojnemu / dvojen) Mmnsgl(dvojnega / dvojen) Mmnsil(dvojnim / dvojen) Mmnsll(dvojnem / dvojen) Mmnsnl(dvojno / dvojen) Mofdal(četrti / četrti) Mofddl(četrtima / četrti) Mofdgl(četrtih / četrti) Mofdil(četrtima / četrti) Mofdll(četrtih / četrti) Mofdnl(četrti / četrti) Mofpal(četrte / četrti) Mofpdl(četrtim / četrti) Mofpgl(četrtih / četrti) Mofpil(četrtimi / četrti) Mofpll(četrtih / četrti) Mofpnl(četrte / četrti) Mofsal(četrto / četrti) Mofsdl(četrti / četrti) Mofsgl(četrte / četrti) Mofsil(četrto / četrti) Mofsll(četrti / četrti) Mofsnl(četrta / četrti) Momdal(četrta / četrti) Momddl(četrtima / četrti) Momdgl(četrtih / četrti) Momdil(četrtima / četrti) Momdll(četrtih / četrti) Momdnl(četrta / četrti) Mompal(četrte / četrti) Mompdl(četrtim / četrti) Mompgl(četrtih / četrti) Mompil(četrtimi / četrti) Mompll(četrtih / četrti) Mompnl(četrti / četrti) Momsal—n(četrti / četrti) Momsal—y(četrtega / četrti) Momsdl(četrtemu / četrti) Momsgl(četrtega / četrti) Momsil(četrtim / četrti) Momsll(četrtem / četrti) Momsnl(četrti / =) Mondal(četrti / četrti) Monddl(četrtima / četrti) Mondgl(četrtih / četrti) Mondil(četrtima / četrti) Mondll(četrtih / četrti) Mondnl(četrti / četrti) Monpal(četrta / četrti) Monpdl(četrtim / četrti) Monpgl(četrtih / četrti) Monpil(četrtimi / četrti) Monpll(četrtih / četrti) Monpnl(četrta / četrti) Monsal(četrto / četrti) Monsdl(četrtemu / četrti) Monsgl(četrtega / četrti) Monsil(četrtim / četrti) Monsll(četrtem / četrti) Monsnl(četrto / četrti) Msfdal(deseteri / deseter) Msfddl(deseterima / deseter) Msfdgl(deseterih / deseter) Msfdil(deseterima / deseter) Msfdll(deseterih / deseter) Msfdnl(deseteri / deseter) Msfpal(desetere / deseter) Msfpdl(deseterim / deseter) Msfpgl(deseterih / deseter) Msfpil(deseterimi / deseter) Msfpll(deseterih / deseter) Msfpnl(desetere / deseter) Msfsal(desetero / deseter) Msfsdl(deseteri / deseter) Msfsgl(desetere / deseter) Msfsil(desetero / deseter) Msfsll(deseteri / deseter) Msfsnl(desetera / deseter) Msmdal(desetera / deseter) Msmddl(deseterima / deseter) Msmdgl(deseterih / deseter) Msmdil(deseterima / deseter) Msmdll(deseterih / deseter) Msmdnl(desetera / deseter) Msmpal(desetere / deseter) Msmpdl(deseterim / deseter) Msmpgl(deseterih / deseter) Msmpil(deseterimi / deseter) Msmpll(deseterih / deseter) Msmpnl(deseteri / deseter) Msmsal—y(deseterega / deseter) Msmsaln–n(deseter / deseter) Msmsaly–n(dvakratni / dvakraten) Msmsdl(deseteremu / deseter) Msmsgl(deseterega / deseter) Msmsil(deseterim / deseter) Msmsll(deseterem / deseter) Msmsnln(deseter / =) Msmsnly(dvakratni / dvakraten) Msndal(deseteri / deseter) Msnddl(deseterima / deseter) Msndgl(deseterih / deseter) Msndil(deseterima / deseter) Msndll(deseterih / deseter) Msndnl(deseteri / deseter) Msnpal(desetera / deseter) Msnpdl(deseterim / deseter) Msnpgl(deseterih / deseter) Msnpil(deseterimi / deseter) Msnpll(deseterih / deseter) Msnpnl(desetera / deseter) Msnsal(desetero / deseter) Msnsdl(deseteremu / deseter) Msnsgl(deseterega / deseter) Msnsil(deseterim / deseter) Msnsll(deseterem / deseter) Msnsnl(desetero / deseter)

6.13 Abbreviation (Y) // Okrajšava (O)

= ============== ============== ====
P ATT-en         VAL-en         C-en
= ============== ============== ====
1 Syntactic_Type                -
- -------------- -------------- ----
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- ----
4 Case                          -
- -------------- -------------- ----
5 Definiteness                  -
====================================

Notes:

• Sem sodijo tako krajšave kot kratice.

Lexicon

1 MSD: Y(št. / =)

6.14 Residual (X) // Neuvrščeno (N)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =

Lexicon

1 MSD: Generation X

Application to Czech

Authors: Vladimír Petkevič, Faculty of Philosophy, Charles University, Prague.

Acknowledgements: The most appreciated base used for the elaboration of the application was constituted by morphosyntactic tables for Czech elaborated by Hana Skoumalová. For all possible errors solely the author is to blame.

7.1 Noun (N)

1.1. Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Czech term
= ============== ============== = ================= =========================
1 Type           common         c kniha             obecné jméno
                 proper         p Petr              vlastní jméno
- -------------- -------------- - ----------------- -------------------------
2 Gender         masculine      m otec              masculinum
                 feminine       f kniha             femininum
                 neuter         n slunce            neutrum
- -------------- -------------- - ----------------- -------------------------
3 Number         singular       s kniha             singulár
                 plural         p knihy             plurál
                 dual           d rukama            duál
- -------------- -------------- - ----------------- -------------------------
4 Case           nominative     n kniha             nominativ
                 genitive       g knihy             genitiv
                 dative         d knize             dativ
                 accusative     a knihu             akuzativ
                 vocative       v kniho!            vokativ
                 locative       l knize             lokál
                 instrumental   i knihou            instrumentál
* ************** ************** * ----------------- -------------------------

Multext-East / Concede  Morphosyntactic specifications V.2— Czech  107

5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -
- -------------- -------------- - ----------------- -------------------------
7 Animate   l.s. no             n hrad              neživotné
            l.s. yes            y otec              životné
- -------------- -------------- - ----------------- -------------------------
8 Owner_Number                  -
- -------------- -------------- -
9 Owner_Person                  -
- -------------- -------------- -
10Owned_Number                  -
=================================
Notes:
1. Czech  distinguishes masculine animate  (Animate=yes above)
   and masculine inanimate  (Animate=no) Gender. Masculine inani-
   mate Nouns always have the same form in the nominative and ac-
   cusative Case,  whereas masculine animate  Nouns have predomi-
   nantly the same form in the genitive and accusative Case. Mas-
   culine animate  Nouns and masculine inanimate  Nouns differ in
   accusative singular and nominative plural only.
2. The dual Number manifests itself only in the instrumental Case
   of several Nouns denoting dual parts of the human body.
3. Czech distinguishes 7 cases, the locative Case being obligato-
   rily prepositional. With most paradigms, the form of the voca-
   tive differs from that of  the nominative in singular; in plu-
   ral the form of the vocative is identical to that of the nomi-
   native for all the paradigms.
4. Verbal nouns are classified as Nouns.
1.2 Combinations
*** **** **** **** **** **** =============================================
PoS Type Gend Numb Case Anim Examples
*** **** **** **** **** **** =============================================
 N   p    m   [sp] any   y   Pavel, Pavlové
 N   p    m   [sp] any   n   Hradec, Hradce
 N   p    f   [sp] any   -   Lenka, Lenky
 N   p    n   [sp] any   -   Bradlo, Ústí
 N   c    m    s    n    -   kluk,      vor
 N   c    m    s    g    -   kluka,     voru
 N   c    m    s    d    -   klukovi,   voru
 N   c    m    s    a    y   kluka
 N   c    m    s    a    n              vor
 N   c    m    s    v    -   kluku,     vore!
 N   c    m    s    l    -   klukovi,   voru
 N   c    m    s    i    -   klukem,    vorem
 N   c    m    p    n    y   kluci
 N   c    m    p    n    n              vory
 N   c    m    p    g    -   kluků      vorů
 N   c    m    p    d    -   klukům     vorům

N   c    m    p    a    -   kluky      vory
N   c    m    p    v    -   kluci!     vory!
N   c    m    p    l    -   klucích    vorech
N   c    m    p    i    -   kluky      vory
N   c   [fn]  s    n    -   sova,      kolo
N   c   [fn]  s    g    -   sovy,      kola
N   c   [fn]  s    d    -   sově       kolu
N   c   [fn]  s    a    -   sovu       kolo
N   c   [fn]  s    v    -   sovo!      kolo!
N   c   [fn]  s    l    -   sově       kole / kolu
N   c   [fn]  s    i    -   sovou      kolem
N   c   [fn]  p    n    -   sovy       kola
N   c   [fn]  p    g    -   sov        kol
N   c   [fn]  p    d    -   sovám      kolům
N   c   [fn]  p    a    -   sovy       kola
N   c   [fn]  p    v    -   sovy!      kola!
N   c   [fn]  p    l    -   sovách     kolech
N   c   [fn]  p    i    -   sovami     koly
N   c   [fn]  d    i    -   rukama     očima    (dual number)
*** **** **** **** **** **** =============================================
Note: In the Combinations above, 'any' is a variable standing for
     any admissible value.

7.2 Verb (V)

2.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C  Example            Czech term
= ============== ============== = =================== ========================
1 Type           main           m dělat               významové sloveso
                auxiliary      a mít                 pomocné sloveso
                modal          o muset               modální sloveso
                copula         c být                 spona (kopula)
- -------------- -------------- - ------------------- ------------------------
2 VForm          indicative     i dělám               indikativ
                imperative     m dělej!              imperativ
                conditional    c by                  podmiňovací částice
                infinitive     n dělat               infinitiv
                participle     p dělal, dělán        příčestí
           l.s. transgressive  t dělaje, udělavši    přechodník
- -------------- -------------- - ------------------- ------------------------
3 Tense          present        p dělám               prézens
                future         f budu, poběžím       futurum
                past           s dělal               čas minulý
- -------------- -------------- - ------------------- ------------------------
4 Person         first          1 dělám               první osoba
                second         2 děláš               druhá osoba

                 third          3 dělá                třetí osoba
- -------------- -------------- - ------------------- ------------------------
5 Number         singular       s dělám               singulár
                 plural         p děláme              plurál
- -------------- -------------- - ------------------- ------------------------
6 Gender         masculine      m dělal               masculinum /rod mužský
                 feminine       f dělala              femininum /rod ženský
                 neuter         n dělalo              neutrum /rod střední
* ************** ************** * ------------------- ------------------------
7 Voice          active         a dělal               aktivum
                 passive        p dělán               pasívum
- -------------- -------------- - ------------------- ------------------------
8 Negative       no             n dělám               kladné sloveso
                 yes            y nedělám             záporné sloveso
- -------------- -------------- - ------------------- ------------------------
9 Definiteness                  -
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- - ------------------- ------------------------
12Animate   l.s. no             n dělaly              neživotné
            l.s. yes            y dělali              životné
- -------------- -------------- - ------------------- ------------------------
13Clitic_s  l.s. no             n udělal              tvar bez klit. -s
            l.s. yes            y udělals             tvar s klit. -s
==============================================================================
Notes
1.  The verb "být" (E. "to be") in all its functions is characte-
    rized as Type=c  (i.e. the copula), which clearly  is an over-
    simplification because  the verb has  more meanings (auxiliary
    etc.).
2.  Auxiliary verbs (Type=a) include  neither the verb "být" (see
    above), nor the modal verbs.
3.  The 'past participle' in Czech is used for expressing compound
    active past Tense and is encoded as:
    Type=p(articiple), Tense=p(ast), Voice=a(ctive).
4.  Passive participles are encoded as:
    Type=participle, Tense='-', Voice=passive.
    This encoding is  used only for the passive  participle in the
    predicative  position,   e.g.  "byl  chycen"   (E.  "(he)  was
    caught").  The case  marked passive  participles, both  in the
    predicative and in the  attributive position, e.g. "padělaný
    rukopis" (E. "forged manuscript"),  are classified as (qua-
    lificative) adjectives.
5.  Adjectival  active and  passive participles,  e.g. "stojící"
    (E. "standing") or "udělaný"  (E. "performed" or "done", cf.
    Note 4 above) are classified as adjectives.
6.  Negative verbs are marked  as Negative=y, whereas non-negative
    verbs are marked as Negative=n.
7.  The term transgressive roughly corresponds to the term 'verbal

participle'. The transgressives distinguish:
 Number (singular, plural),
 Gender (masculine, non-masculine)
 Tense (past, present)
The Gender opposition of  masculine animate vs. masculine ina-
nimate is  neutralized in transgressives;  thus, there is  one
masculine form and one non-masculine (the form of feminine and
neuter is neutralized).
8.  Gender and Animate values  correspond to those associated with
Nouns and are necessary to account properly for agreement.
9.  Gender  manifests itself  in past  participles, passive parti-
ciples and transgressives only.
10. Normally,  Verbs  form  the  future  Tense periphrastically by
auxiliary  "být" (E.  "to be")  plus infinitive  of the  main
Verb.  In addition  to the  copula, there  are, however,  some
Verbs  which  form  future  Tense  non-periphrastically,  i.e.
synthetically (Verbs of motion).  Such verbal forms are marked
as Tense=f.
11. The  'yes' value  of the  Clitic_s attribute  denotes a verbal
form having the clitic morpheme 's' appended as a suffix. This
's' morpheme  expresses 2nd Person  singular present Tense  of
the auxiliary Verb  "být" (i.e. the form "jsi").  There is no
intermediate  hyphen  between  the  verbal  form  and  the 's'
morpheme.
12. The  Clitic_s  attribute  is  specified  for  VForm=infinitive
(VForm=n) and Vform=p(articiple) only.
13. Some  modal and  auxiliary Verbs  do not  form imperative  and
transgressive.
14. The   voice  value   ('a'  or   'p')  is   not  specified  for
VForm=c(onditional)  and VForm=t(ransgressive),  in which case
VForm='-'.

2.2 Combinations
*** **** **** **** **** **** ***** ---- ---- ---- ---- ========================
PoS Type VFrm Tens Pers Numb Gend  Voic Neg  Anim Cl_s Examples
*** **** **** **** **** **** ***** ---- ---- ---- ---- ========================
V   m     n    -    -    -    -   [na] [ny]   -  [ny] prát, neprat
V   a     n    -    -    -    -   [na] [ny]   -  [ny] mít, nemít
V   o     n    -    -    -    -   [na] [ny]   -  [ny] muset, nemuset
V   c     n    -    -    -    -   [na] [ny]   -  [ny] být, nebýt
V   c     c    -    3    s    -     -    -    -    -  by
V   c     c    -  [12]   s    -     -    -    -    -  bych, bys
V   c     c    -  [12]   p    -     -    -    -    -  bychom, byste
V   c     i    f  [123]  s    -     a    n    -    -  budu, budeš, bude
V   c     i    f  [123]  s    -     a    y    -    -  nebudu, nebudeš, nebude
V   c     i    f  [123]  p    -     a    n    -    -  budeme, budete, budou
V   c     i    f  [123]  p    -     a    y    -    -  nebudeme, nebudete, nebudou
V   m     i    f  [123]  s    -     a    n    -    -  poletím, poletíš, poletí
V   m     i    f  [123]  s    -     a    y    -    -  nepoletím, nepoletíš,
                                                   nepoletí
V   m     i    f  [123]  p    -     a    n    -    -  poletíme, poletíte, poletí

V   m     i    f  [123]  p    -     a    y    -    -  nepoletíme, nepoletíte,
                                                      nepoletí
V   c     i    p  [123]  s    -     a    n    -    -  jsem, jsi, je
V   c     i    p  [123]  s    -     a    y    -    -  nejsem, nejsi, není
V   c     i    p  [123]  p    -     a    n    -    -  jsme, jste, jsou
V   c     i    p  [123]  p    -     a    y    -    -  nejsme, nejste, nejsou
V   m     i    p  [123]  s    -     a    n    -    -  třu, třeš, tře
V   m     i    p  [123]  s    -     a    y    -    -  netřu, netřeš, netře
V   m     i    p  [123]  p    -     a    n    -    -  třeme, třete, třou
V   m     i    p  [123]  p    -     a    y    -    -  netřeme, netřete, netřou
V   a     i    p  [123]  s    -     a    n    -    -  mám, máš, má
V   a     i    p  [123]  s    -     a    y    -    -  nemám, nemáš, nemá
V   a     i    p  [123]  p    -     a    n    -    -  máme, máte, mají
V   a     i    p  [123]  p    -     a    y    -    -  nemáme, nemáte, nemají
V   o     i    p  [123]  s    -     a    n    -    -  musím, musíš, musí
V   o     i    p  [123]  s    -     a    y    -    -  nemusím, nemusíš,
                                                      nemusí
V   o     i    p  [123]  p    -     a    n    -    -  musíme, musíte,
                                                      musejí/musí
V   o     i    p  [123]  p    -     a    y    -    -  nemusíme, nemusíte,
                                                      nemusejí
V   c     m    p    1    p    -     a   [ny]  -    -  buďme!, nebuďme!
V   c     m    p    2   [sp]  -     a    n    -    -  buď!, buďte!
V   c     m    p    2   [sp]  -     a    y    -    -  nebuď!, nebuďte!
V   m     m    p    1    p    -     a   [ny]  -    -  pracujme!, nepracujme!
V   m     m    p    2   [sp]  -     a    n    -    -  pracuj!, pracujte!
V   m     m    p    2   [sp]  -     a    y    -    -  nepracuj!, nepracujte!
V   c     p    s    -    s  [mfn]   a    n    -    n  byl, byla, bylo
V   c     p    s    -    s  [mfn]   a    y    -    n  nebyl, nebyla, nebylo
V   c     p    s    -    s  [mfn]   a    n    -    y  byls, bylas, bylos
V   c     p    s    -    s  [mfn]   a    y    -    y  nebyls, nebylas,
                                                      nebylos
V   c     p    s    -    p   m      a    n  [yn]   n  byli, byly
V   c     p    s    -    p   m      a    y  [yn]   n  nebyli, nebyly
V   c     p    s    -    p  [fn]    a    n    -    n  byly, byla
V   c     p    s    -    p  [fn]    a    y    -    n  nebyly, nebyla
V   a     p    s    -    s  [mfn]   a    n    -    n  měl, měla, mělo
V   a     p    s    -    s  [mfn]   a    y    -    n  neměl, neměla,
                                                      nemělo
V   a     p    s    -    s  [mfn]   a    n    -    y  měls, mělas, mělos
V   a     p    s    -    s  [mfn]   a    y    -    y  neměls, nemělas,
                                                      nemělos
V   a     p    s    -    p   m      a    n  [yn]   n  měli, měly
V   a     p    s    -    p   m      a    y  [yn]   n  neměli, neměly
V   a     p    s    -    p  [fn]    a    n    -    n  měly, měla
V   a     p    s    -    p  [fn]    a    y    -    n  neměly, neměla
V   o     p    s    -    s  [mfn]   a    n    -    n  musel, musela, muselo
V   o     p    s    -    s  [mfn]   a    y    -    n  nemusel, nemusela,
                                                      nemuselo
V   o     p    s    -    s  [mfn]   a    n    -    y  musels, muselas,
                                                      muselos
V   o     p    s    -    s  [mfn]   a    y    -    y  nemusels, nemuselas,
                                                      nemuselos

V   o     p    s    -    p   m      a    n  [yn]   n  museli, musely
V   o     p    s    -    p   m      a    y  [yn]   n  nemuseli, nemusely
V   o     p    s    -    p  [fn]    a    n    -    n  musely, musela
V   o     p    s    -    p  [fn]    a    y    -    n  nemusely, nemusela
V   m     p    s    -    s  [mfn]   a    n    -    n  dělal, dělala,
                                                      dělalo
V   m     p    s    -    s  [mfn]   a    y    -    n  nedělal, nedělala,
                                                      nedělalo
V   m     p    s    -    s  [mfn]   a    n    -    y  dělals, dělalas,
                                                      dělalos
V   m     p    s    -    s  [mfn]   a    y    -    y  nedělals, nedělalas,
                                                      nedělalos
V   m     p    s    -    p   m      a    n  [yn]   n  dělali, dělaly
V   m     p    s    -    p   m      a    y  [yn]   n  nedělali, nedělaly
V   m     p    s    -    p  [fn]    a    n    -    n  dělaly, dělala
V   m     p    s    -    p  [fn]    a    y    -    n  nedělaly, nedělala
V   m     p    -    -    s  [mfn]   p    n    -    n  udělán, udělána
                                                      uděláno
V   m     p    -    -    s  [mfn]   p    y    -    n  neudělán, neudělána
                                                      neuděláno
V   m     p    -    -    p   m      p    n  [yn]   n  uděláni, udělány
V   m     p    -    -    p   m      p    y  [yn]   n  neuděláni, neudělány
V   m     p    -    -    p  [fn]    p    n    -    n  udělány, udělána
V   m     p    -    -    p  [fn]    p    y    -    n  neudělány, neudělána
V   c     t    p    -    s   m      -  [ny]   -    -  jsa, nejsa
V   c     t    p    -    s  [fn]    -  [ny]   -    -  jsouc, nejsouc
V   c     t    p    -    p   -      -  [ny]   -    -  jsouce, nejsouce
V   m     t    p    -    s   m      -  [ny]   -    -  dělaje, nedělaje
V   m     t    p    -    s  [fn]    -  [ny]   -    -  dělajíc, nedělajíc
V   m     t    p    -    p   -      -  [ny]   -    -  dělajíce nedělajíce
V   c     t    s    -    s   m      -  [ny]   -    -  byv, nebyv
V   c     t    s    -    s  [fn]    -  [ny]   -    -  byvši, nebyvši
V   c     t    s    -    p   -      -  [ny]   -    -  byvše, nebyvše
V   m     t    s    -    s   m      -  [ny]   -    -  udělav, neudělav
V   m     t    s    -    s  [fn]    -  [ny]   -    -  udělavši, neudělavší
V   m     t    s    -    p   -      -  [ny]   -    -  udělavše, neudělavše
*** **** **** **** **** **** ***** ---- ---- ---- ---- =======================

7.3 Adjective (A)

3.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           qualificative  f dobrý               kvalitativní adjektivum
                possessive     s matčin              posesívní adjektivum
- -------------- -------------- --------------------- ------------------------

2 Degree         positive       p dobrý               pozitiv / 1. stupeň
                 comparative    c lepší               komparativ / 2. stupeň
                 superlative    s nejlepší            superlativ / 3. stupeň
- -------------- -------------- -  ------------------ ------------------------
3 Gender         masculine      m dobrý               masculinum / rod mužský
                 feminine       f dobrá               femininum / rod ženský
                 neuter         n dobré               neutrum  / rod střední
- -------------- -------------- --------------------- ------------------------
4 Number         singular       s dobrý               singulár
                 plural         p dobří               plurál
                 dual           d dobrýma             duál
- -------------- -------------- --------------------- ------------------------
5 Case           nominative     n dobrý               nominativ
                 genitive       g dobrého             genitiv
                 dative         d dobrému             dativ
                 accusative     a dobrého             akuzativ
                 vocative       v dobrý!              vokativ
                 locative       l dobrém              lokál
                 instrumental   i dobrým              instrumentál
* ***************************** * ------------------- ------------------------
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- - ------------------- ------------------------
8 Animate        no             n dobré               neživotné
                 yes            y dobří               životné
- -------------- -------------- - ------------------- ------------------------
9 Formation l.s. nominal        n stár                jmenný tvar
            l.s. compound       c starý               složený tvar
- -------------- -------------- --------------------- ------------------------
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================
Notes
1. Three  deverbative  adjectival  participles,  i.e. past active
   participle, passive  participle and present  active participle
   are not  distinguished. They are conflated  in the 'qualifica-
   tive' value of the Type attribute (Type=f).
2. Only qualificative (and passive  participle) Adjectives can be
   specified for Degree and for  nominal (short) Form.
3. The attributes Gender,  Number, Case and  Animate correspond to
   the same categories  within the Nouns. They  are necessary for
   the proper account of agreement of Adjectives with Nouns.
4. The Formation  attribute distinguishes a  nominal (short) form
   from a so-called compound (long) form of an Adjective. The no-
   minal form can be used in the predicative function only. It is
   specified for nominative and accusative Case only.
5. The Adjectives  which have no  degrees of comparison  have the

  Degree value equal to p(ositive).
3.2 Combinations
*** **** ***** **** **** **** **** **** ======================================
PoS Type Degr  Gend Numb Case Anim Form Examples
*** **** ***** **** **** **** **** **** ======================================
A   s     -    m     s    n    -   -   otcův
A   s     -    m     s    g    -   -   otcova
A   s     -    m     s    d    -   -   otcovu
A   s     -    m     s    a    y   -   otcova
A   s     -    m     s    a    n   -            otcův
A   s     -    m     s    v    -   -   otcův!
A   s     -    m     s    l    -   -   otcově / otcovu
A   s     -    m     s    i    -   -   otcovým
A   s     -    m     p    n    y   -   otcovi
A   s     -    m     p    n    n   -            otcovy
A   s     -    m     p    g    -   -   otcových
A   s     -    m     p    d    -   -   otcovým
A   s     -    m     p    a    -   -   otcovy
A   s     -    m     p    v    y   -   otcovi!
A   s     -    m     p    v    n   -             otcovy!
A   s     -    m     p    l    -   -   otcových
A   s     -    m     p    i    -   -   otcovými
A   s     -   [fn]   s    n    -   -   otcova,  otcovo
A   s     -   [fn]   s    g    -   -   otcovy,  otcova
A   s     -   [fn]   s    d    -   -   otcově, otcovu
A   s     -   [fn]   s    a    -   -   otcovu,  otcovo
A   s     -   [fn]   s    v    -   -   otcova!, otcovo
A   s     -   [fn]   s    l    -   -   otcově,  otcově / otcovu
A   s     -   [fn]   s    i    -   -   otcovou, otcovým
A   s     -   [fn]   p    n    -   -   otcovy,  otcova
A   s     -   [fn]   p    g    -   -   otcových
A   s     -   [fn]   p    d    -   -   otcovým
A   s     -   [fn]   p    a    -   -   otcovy,  otcova
A   s     -   [fn]   p    v    -   -   otcovy!  otcova!
A   s     -   [fn]   p    l    -   -   otcových
A   s     -   [fn]   p    i    -   -   otcovými
A   s     -   [fn]   d    i    -   -   otcovýma
A   f   [pcs]  m     s    n    -   c   dobrý
A   f   [pcs]  m     s    g    -   c   dobrého
A   f   [pcs]  m     s    d    -   c   dobrému
A   f   [pcs]  m     s    a    y   c   dobrého
A   f   [pcs]  m     s    a    n   c   dobrý
A   f   [pcs]  m     s    v    -   c   dobrý!
A   f   [pcs]  m     s    l    -   c   dobrém
A   f   [pcs]  m     s    i    -   c   dobrým
A   f   [pcs]  m     p    n    y   c   dobří
A   f   [pcs]  m     p    n    n   c            dobré
A   f   [pcs]  m     p    g    -   c   dobrých
A   f   [pcs]  m     p    d    -   c   dobrým
A   f   [pcs]  m     p    a    -   c   dobré
A   f   [pcs]  m     p    v    y   c   dobří

A   f   [pcs]  m     p    v    n   c            dobré
A   f   [pcs]  m     p    l    -   c   dobrých
A   f   [pcs]  m     p    i    -   c   dobrými
A   f   [pcs] [fn]   s    n    -   c   dobrá,  dobré
A   f   [pcs] [fn]   s    g    -   c   dobré,  dobrého
A   f   [pcs] [fn]   s    d    -   c   dobré,  dobrému
A   f   [pcs] [fn]   s    a    -   c   dobrou,  dobré
A   f   [pcs] [fn]   s    v    -   c   dobrá!, dobré
A   f   [pcs] [fn]   s    l    -   c   dobré,  dobrém
A   f   [pcs] [fn]   s    i    -   c   dobrou,  dobrým
A   f   [pcs] [fn]   p    n    -   c   dobré,  dobrá
A   f   [pcs] [fn]   p    g    -   c   dobrých
A   f   [pcs] [fn]   p    d    -   c   dobrým
A   f   [pcs] [fn]   p    a    -   c   dobré,  dobrá
A   f   [pcs] [fn]   p    v    -   c   dobré,  dobrá
A   f   [pcs] [fn]   p    l    -   c   dobrých
A   f   [pcs] [fn]   p    i    -   c   dobrými
A   f   [pcs] [fn]   d    i    -   c   dobrýma
A   f     -   [mfn]  s    n    -   n   stár, stára, stáro
A   f     -   [mfn]  s    a    -   n   stára, stáru, stáro
A   f     -     m    p    n   [ny] n   stáři, stáry
A   f     -   [fn]   p    n    -   n   stáry, stára
A   f     -   [mfn]  p    a    -   n   stáry, stáry, stára
*** **** ***** **** **** **** **** **** ======================================

7.4 Pronoun (P)

4.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           personal       p já                  osobní
                demonstrative  d ten                 ukazovací
                indefinite     i někdo               neurčité
                possessive     s její                posesívní
                interrogative  q kdo                 tázací
                relative       r jenž                vztažné / relativní
                reflexive      x se                  zvratné  / reflexívní
                negative       z nikdo               záporná / negativní
                general        g každý               totální
- -------------- -------------- --------------------- ------------------------
2 Person         first          1 já                  první osoba
                second         2 ty                  druhá osoba
                third          3 on                  třetí osoba
- -------------- -------------- --------------------- ------------------------
3 Gender         masculine      m on                  masculinum
                feminine       f ona                 femininum

                 neuter         n to                  neutrum
- -------------- -------------- --------------------- ------------------------
4 Number         singular       s ty                  singulár
                 plural         p oni                 plurál
                 dual           d těma                duál
- -------------- -------------- --------------------- ------------------------
5 Case           nominative     n ty                  nominativ
                 genitive       g tebe                genitiv
                 dative         d tobě,ti             dativ
                 accusative     a tebe, tě            akuzativ
                 locative       l tobě                lokál
                 instrumental   i tebou               instrumentál
- -------------- -------------- - ------------------- ------------------------
6 Owner_Number   singular       s můj                 singulár
                 plural         p jejich              plurál
- -------------- -------------- - ------------------- ------------------------
7 Owner_Gender   masculine      m jehož               masculinum
                 feminine       f jejíhož             femininum
                 neuter         n jehož               neutrum
********************************* ------------------- ------------------------
8 Clitic         no             n tobě                nepříklonné
                 yes            y ti                  příklonka
- -------------- -------------- - ------------------- ------------------------
9 Referent_Type  personal       p se, sebe, si        osobní zvratné
                 possessive     s svůj                zvratné posesívní
- -------------- -------------- - ------------------- ------------------------
10Syntactic_Type nominal        n on                  syntaktické substantivum
                 adjectival     a který               syntaktické adjektivum
- -------------- -------------- - ------------------- ------------------------
11Definiteness                  -
- -------------- -------------- - ------------------- ------------------------
12Animate        no             n který               neživotné
                 yes            y kteří, kterého      životné
- -------------- -------------- - ------------------- ------------------------
13Clitic_s  l.s. yes            y tys,ses             zájmeno s klitickým -s
            l.s. no             n ty                  zájmeno bez klitic. -s
- -------------- -------------- - ------------------- ------------------------
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person                  -
- -------------- -------------- -
16Owned_Number                  -
================ ================
Notes
1. Gender,  Number,  Case  and  Animate  correspond  to  the same
   categories as specified for Nouns.  They are  necessary for
   the proper account of agreement of adjectival Pronouns with
   Nouns.
2. Type=reflexive ecompasses all reflexive pronouns (``se'', ``sebe'',
   ``si'', ``svůj'') as well as "se" in  its role as the obligatory
   particle of  reflexive verbs.  Personal and  possessive reflexives

  are further distinguished via the Referent_Type attribute.
  "se" in all its roles will be marked as the reflexive personal
  clitic pronoun.
3. Pronouns  are distinguished  between having  a (syntactically)
  nominal and  (syntactically) adjectival function.  All
  pronominal types except the demonstrative and  possessive one
  can be nominal, and all except for the personal one can be
  adjectival.
5. Referent_Type  is  used  to  distinguish  personal  reflexives
  (which include  "se" in all its  functions) from the possesive
  reflexives ("svůj").
6. Negative and general Pronouns ("general" Pronouns concern the
  Pronouns like  "všichni" [E. "all"],  "každý" [E. "every"]
  etc.)  are important  from  the  viewpoint of  their syntactic
  distribution.
7. The  Clitic attribute  distinguishes clitical  vs. nonclitical
  pronominal forms, e.g. "ti" vs. "tobě".
8. The 'yes' value of the Clitic_s attribute denotes a pronominal
  form having the clitic morpheme 's' appended as a suffix. This
  morpheme expresses  2nd Person singular  present Tense of  the
  auxiliary Verb  "být" (E. "to be").  There is no intermediate
  hyphen between the Pronoun and the 's' morpheme.
9. Owner_Number and  Owner_Gender concern the  possessor's Number
  and Gender, respectively.

4.2 Combinations
*** **** ***** ***** **** **** **** **** ---- --- --- ----- ---- =============
PoS Type Pers  Gend  Numb Case Ow_N Ow_G Clit Ref Syn Anim  Cl_s  Examples
*** **** ***** ***** **** **** **** **** ---- --- --- ----- ---- =============
P    x    -     -     -  [ga]   -    -    y   p   n    -    n   se
P    x    -     -     -  [ga]   -    -    y   p   n    -    y   ses
P    x    -     -     -  [ga]   -    -    n   p   n             sebe
P    x    -     -     -   d     -    -    y   p   n    -    n   si
P    x    -     -     -   d     -    -    y   p   n    -    y   sis
P    x    -     -     -  [dl]   -    -    n   p   n    -   [ny] sobě
P    x    -     -     -   i     -    -    n   p   n    -   [ny] sebou
P    x    -   [mfn]  all any    -    -    n   s   a  [ny-]  n   svůj
P    p    1     -   [sp] all    -    -  [ny]  -   n    -   [ny] já, my
P    p    2     -   [sp] all    -    -  [ny]  -   n    -    n   ty, vy
P    p    2     -     s   n     -    -    n   -   n    -    y   tys
P    p    3   [mfn] [sp] all    -    -  [ny]  -   n    -   [ny] on, oni,
                                                                ony, ona
P    s    1   [mfn]  all all    s    -    n   -   a  [ny-] [ny] můj
P    s    2   [mfn]  all [na]   s    -    n   -   a  [ny-] [ny] tvůj
P    s    3     -     -   -     s  [mn]   n   -   a    -   [ny] jeho
P    s    3   [mfn]  all all    s    f    n   -   a    -   [ny] její
P    s    3     -     -   -     p  [mfn]  n   -   a    -    -   jejich
P    d    -   [mfn]  all all    -    -    n   -   a  [ny-] [ny] 1)
P    i    -   [mn]    s  all    -    -    n   -   n    -    -   2)
P    i    -   [mfn] [sp] all    -    -    n   -   a  [ny-]  -   3)
P   [qr]  -   [mn]    s  all    -    -    n   -   n    -    -   4)

 P   [qr]  -   [mfn] [sp] all    -    -    n   -   a  [ny-]  -   5)
 P    z    -   [mn]    s  all    -    -    n   -   n    -    -   6)
 P    z    -   [mfn] [sp] all    -    -    n   -   a  [ny-]  -   7)
 P    g    -   [mfn] [sp] all    -    -    n   -   n  [ny-]  -   8)
*** **** ***** ***** **** **** **** **** ---- --- --- ----- ---- =============
1) ten, tento, tamten, tenhle, tenhleten, takový, onaký
2) někdo, nějaký, kdovíkdo, leckdo, leccos, něco
3) jaký, který, kdovíjaký, lecjaký
4) kdo, co
5) jaký, který, jenž, což
6) nikdo, nic
7) žádný, nijaký
8) všechen, každý

7.5 Determiner (D)

Not applicable.

7.6 Article (T)

Not applicable.

7.7 Adverb (R)

7.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           general        g dobře               příslovce
- -------------- -------------- --------------------- ------------------------
2 Degree         positive       p málo                pozitiv / 1. stupeň
                 comparative    c méně                komparativ / 2. stupeň
                 superlative    s nejméně             superlativ / 3. stupeň
* ************** ************** * ------------------- ------------------------
3 Clitic                        -
- -------------- -------------- -
4 Number                        -

- -------------- -------------- -
5 Person                        -
= ============== ============== =
Notes
1. The  classification  given  above  is considerably simplified  as
   compared  to traditional  Czech grammars.  There can  be other
   classifications as well, eg.:
   (a) syntactic: Adverb modifying a Verb
                  Adverb modifying an Adjective
                  Adverb modifying another Adverb
                  Adverb modifying a Noun from the left
                  Adverb modifying a Noun from the right
   (b) semantic   temporal
                  spatial
                  qualitative
                   ...
   We refrained from this typology here.
2. Particles form a separate part of speech category (see below)
   as is customary in Czech grammars.
3. The Adjectives  which have no  degrees of comparison  have the
   Degree value equal to p(ositive) similarly as Adjectives.

7.2 Combinations
*** **** **** ================================================================
Pos Type Deg  Examples
*** **** **** ================================================================
 R   g    p   dobře
 R   g    c   lépe
 R   g    s   nejlépe
 R   g    p   tady
*** **** **** ================================================================

7.8 Adposition (S)

8.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           preposition    p nad                 předložka
- -------------- -------------- - ------------------- ------------------------
2 Formation      simple         s na, v               predložka (pravá)
                 compound       c naň, oč             kontrakce předl. + zájm.
* ************** ************** * ------------------- ------------------------

3 Case
  (req.by prep.) genitive       g bez                 genitiv
                 dative         d proti               dativ
                 accusative     a pro                 akuzativ
                 locative       l v                   lokál
                 instrumental   i s                   instrumentál
- -------------- -------------- - ---- -------------- ------------------------
4 Clitic                        -
=================================
Notes
1. Czech has only Prepositions, no postpositions.
2. For the  disambiguation of word forms  belonging to declension
   parts of speech it seems  necessary to include the information
   about the case which each Preposition requires.
3. A preposition  can be  contracted with  a pronoun;  such a
   preposition has Formation=c(ompound).

8.2 Combinations
*** **** **** ---- ===========================================================
PoS Type Form Case Examples
*** **** **** ---- ===========================================================
 S    p    s    g  bez, kromě
 S    p    s    d  k, proti
 S    p    s    a  pro, mimo
 S    p    s    l  o, v, při
 S    p    s    i  s, pod
 S    p    c    -  naň, zač
*** **** **** ---- ===========================================================

7.9 Conjunction (C)

9.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           coordinating   c a, ani              spojka souřadicí
                 subordinating  s aby, protože        spojka podřadicí
* ************** ************** * ------------------- ------------------------
2 Formation                     -
- -------------- -------------- -
3 Coord_Type                    -
- -------------- -------------- -
4 Sub_Type                      -

- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- - ------------------- ------------------------
6 Number         singular       s kdybych             spojka 'kdyby'+sg. od 'by'
                 plural         p kdybyste            spojka 'kdyby'+pl. od 'by'
- -------------- -------------- - ------------------- ------------------------
7 Person         first          1 abych               spojka 'aby'+1.os. od 'by'
                 second         2 abyste              spojka 'aby'+2.os. od 'by'
                 third          3 aby                 spojka 'aby'+3.os. od 'by'
================================================================================
Notes
1. Conjunctions "abych", "abys", ..., "kdybych", ... are composed
   of  the Conjunction  "aby" or  "kdyby" and the conditional
   particle ``by'' ("bych", "bys",  "by", "bychom", "byste").
   These Conjunctions are therefore specified for Person  and
   Number for the sake of agreement relation.
2. The class of two-part Conjunctions has not been introduced.

9.2 Combinations
*** **** ---- ---- ===========================================================
PoS Type Numb Pers Examples
*** **** ---- ---- ===========================================================
 C    c    -   -   a, i, ba, neboť, ani
 C    s    -   -   že, protože, zda
 C    s    -   3   aby, kdyby
 C    s    s  [12] abych, abys, kdybych, kdybys
 C    s    p  [12] abychom, abyste, kdybychom, kdybyste
*** **** =====================================================================

7.10 Numeral (M)

10.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           cardinal       c dva                 číslovka základní
                 ordinal        o pátý                číslovka řadová
                 multiple       m dvakrát             číslovka násobná
                 special        s dvojí               číslovka druhová
- -------------- -------------- - ------------------- ------------------------
2 Gender         masculine      m dva                 masculinum
                 feminine       f dvě                 femininum
                 neuter         n dvě                 neutrum

- -------------- -------------- - ------------------- ------------------------
3 Number         singular       s druhý               singulár
                 plural         p druzí               plurál
                 dual           d druhýma             duál
- -------------- -------------- --------------------- ------------------------
4 Case           nominative     n dva, dvě            nominativ
                 genitive       g dvou                genitiv
                 dative         d dvěma               dativ
                 accusative     a dva, dvě            akuzativ
                 locative       l dvou                lokál
                 instrumental   i dvěma               instrumentál
- -------------- -------------- - ------------------- ------------------------
5 Form           digit          d 5                   číslo (arabské)
                 roman          r MCMVIII             číslo (římské)
                 letter         l šedesát             číslovka slovem
* ************** ************** * ------------------- ------------------------
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- - ------------------- ------------------------
8 Class          definite1      1 jeden
                 definite2      2 dva
                 definite34     3 tři
                 definite       f pět
                 demonstrative  d tolik
                 indefinite     i několik
                 interrogative  q kolik
                 relative       r kolik
---------------- -------------- - ------------------- ------------------------
9 Animate        no             n páté                neživotné
                 yes            y pátí                životné
- -------------- -------------- - ------------------- ------------------------
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================
Notes
1. Numerals have been specified as a separate category because of
   their specific  syntactic distribution. We  have specified two
   syntactic classifications by means  of the attributes Type and
   Class;  they  concern  different  syntactic distributions. For
   instance "několik" (E. "several") will be characterized as:
   Type:  cardinal
   Class: indefinite
2. Among   the  definite   numbers  there   are  four  subclasses
   (definite1, definite2,  definite34, definite) which  differ in
   their syntactic distribution and contain the following Numerals:
   {1}, {2}, {3,4}, {5,6,...}
3. Gender,  Number and  Case  correspond  to the  same categories

   as specified for Nouns. They  are necessary  for the  proper
   account of agreement of Numerals with Nouns.
4. In direct cases (nominative, accusative) the cardinal numerals
   belonging to the class with one  of the values {fdiqr} (see above)
   behave like neuter Nouns in singular, whereas in oblique cases
   they behave like adjectival attributes in plural.

10.2 Combinations
*** **** ****  **** **** ---- ------ ----- ===================================
PoS Type Gend  Numb Case Form Class  Anim  Examples
*** **** ****  **** **** ---- ------ ----- ===================================
 M    -    -     -    -    d  [123f]   -   56
 M    -    -     -    -    r  [123f]   -   MVIII
 M    c  [mfn]   s   any   l     1     -   jeden, jedna, jedno
 M    c  [mfn]   p   any   l     2     -   dva, dvě
 M    c    -     p   any   l     3     -   tři, čtyři
 M    c    n    [sp] any   l     f     -   pět, deset, padesát
 M    c    n    [sp] any   l  [diqr]   -   tolik, několik, kolik
 M    o  [mfn] [spd] any   l    any  [ny-] první, druhý, několikátý
 M    m    -     -    -    l    any    -   dvakrát, pětkrát, několikrát
 M    s  [mfn] [spd]  n    l    any  [ny-] dvojí, troje, paterý
*** **** ****  **** **** ---- ------ ----- ===================================
Note: In the Combinations above, 'any' is a variable standing for
      any admissible value.

7.11 Interjection (I)

= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type                          - au!, och!           citoslovce
- -------------- -------------- - ------------------- ------------------------
2 Formation                     -
=================================

7.12 Residual (X)

= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
                                - sic, a+b            Zbývající elementy

= ============== ============== =

7.13 Abbreviation (Y)

= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Syntactic_Type                - atd.                zkratka
- -------------- -------------- - ------------------- ------------------------
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
- -------------- -------------- -
5 Definiteness                  -
=================================

7.14 Particle (Q)

= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type                          - ano, ne             partikule / částice
- -------------- -------------- - ------------------- ------------------------
2 Formation                     -
- -------------- -------------- -
3 Clitic                        -
=================================

Application to Bulgarian

Authors: Radoslav Pavlov and Ludmila Dimitrova (IM-BAS), Lydia Sinapova (IIT-BAS), and Kiril Iv. Simov (LML-BAS).

8.1 Noun (N)

= ============== ============== ============= =
P ATT            VAL              Example     C
= ============== ============== ============= =
1 Type           common           kniga       c
                 proper           Ivan        p
- -------------- -------------- ------------- -
2 Gender         masculine        stol        m
                 feminine         masa        f
                 neuter           vreteno     n
- -------------- -------------- ------------- -
3 Number         singular         momtche     s
                 plural           stolove     p
          l.s.   count            (dva) stola t
- -------------- -------------- ------------- -
4 Case           nominative       narod       n
                 vocative         narode      v
* ***************************** ************* *
5 Definiteness   no               utchitel    n
                 yes              zhenata     y
           l.s.  short_art        moliva      s
           l.s.  full_art         molivyt     f
- -------------- -------------- ------------- -
6 Clitic                            -         -
- -------------- -------------- ------------- -
7 Animate                           -         -
- -------------- -------------- ------------- -
8 Owner_Number                      -         -
- -------------- -------------- ------------- -
9 Owner_Person                      -         -

Multext-East / Concede  Morphosyntactic specifications V.2— Bulgarian  126

---------------- -------------- ------------- -
10Owned_Number                      -         -
=============================== ============= =

Combinations
=== ====== ====== ===== ==== === ==== ==== ==== ==== ==== =======
POS  Type   Gend   Numb Case Def Clit Anim OwnN OwnP OwdN Example
=== ====== ====== ===== ==== === ==== ==== ==== ==== ==== =======
N    [cp]  [mfn]   [sp]  -    n    -   -    -    -    -       1.
N    [cp]    m      s    -   [sf]  -   -    -    -    -       2.
N    [cp]  [mf]     -    v    -    -   -    -    -    -       3.
N    [cp]  [mfn]    p    -    y    -   -    -    -    -       4.
N    [cp]  [fn]     s    -    y    -   -    -    -    -       5.
N     c      m      t    -    -    -   -    -    -    -       6.
=== ====== ======= ==== ==== === ==== ==== ==== ==== ==== =======
Examples:
1. narod, zhena, selo, Ivan, Penka
2. naroda, narodyt
3. narode, zheno
4. narodite, zhenite, selata
5. zhenata, seloto
6. naroda

8.2 Verb (V)

= ============== ============== ============== =
P ATT            VAL             Example       C
= ============== ============== ============== =
1 Type           main            govorya       m
                 auxiliary       sym           a
- -------------- -------------- -------------- -
2 VForm          indicative      govorya       i
                 imperative      govorete      m
                 participle      govoril       p
                 gerund          govorejki     g
- -------------- -------------- -------------- -
3 Tense          present         govorya       p
                 imperfect       govoreh       i
                 past            govoreno      s
           l.s.  aorist          govorih       a
- -------------- -------------- -------------- -
4 Person         first           govorya       1
                 second          govorish      2
                 third           govori        3
- -------------- -------------- -------------- -
5 Number         singular        govorya       s
                 plural          govoryat      p

- -------------- -------------- -------------- -
6 Gender         masculine       govoril       m
                 feminine        govorila      f
                 neuter          govorilo      n
************************************************
7 Voice          active          govorest      a
                 passive         govoreno      p
- -------------- -------------- -------------- -
8 Negative                                     -
- -------------- -------------- -------------- -
9 Definiteness   no             govoril        n
                 yes            govorilite     y
                 short_art      govoriliya     s
                 full_art       govoriliyat    f
- -------------- -------------- -------------- -
10Clitic                                       -
- -------------- -------------- -------------- -
11Case                                         -
- -------------- -------------- -------------- -
12Animate                                      -
- -------------- -------------- -------------- -
13Clitic_s                                     -
============================================== =

Combinations
=== ===== ==== ==== ==== ===== ====  === === === === ==== == === =======
PoS  Type VFrm Tens Pers Numb  Gend Voic Neg Def Cl1 Case An Cl2 Example
=== ===== ==== ==== ==== ===== ====  === === === === ==== == === =======
V    [ma]  i   [pai][123] [sp]   -    -  -    -   -   -    -  -    1.
V    [ma]  m    -     2   [sp]   -    -  -    -   -   -    -  -    2.
V     m    p   [pai]  -    s    [mfn] a  -    n   -   -    -  -    3.
V     m    p   [pa]   -    s     m    a  -   [sf] -   -    -  -    4.
V     m    p   [pa]   -    s    [fn]  a  -    y   -   -    -  -    5.
V     m    p   [pai]  -    p     -    a  -    n   -   -    -  -    6.
V     m    p   [pa]   -    p     -    a  -    y   -   -    -  -    7.
V     m    p    s     -    s    [mfn] p  -    n   -   -    -  -    8.
V     m    p    s     -    s     m    p  -   [sf] -   -    -  -    9.
V     m    p    s     -    s    [fn]  p  -    y   -   -    -  -   10.
V     m    p    s     -    p     -    p  -   [ny] -   -    -  -   11.
V     a    p   [ai]   -    s    [mfn] a  -    n   -   -    -  -   12.
V     a    p   [a]    -    s     m    a  -   [sf] -   -    -  -   13.
V     a    p   [a]    -    s    [fn]  a  -    y   -   -    -  -   14.
V     a    p   [ai]   -    p     -    a  -    n   -   -    -  -   15.
V     a    p   [a]    -    p     -    a  -    y   -   -    -  -   16.
V    [ma]  g    -     -    -     -    -  -    -   -   -    -  -   17.
=== ===== ==== ==== ==== ===== ====  === === === === ==== == === =======

Examples:
 1. govorya, sym
 2. govori, bydete

 3. govorest, govorila
 4. govorestiya, govorestiyat
 5. govorestta, govoriloto
 6. govoresti, govoreli
 7. govoresti, govorilite
 8. govoren, govorena
 9. govoreniya, govoreniyat
10. govorenata, govorenoto
11. govoreni, govorenite
12. bil, bila
13. biliya, biliyat
14. bilata, biloto
15. bili
16. bilite
17. govorejki, bidejki

8.3 Adjective (A)

= ============== ============== =============== ==
P ATT            VAL             Example        CC
= ============== ============== =============== ==
1 Type                             -             -
- -------------- -------------- --------------- --
2 Degree                           -             -
- -------------- -------------- --------------- --
3 Gender         masculine       bedniyat        m
                 feminine        bedna           f
                 neuter          bednoto         n
- -------------- -------------- --------------- --
4 Number         singular        bedniya         s
                 plural          bednite         p
- -------------- -------------- --------------- --
5 Case                                           -
* ************************************************
6 Definiteness   no              bedna           n
                 yes             bednata         y
           l.s.  short_art       bedniya         s
           l.s.  full_art        bedniyat        f
- -------------- -------------- --------------- --
7 Clitic                                         -
- -------------- -------------- --------------- --
8 Animate                                        -
- -------------- -------------- --------------- --
9 Formation                                      -
- -------------- -------------- --------------- --
10Owner_Number                                   -
- -------------- -------------- --------------- --
11Owner_Person                                   -
---------------- -------------- --------------- --
12Owned_Number                                   -

==================================================

Combinations
=== ==== ==== ==== ==== ==== === ==== === === === === === =======
PoS Type Degr Gend Numb Case Def Clit  An Frm OnN OnP OdN Example
=== ==== ==== ==== ==== ==== === ==== === === === === === =======
A    -    -   [mfn] s   -    n   -    -   -   -   -   -  beden, bedna
A    -    -    m    s   -   [sf] -    -   -   -   -   -  bedniya, bedniyat
A    -    -   [fn]  s   -    y   -    -   -   -   -   -  bednata, bednoto
A    -    -    -    p   -   [ny] -    -   -   -   -   -  bedni, bednite
=== ==== ==== ==== ==== ==== === ==== === === === === === =======

8.4 Pronoun (P)

= ============== ============== =========== =
P ATT            VAL            Example     C
= ============== ============== =========== =
1 Type           personal        az         p
                 demonstrative   tozi       d
                 indefinite      nyakoj     i
                 possessive      moj        s
                 interrogative   koi        q
                 relative        kojto      r
                 reflexive       sebe       x
                 negative        nikoj      z
                 general         vseki      g
- -------------- -------------- ----------- -
2 Person         first           az         1
                 second          ti         2
                 third           toj        3
- -------------- -------------- ----------- -
3 Gender         masculine       moj        m
                 feminine        moya       f
                 neuter          moe        n
- -------------- -------------- ----------- -
4 Number         singular        az         s
                 plural          nie        p
- -------------- -------------- ----------- -
5 Case           nominative      toj        n
                 dative          nemu       d
                 accusative      nego       a
- -------------- -------------- ----------- -
6 Owner_Number                              -
- -------------- -------------- ----------- -
7 Owner_Gender                              -
*********************************************
8 Clitic         no              sebe_si    n
                 yes             mi         y

- -------------- -------------- ----------- -
9 Referent_Type  personal        koj        p
                 possessive      tchij      s
                 attributive     kakva      a
                 quantitative    kolko      q
- -------------- -------------- ----------- -
10Syntactic_Type                            -
- -------------- -------------- ----------- -
11Definiteness   no              svoj       n
                 yes             svoyata    y
                 short_art       svoya      s
                 full_art        svoyat     f
- -------------- -------------- ----------- -
12Animate                                   -
- -------------- -------------- ----------- -
13Clitic_s                                  -
- -------------- -------------- ----------- -
14Pronoun_Form                              -
- -------------- -------------- ----------- -
15Owner_Person                              -
- -------------- -------------- ----------- -
16Owned_Number                              -
=============================================

Combinations
Personal pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    p  [12]   -    s   n    -   -   -   -    -     -   1.
P    p   3    [mfn] s   n    -   -   -   -    -     -   2.
P    p  [123]  -    p   n    -   -   -   -    -     -   3.
P    p  [12]   -    s  [ad]  -   -  [yn] -    -     -   4.
P    p   3    [mnf] s  [ad]  -   -  [yn] -    -     -   5.
P    p  [123]  -    p  [ad]  -   -  [yn] -    -     -   6.
Examples:
1. az, ti
2. toj, tya, to
3. nie, vie, te
4. mene, me, mi, mene_me
5. nego, go, nego_go
6. nas, ni, nas_ni, nam

Demonstrative pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    d    -  [mfn]  s   -    -   -   -   p    -    -   tozi,tova
P    d    -   -     p   -    -   -   -   p    -    -   tezi

P    d    -   -     -   -    -   -   -   q    -    -   tolkova

Indefinite, Interrogative, Relative, Negative pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P [iqrz]  -    m    s  [nad] -   -   -   p    -   -     1.
P [iqrz]  -    m    s   -    -   -   -  [as]  -   -     2.
P [iqrz]  -   [fn]  s   -    -   -   -  [pas] -   -     3.
P [iqrz]  -    -    p   -    -   -   -  [pas] -   -     4.
P [iqrz]  -    -    -   -    -   -   -   q    -   -     5.
Examples:
1. nyakoj, nyakogo, nyakomu, nikoj
2. nyakakyv, netchij, nikakyv
3. nyakoya, nyakakva, nechiya
4. nyakoi, nyakavi, netchii
5. nyakolko, nikolko, kolko

Possessive pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    s   [123]  m   s   -   -    -   n   -    -  [nsf] 1.
P    s   [123] [fn] s   -   -    -   n   -    -  [ny]  2.
P    s   [123]  -   p   -   -    -   n   -    -  [ny]  3.
P    s   [123]  -  [sp] -   -    -   y   -    -   -    4.
Examples:
1. moj, moya, moyat
2. moya, moyata, tvoya
3. moi, moite, tvoi
4. mi, ti, mu, j, ni, vi, im

Reflexive pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    x    -    -    -  [ad]  -   -  [ny] p    -    -    1.
P    x    -    m    s   -    -   -    n  s    -  [nsf]  2.
P    x    -   [fn]  s   -    -   -    n  s    -  [ny]   3.
P    x    -    -    p   -    -   -    n  s    -  [ny]   4.
P    x    -    -    -   -    -   -    y  s    -    -    5.
Examples:
1. sebe, se, si, sebe_si
2. svoj, svoyat, svoya
3. svoya, svoyata
4. svoi, svoite

5. si

General pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    g    -    m    s  [nad] -   -   -   p    -    -    1.
P    g    -    m    s    -   -   -   -   a    -    -    2.
P    g    -   [fn]  s    -   -   -   -  [pa]  -    -    3.
P    g    -    -    p    -   -   -   -  [pa]  -    -    4.
P    g    -    m    s    -   -   -   -   q    -   [sf]  5.
P    g    -   [fn]  s    -   -   -   -   q    -   [ny]  6.
P    g    -    -    p    -   -   -   -   q    -   [ny]  7.
Examples:
1. vseki, vsekigo, vsekimu
2. vsyakakyv
3. vsyaka, vsyakakva
4. vsyakoi, vsyakakvi
5. vsitchkiya, vsitchikiyat
6. vsitchka, vsitchkata
7. vsitchki, vsitchkite

8.5 Determiner (D)

Not applicable.

8.6 Article (T)

Not applicable.

8.7 Adverb (R)

= ============== ============== ============ =
P ATT            VAL             Example     C
= ============== ============== ============ =
1 Type           general          tuk        g
           l.s.  adjectival       umno       a
- -------------- --------------- ----------- -
2 Degree                                     -
**********************************************
3 Clitic                                     -
- -------------- -------------- ------------ -
4 Number                                     -

- -------------- -------------- ------------ -
5 Person                                     -
==============================================

Combinations
===== ====== ==== ==== ==== ===== ======================
POS    Type  Degr Clct Numb Pers  Example
===== ====== ==== ==== ==== ===== ======================
R        g    -    -    -    -     tuk, mnogo
R        a    -    -    -    -     umno, veselo, studeno
========================================================
Note: Adverbs of type adjectival have the same form as adjectives in
Gender = neuter, Person = 3, Number = singular.

8.8 Adposition (S)

= ============== ============== ============= ==
P ATT            VAL             Example       C
= ============== ============== ============= ==
1 Type           preposition     na, v         p
- -------------- -------------- ------------- --
2 Formation                                    -
************************************************
3 Case                                         -
- -------------- -------------- ------------- --
4 Clitic                                       -
================================================

Combinations
=== ===== ==== ==== ==== ============================
POS Type  Form Case Cltc  Examples
=== ===== ==== ==== ==== ============================
S    p     -    -    -    na, za, v, okolo, spored
=== ===== ===========================================

8.9 Conjunction (C)

= ============== ============== ============== ==
P ATT            VAL            Example         C
= ============== ============== ============== ==
1 Type           coordinating   a, ala, ili     c
                 subordinating  tche, da        s
*************************************************

2 Formation      simple         a, i, tche      s
                 compound       i_da, za_da     c
- -------------- -------------- -------------- --
3 Coord_Type                                    -
- -------------- -------------- -------------- --
4 Sub_Type                                      -
- -------------- -------------- -------------- --
5 Clitic                                        -
- -------------- -------------- -------------- --
6 Number                                        -
- -------------- -------------- -------------- --
7 Person                                        -
=================================================

Combinations
=== ==== ==== ==== ==== ==== ==== ===== ==========================
PoS Type Form CTyp SubT Clit Numb Pers  Examples
=== ==== ==== ==== ==== ==== ==== ===== ==========================
C    c   [sc]  -    -     -   -    -    i, a, ami, tche, tyj_tche
C    s   [sc]  -    -     -   -    -    tche, da, kato, za_da
=== ==== ==== ==== ==== ==== ==== ===== ==========================

8.10 Numeral (M)

= ============== ============== ============ =
P ATT            VAL             Example     C
= ============== ============== ============ =
1 Type           cardinal         edin       c
                 ordinal          vtori      o
- -------------- -------------- ------------ -
2 Gender         masculine        edin       m
                 feminine         edna       f
                 neuter           edno       n
- -------------- -------------- ------------ -
3 Number         singular         edin       s
                 plural           edni       p
- -------------- -------------- ------------ -
4 Case                            -          -
**********************************************
5 Form           digit            1984       d
                 roman            IX         r
                 letter           dva        l
           l.s.  m_form           dvama      m
           l.s.  approx           stotina    a
- -------------- -------------- ------------ -
6 Definiteness   no               edin       n
                 yes              ednata     y
                 short_art        ediniya    s

                 full_art         ediniyat   f
- -------------- -------------- ------------ -
7 Clitic                                     -
- -------------- -------------- ------------ -
8 Class                                      -
- -------------- -------------- ------------ -
9 Animate                                    -
- -------------- -------------- ------------ -
10Owner_Number                               -
- -------------- -------------- ------------ -
11Owner_Person                               -
- -------------- -------------- ------------ -
12Owned_Number                               -
===============================================

Combinations
=== ===== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ===========
PoS Type  Gend Numb Case Frm  Def  Clit Clas Anim OwnN OwnP OwdN   Example
=== ===== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ===========
Cardinal
M     c    -    -    -    d    -    -    -    -    -    -    -       1.
M     c    -    -    -    r    -    -    -    -    -    -    -       2.
M     c    m    s    -    l   [nsf] -    -    -    -    -    -       3.
M     c   [fn]  s    -    l   [ny]  -    -    -    -    -    -       4.
M     c    -    p    -    l   [ny]  -    -    -    -    -    -       5.
M     c   [mfn] p    -    l   [ny]  -    -    -    -    -    -       6.
M     c    m    p    -    m   [ny]  -    -    -    -    -    -       7.
M     c    -    p    -    l   [ny]  -    -    -    -    -    -       8.
M     c    -    p    -    a   [ny]  -    -    -    -    -    -       9.
Ordinal
M     o    m    s    -    -   [nsf] -    -    -    -    -    -      10.
M     o   [fn]  s    -    -   [ny]  -    -    -    -    -    -      11.
M     o    -    p    -    -   [ny]  -    -    -    -    -    -      12.
=== ===== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ===========
Examples:
1. 1984
2. IX
3. edin, ediniya, ediniyat
4. edna, ednata, edno, ednoto
5. edni, ednite
6. dva, dvata, dve, dvete
7. dvama, dvamata, trima
8. tri, trite, chetiri
9. stotina, stotinata
10. pyrvi, pyrviya, pyrviyat
11. pyrva, pyrvata
12. pyrvi, pyrvite

8.11 Interjection (I)

= ============== ============== ============ ==
P ATT            VAL             Example      C
= ============== ============== ============ ==
1 Type                                        -
- -------------- -------------- ------------ --
2 Formation      simple          ah           s
                 compound        bozhe moj!   c
============================================ ==

Combinations
=== ==== ========== =================================================
POS Type Formation  Examples
=== ==== ========== =================================================
I    -    s         o, hm-m, ha-ha, ah-lele, uvi
I    -    c         diavol_da_go_vzeme, ama_rabota, bozhe_moj
=====================================================================

8.12 Residual (X)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
P ATT                           -
=================================

8.13 Abbreviation (Y)

= ============== ============== ============ =
P ATT            VAL             Example     C
= ============== ============== ============ =
1 Syntactic_Type                             -
- -------------- -------------- ------------ -
2 Gender
- -------------- -------------- ------------ -
3 Number
- -------------- -------------- ------------ -
4 Case
- -------------- -------------- ------------ -
5 Definiteness
==============================================
Note: As there are no Attributes for Abbreviations, this section
      is has no examples

Combinations
==== ==== ==== ==== === ===========================
PoS  SynT Gend Numb Case Def  Examples
==== ==== ==== ==== === ===========================
Y     -    -   -     -   -    t.n., vzh., lat., razg...
===================================================

8.14 Particle (Q)

= ============== ============== =========== ==
P ATT            VAL             Example     C
= ============== ============== =========== ==
1 Type           negative        ne, ni      z
                 general         a, be       g
                 comparative     po, naj     c
                 verbal          da, ste     v
                 interrogative   li, dali    q
                 modal           da, dano    o
- -------------- -------------- ----------- --
2 Formation      simple          a, ne       s
                 compound        hajde_de    c
- -------------- -------------- ----------- --
3 Clitic                                     -
==============================================

Combinations
=== ============ ======= =======================
POS  Type         Form   Cltc  Examples
=== ============ ======= =======================
Q    [go]         [sc]    -    a, be, neka,hajde_de
Q    [cvzq]        s      -    po, naj, ne, li, nali
=== ============ ======= =======================
Note: In Bulgarian superlative particles  "po" and "naj" precede
      the superlative Forms of adjectives and adverbs:
      "po-hubav", "naj-hubav", "po-bavno", "naj-bavno".

Application to Estonian

Authors: H.J. Kaalep.

Notes:

1. Two additional tables are provided for every POS: one for definitions and one for examples. For nouns, adjectives, numerals and pronouns the table of examples contains more rows than the table of definitions. This is because in Estonian there are two equal ways for forming 11–12 plural cases for these POS and both are given.

2. The ordering of attributes is the same as in the synoptic tables of the proposal.

9.1 Nouns (N)

1.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           common         c jalg
                 proper         p London
- -------------- -------------- - ------------
2 Gender                        -
- -------------- -------------- - ------------
3 Number         singular       s jalg
                 plural         p jalad
- -------------- -------------- - ------------
4 Case           nominative     n jalg
                 genitive       g jala
           l.s.  partitive      1 jalga
                 illative       x jalasse
                 inessive       2 jalas
                 elative        e jalast
                 allative       t jalale

Multext-East / Concede  Morphosyntactic specifications V.2— Estonian  139

                 adessive       3 jalal
                 ablative       b jalalt
           l.s.  translative    4 jalaks
                 terminative    9 jalani
                 essive         w jalana
           l.s.  abessive       5 jalata
           l.s.  komitative     k jalaga
           l.s.  aditive        7 jalga
* ***************************** * ************
5 Definiteness                  -
- -------------- -------------- - ------------
6 Clitic                        -
- -------------- -------------- - ------------
7 Animate                       -
- -------------- -------------- - ------------
8 Owner_Number                  -
- -------------- -------------- - ------------
9 Owner_Person                  -
---------------- -------------- - ------------
10Owned_Number                  -
================================= ============

1.2. Definitions
Corpus tags       Lexicon       Definitions
---------------------------------------------------------
NCSX    Nc      Noun common
NCSN    Nc-sn   Noun common singular nominative
NCSG    Nc-sg   Noun common singular genitive
NCS1    Nc-s1   Noun common singular partitive
NCSA    Nc-sx   Noun common singular illative
NCSA    Nc-s2   Noun common singular inessive
NCSA    Nc-se   Noun common singular elative
NCSA    Nc-st   Noun common singular allative
NCSA    Nc-s3   Noun common singular adessive
NCSA    Nc-sb   Noun common singular ablative
NCS     Nc-s4   Noun common singular translative
NCS     Nc-s9   Noun common singular terminative
NCS     Nc-sw   Noun common singular essive
NCS     Nc-s5   Noun common singular abessive
NCS     Nc-sk   Noun common singular komitative
NCSA    Nc-s7   Noun common singular aditive
NCSN    Nc-pn   Noun common plural nominative
NCSG    Nc-pg   Noun common plural genitive
NCS1    Nc-p1   Noun common plural partitive
NCSA    Nc-px   Noun common plural illative
NCSA    Nc-p2   Noun common plural inessive
NCSA    Nc-pe   Noun common plural elative
NCSA    Nc-pt   Noun common plural allative
NCSA    Nc-p3   Noun common plural adessive
NCSA    Nc-pb   Noun common plural ablative
NCS     Nc-p4   Noun common plural translative

NCS     Nc-p9   Noun common plural terminative
NCS     Nc-pw   Noun common plural essive
NCS     Nc-p5   Noun common plural abessive
NCS     Nc-pk   Noun common plural komitative
NPSN    Np      Noun proper
NPSN    Np-sn   Noun proper singular nominative
NPSG    Np-sg   Noun proper singular genitive
NPS1    Np-s1   Noun proper singular partitive
NPSA    Np-sx   Noun proper singular illative
NPSA    Np-s2   Noun proper singular inessive
NPSA    Np-se   Noun proper singular elative
NPSA    Np-st   Noun proper singular allative
NPSA    Np-s3   Noun proper singular adessive
NPSA    Np-sb   Noun proper singular ablative
NPS     Np-s4   Noun proper singular translative
NPS     Np-s9   Noun proper singular terminative
NPS     Np-sw   Noun proper singular essive
NPS     Np-s5   Noun proper singular abessive
NPS     Np-sk   Noun proper singular komitative
NPSA    Np-s7   Noun proper singular aditive
NPSN    Np-pn   Noun proper plural nominative
NPSG    Np-pg   Noun proper plural genitive
NPS1    Np-p1   Noun proper plural partitive
NPSA    Np-px   Noun proper plural illative
NPSA    Np-p2   Noun proper plural inessive
NPSA    Np-pe   Noun proper plural elative
NPSA    Np-pt   Noun proper plural allative
NPSA    Np-p3   Noun proper plural adessive
NPSA    Np-pb   Noun proper plural ablative
NPS     Np-p4   Noun proper plural translative
NPS     Np-p9   Noun proper plural terminative
NPS     Np-pw   Noun proper plural essive
NPS     Np-p5   Noun proper plural abessive
NPS     Np-pk   Noun proper plural komitative

1.3. Examples
Corpus tags      Lexicon       Examples
---------------------------------------------
NCSX    Nc      kesk-
NCSN    Nc-sn   jalg
NCSG    Nc-sg   jala
NCS1    Nc-s1   jalga
NCSA    Nc-sx   jalasse
NCSA    Nc-s2   jalas
NCSA    Nc-se   jalast
NCSA    Nc-st   jalale
NCSA    Nc-s3   jalal
NCSA    Nc-sb   jalalt
NCS     Nc-s4   jalaks
NCS     Nc-s9   jalani

NCS     Nc-sw   jalana
NCS     Nc-s5   jalata
NCS     Nc-sk   jalaga
NCSA    Nc-s7   jalga
NCSN    Nc-pn   jalad,
NCSG    Nc-pg   jalgade,
NCS1    Nc-p1   jalgasid, jalgu
NCSA    Nc-px   jalgadesse, jalusse
NCSA    Nc-p2   jalgades, jalus
NCSA    Nc-pe   jalgadest, jalust
NCSA    Nc-pt   jalgadele, jalule
NCSA    Nc-p3   jalgadel, jalul
NCSA    Nc-pb   jalgadelt, jalult
NCS     Nc-p4   jalgadeks, jaluks
NCS     Nc-p9   jalgadeni
NCS     Nc-pw   jalgadena
NCS     Nc-p5   jalgadeta
NCS     Nc-pk   jalgadega
NPSN    Np      Kesk-
NPSN    Np-sn   Julia
NPSG    Np-sg   Julia
NPS1    Np-s1   Juliat
NPSA    Np-sx   Juliasse
NPSA    Np-s2   Julias
NPSA    Np-se   Juliast
NPSA    Np-st   Juliale
NPSA    Np-s3   Julial
NPSA    Np-sb   Julialt
NPS     Np-s4   Juliaks
NPS     Np-s9   Juliani
NPS     Np-sw   Juliana
NPS     Np-s5   Juliata
NPS     Np-sk   gJuliaga
NPSA    Np-s7   Londoni
NPSN    Np-pn   Juliad,
NPSG    Np-pg   Juliate,
NPS1    Np-p1   Juliaid,
NPSA    Np-px   Juliatesse, Juliaisse
NPSA    Np-p2   Juliates, Juliais
NPSA    Np-pe   Juliatest, Juliaist
NPSA    Np-pt   Juliatele, Juliaile
NPSA    Np-p3   Juliatel, Juliail
NPSA    Np-pb   Juliatelt, Juliailt
NPS     Np-p4   Juliateks, Juliaiks
NPS     Np-p9   Juliateni
NPS     Np-pw   Juliatena
NPS     Np-p5   Juliateta
NPS     Np-pk   Juliatega

9.2 Verbs (V)

2.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           main           m lugema
                 auxiliary      a olema
                 modal          o saama
- -------------- -------------- - ------------
2 VForm          indicative     i loen
                 imperative     m loe
                 conditional    c loeksin
                 infinitive     n lugeda
                 participle     p lugev
                 gerund         g lugedes
                 supine         u lugema
           l.s.  quotative      q lugevat
- -------------- -------------- - ------------
3 Tense          present        p loen
                 imperfect      i lugesin
                 past           s lugenuvat
- -------------- -------------- - ------------
4 Person         first          1 loen
                 second         2 loed
                 third          3 loeb
- -------------- -------------- - ------------
5 Number         singular       s loen
                 plural         p loeme
- -------------- -------------- - ------------
6 Gender                        -
********************************* ************
7 Voice          active         a loen
                 passive        p loetakse
- -------------- -------------- - ------------
8 Negative       no             n loen
                 yes            y (ei) + loeta
- -------------- -------------- - ------------
9 Definiteness                  -
- -------------- -------------- - ------------
10Clitic                        -
- -------------- -------------- - ------------
11Case
                 illative       x lugema
                 inessive       2 lugemas
                 elative        e lugemast
                 translative    4 lugemaks
                 abessive       5 lugemata
- -------------- -------------- - ------------
12Animate                       -

- -------------- -------------- - ------------
13Clitic_s                      -
================================= ============

2.2. Definitions
Corpus tags       Lexicon       Definitions
------------------------------------------------------------------------------------
VMS     Vm------n       Verb main afirmative
VM1     Vmip1s-an       Verb main indic present first singular active afirmative
VM2     Vmip2s-an       Verb main indic present second singular active afirmative
VM3     Vmip3s-an       Verb main indic present third singular active afirmative
VM1     Vmip1p-an       Verb main indic present first plural active afirmative
VM2     Vmip2p-an       Verb main indic present second plural active afirmative
VM3     Vmip3p-an       Verb main indic present third plural active afirmative
VMK     Vmip---ay       Verb main indic present active negative
VMP     Vmip---pn       Verb main indic present passive afirmative
VMN     Vmip---py       Verb main indic present passive negative
VM1     Vmii1s-an       Verb main indic imperfect first singular active afirmative
VM2     Vmii2s-an       Verb main indic imperfect second singular active afirmative
VM3     Vmii3s-an       Verb main indic imperfect third singular active afirmative
VM1     Vmii1p-an       Verb main indic imperfect first plural active afirmative
VM2     Vmii2p-an       Verb main indic imperfect second plural active afirmative
VM3     Vmii3p-an       Verb main indic imperfect third plural active afirmative
VMAS    Vmii---ay       Verb main indic imperfect active negative
VMP     Vmii---pn       Verb main indic imperfect passive afirmative
VMK     Vmmp2s-an       Verb main imper present second singular active afirmative
VMK     Vmmp2s-ay       Verb main imper present second singular active negative
VMK     Vmip---ay       Verb main indic present active negative
VM3     Vmmp3s-an       Verb main imper present third singular active afirmative
VMK     Vmmp1p-an       Verb main imper present first plural active afirmative
VMK     Vmmp2p-an       Verb main imper present second plural active afirmative
VM3     Vmmp3p-an       Verb main imper present third plural active afirmative
VMP     Vmmp---pn       Verb main imper present passive afirmative
VM3     Vmmp3s-ay       Verb main imper present third singular active negative
VMK     Vmmp1p-ay       Verb main imper present first plural active negative
VMK     Vmmp2p-ay       Verb main imper present second plural active negative
VM3     Vmmp3p-ay       Verb main imper present third plural active negative
VMP     Vmmp---py       Verb main imper present passive negative
VMS     Vmcp1s-an       Verb main condit present first singular active afirmative
VMS     Vmcp2s-an       Verb main condit present second singular active afirmative
VMS     Vmcp1s-an       Verb main condit present first singular active afirmative
VMS     Vmcp2s-an       Verb main condit present second singular active afirmative
VMS     Vmcp3s-an       Verb main condit present third singular active afirmative
VMS     Vmcp1p-an       Verb main condit present first plural active afirmative
VMS     Vmcp2p-an       Verb main condit present second plural active afirmative
VMS     Vmcp3p-an       Verb main condit present third plural active afirmative
VMS     Vmcp1p-an       Verb main condit present first plural active afirmative
VMS     Vmcp2p-an       Verb main condit present second plural active afirmative
VMS     Vmcp3p-an       Verb main condit present third plural active afirmative
VMS     Vmcp---ay       Verb main condit present active negative
VMS     Vmcp---ay       Verb main condit present active negative

VMP     Vmcp---pn       Verb main condit present passive afirmative
VMS     Vmcs1s-anVerb main condit past first singular active afirmative
VMS     Vmcs2s-an       Verb main condit past second singular active afirmative
VMS     Vmcs1s-an       Verb main condit past first singular active afirmative
VMS     Vmcs2s-an       Verb main condit past second singular active afirmative
VMS     Vmcs3s-an       Verb main condit past third singular active afirmative
VMS     Vmcs1p-an       Verb main condit past first plural active afirmative
VMS     Vmcs2p-an       Verb main condit past second plural active afirmative
VMS     Vmcs3p-an       Verb main condit past third plural active afirmative
VMS     Vmcs1p-an       Verb main condit past first plural active afirmative
VMS     Vmcs2p-an       Verb main condit past second plural active afirmative
VMS     Vmcs3p-an       Verb main condit past third plural active afirmative
VMS     Vmcs---ay       Verb main condit past active negative
VMS     Vmcs---ay       Verb main condit past active negative
VMP     Vmcs---pn       Verb main condit past passive afirmative
VMP     Vmcp---py       Verb main condit present passive negative
VMP     Vmcs---py       Verb main condit past passive negative
VMD     Vmn     Verb main infinit
VMAP    Vmpp---a        Verb main partic present active
VMAP    Vmpp---p        Verb main partic present passive
VMAS    Vmps---a        Verb main partic past active
VMAS    Vmps---p        Verb main partic past passive
VMAS    Vmii---ay       Verb main indic imperfect active negative
VMAS    Vmii---py       Verb main indic imperfect passive negative
VMG     Vmg     Verb main gerund
VMM     Vmu----a---x    Verb main supine active illative
VMM     Vmu----a---2    Verb main supine active inessive
VMM     Vmu----a---e    Verb main supine active elative
VMG     Vmu----a---4    Verb main supine active translative
VMASS   Vmu----a---5    Verb main supine active abessive
VMM     Vmu----p        Verb main supine passive
VMQ     Vmqp---an       Verb main quotat present active afirmative
VMQ     Vmqp---ay       Verb main quotat present active negative
VMQ     Vmqp---pn       Verb main quotat present passive afirmative
VMQ     Vmqs---an       Verb main quotat past active afirmative
VMQ     Vmqs---pn       Verb main quotat past passive afirmative
VMQ     Vmqp---ay       Verb main quotat present active negative
VMQ     Vmqp---py       Verb main quotat present passive negative
VMQ     Vmqs---ay       Verb main quotat past active negative
VMQ     Vmqs---py       Verb main quotat past passive negative
VM1     Vaip1s-an       Verb aux indic present first singular active afirmative
VM2     Vaip2s-an       Verb aux indic present second singular active afirmative
VM3     Vaip3s-an       Verb aux indic present third singular active afirmative
VM1     Vaip1p-an       Verb aux indic present first plural active afirmative
VM2     Vaip2p-an       Verb aux indic present second plural active afirmative
VM3     Vaip3p-an       Verb aux indic present third plural active afirmative
VMK     Vaip---ay       Verb aux indic present active negative
VMP     Vaip---pn       Verb aux indic present passive afirmative
VMN     Vaip---py       Verb aux indic present passive negative
VM1     Vaii1s-an       Verb aux indic imperfect first singular active afirmative
VM2     Vaii2s-an       Verb aux indic imperfect second singular active afirmative
VM3     Vaii3s-an       Verb aux indic imperfect third singular active afirmative

VM1     Vaii1p-an       Verb aux indic imperfect first plural active afirmative
VM2     Vaii2p-an       Verb aux indic imperfect second plural active afirmative
VM3     Vaii3p-an       Verb aux indic imperfect third plural active afirmative
VMAS    Vaii---ay       Verb aux indic imperfect active negative
VMP     Vaii---pn       Verb aux indic imperfect passive afirmative
VMK     Vamp2s-an       Verb aux imper present second singular active afirmative
VMK     Vamp2s-ay       Verb aux imper present second singular active negative
VMK     Vaip---ay       Verb aux indic present active negative
VM3     Vamp3s-an       Verb aux imper present third singular active afirmative
VMK     Vamp1p-an       Verb aux imper present first plural active afirmative
VMK     Vamp2p-an       Verb aux imper present second plural active afirmative
VM3     Vamp3p-an       Verb aux imper present third plural active afirmative
VMP     Vamp---pn       Verb aux imper present passive afirmative
VMS     Vamp3s-ay       Verb aux imper present third singular active negative
VMS     Vamp1p-ay       Verb aux imper present first plural active negative
VMS     Vamp2p-ay       Verb aux imper present second plural active negative
VMS     Vamp3p-ay       Verb aux imper present third plural active negative
VMS     Vamp---py       Verb aux imper present passive negative
VMP     Vamp---py       Verb aux imper present passive negative
VMK     Vamp2s-ay       Verb aux imper present second singular active negative
VMS     Vamp3s-ay       Verb aux imper present third singular active negative
VMS     Vamp1p-ay       Verb aux imper present first plural active negative
VMS     Vamp2p-ay       Verb aux imper present second plural active negative
VMS     Vamp3p-ay       Verb aux imper present third plural active negative
VMS     Vacp1s-an       Verb aux condit present first singular active afirmative
VMS     Vacp2s-an       Verb aux condit present second singular active afirmative
VMS     Vacp1s-an       Verb aux condit present first singular active afirmative
VMS     Vacp2s-an       Verb aux condit present second singular active afirmative
VMS     Vacp3s-an       Verb aux condit present third singular active afirmative
VMS     Vacp1p-an       Verb aux condit present first plural active afirmative
VMS     Vacp2p-an       Verb aux condit present second plural active afirmative
VMS     Vacp3p-an       Verb aux condit present third plural active afirmative
VMS     Vacp1p-an       Verb aux condit present first plural active afirmative
VMS     Vacp2p-an       Verb aux condit present second plural active afirmative
VMS     Vacp3p-an       Verb aux condit present third plural active afirmative
VMS     Vacp---ay       Verb aux condit present active negative
VMS     Vacp---ay       Verb aux condit present active negative
VMP     Vacp---pn       Verb aux condit present passive afirmative
VMS     Vacs1s-an       Verb aux condit past first singular active afirmative
VMS     Vacs2s-an       Verb aux condit past second singular active afirmative
VMS     Vacs1s-an       Verb aux condit past first singular active afirmative
VMS     Vacs2s-an       Verb aux condit past second singular active afirmative
VMS     Vacs3s-an       Verb aux condit past third singular active afirmative
VMS     Vacs1p-an       Verb aux condit past first plural active afirmative
VMS     Vacs2p-an       Verb aux condit past second plural active afirmative
VMS     Vacs3p-an       Verb aux condit past third plural active afirmative
VMS     Vacs1p-an       Verb aux condit past first plural active afirmative
VMS     Vacs2p-an       Verb aux condit past second plural active afirmative
VMS     Vacs3p-an       Verb aux condit past third plural active afirmative
VMS     Vacs---ay       Verb aux condit past active negative
VMS     Vacs---ay       Verb aux condit past active negative
VMP     Vacs---pn       Verb aux condit past passive afirmative
VMP     Vacp---py       Verb aux condit present passive negative

VMP     Vacs---py       Verb aux condit past passive negative
VMD     Van     Verb aux infinit
VMAP    Vapp---a        Verb aux partic present active
VMAP    Vapp---p        Verb aux partic present passive
VMAS    Vaps---a        Verb aux partic past active
VMAS    Vaps---p        Verb aux partic past passive
VMAS    Vaii---ay       Verb aux indic imperfect active negative
VMAS    Vaii---py       Verb aux indic imperfect passive negative
VMG     Vag     Verb aux gerund
VMM     Vau----a---x    Verb aux supine active illative
VMM     Vau----a---2    Verb aux supine active inessive
VMM     Vau----a---e    Verb aux supine active elative
VMG     Vau----a---4    Verb aux supine active translative
VMASS   Vau----a---5    Verb aux supine active abessive
VMM     Vau----p        Verb aux supine passive
VMQ     Vaqp---an       Verb aux quotat present active afirmative
VMQ     Vaqp---ay       Verb aux quotat present active negative
VMQ     Vaqp---pn       Verb aux quotat present passive afirmative
VMQ     Vaqs---an       Verb aux quotat past active afirmative
VMQ     Vaqs---pn       Verb aux quotat past passive afirmative
VMQ     Vaqp---ay       Verb aux quotat present active negative
VMQ     Vaqp---py       Verb aux quotat present passive negative
VMQ     Vaqs---ay       Verb aux quotat past active negative
VMQ     Vaqs---py       Verb aux quotat past passive negative
VME     Va------y       Verb aux negative
VM1     Voip1s-an       Verb modal indic present first singular active afirmative
VM2     Voip2s-an       Verb modal indic present second singular active afirmative
VM3Voip3s-an    Verb modal indic present third singular active afirmative
VM1     Voip1p-an       Verb modal indic present first plural active afirmative
VM2     Voip2p-an       Verb modal indic present second plural active afirmative
VM3     Voip3p-an       Verb modal indic present third plural active afirmative
VMP     Voip---pn       Verb modal indic present passive afirmative
VMN     Voip---py       Verb modal indic present passive negative
VM1     Voii1s-an       Verb modal indic imperfect first singular active afirmative
VM2     Voii2s-an       Verb modal indic imperfect second singular active afirmative
VM3     Voii3s-an       Verb modal indic imperfect third singular active afirmative
VM1     Voii1p-an       Verb modal indic imperfect first plural active afirmative
VM2     Voii2p-an       Verb modal indic imperfect second plural active afirmative
VM3     Voii3p-an       Verb modal indic imperfect third plural active afirmative
VMP     Voii---pn       Verb modal indic imperfect passive afirmative
VMK     Vomp2s-an       Verb modal imper present second singular active afirmative
VMK     Vomp2s-ay       Verb modal imper present second singular active negative
VMK     Voip---ay       Verb modal indic present active negative
VM3     Vomp3s-an       Verb modal imper present third singular active afirmative
VMK     Vomp1p-an       Verb modal imper present first plural active afirmative
VMK     Vomp2p-an       Verb modal imper present second plural active afirmative
VM3     Vomp3p-an       Verb modal imper present third plural active afirmative
VMP     Vomp---pn       Verb modal imper present passive afirmative
VM3     Vomp3s-ay       Verb modal imper present third singular active negative
VMK     Vomp1p-ay       Verb modal imper present first plural active negative
VMK     Vomp2p-ay       Verb modal imper present second plural active negative
VM3     Vomp3p-ay       Verb modal imper present third plural active negative

VMP     Vomp---py       Verb modal imper present passive negative
VMS     Vocp1s-an       Verb modal condit present first singular active afirmative
VMS     Vocp2s-an       Verb modal condit present second singular active afirmative
VMS     Vocp1s-an       Verb modal condit present first singular active afirmative
VMS     Vocp2s-an       Verb modal condit present second singular active afirmative
VMS     Vocp3s-an       Verb modal condit present third singular active afirmative
VMS     Vocp1p-an       Verb modal condit present first plural active afirmative
VMS     Vocp2p-an       Verb modal condit present second plural active afirmative
VMS     Vocp3p-an       Verb modal condit present third plural active afirmative
VMS     Vocp1p-an       Verb modal condit present first plural active afirmative
VMS     Vocp2p-an       Verb modal condit present second plural active afirmative
VMS     Vocp3p-an       Verb modal condit present third plural active afirmative
VMS     Vocp---ay       Verb modal condit present active negative
VMP     Vocp---pn       Verb modal condit present passive afirmative
VMS     Vocs1s-an       Verb modal condit past first singular active afirmative
VMS     Vocs2s-an       Verb modal condit past second singular active afirmative
VMS     Vocs1s-an       Verb modal condit past first singular active afirmative
VMS     Vocs2s-an       Verb modal condit past second singular active afirmative
VMS     Vocs3s-an       Verb modal condit past third singular active afirmative
VMS     Vocs1p-an       Verb modal condit past first plural active afirmative
VMS     Vocs2p-an       Verb modal condit past second plural active afirmative
VMS     Vocs3p-an       Verb modal condit past third plural active afirmative
VMS     Vocs1p-an       Verb modal condit past first plural active afirmative
VMS     Vocs2p-an       Verb modal condit past second plural active afirmative
VMS     Vocs3p-an       Verb modal condit past third plural active afirmative
VMS     Vocs---ay       Verb modal condit past active negative
VMP     Vocs---pn       Verb modal condit past passive afirmative
VMP     Vocp---py       Verb modal condit present passive negative
VMP     Vocs---py       Verb modal condit past passive negative
VMD     Von     Verb modal infinit
VMAP    Vopp---a        Verb modal partic present active
VMAP    Vopp---p        Verb modal partic present passive
VMAS    Vops---a        Verb modal partic past active
VMAS    Vops---p        Verb modal partic past passive
VMAS    Voii---ay       Verb modal indic imperfect active negative
VMAS    Voii---py       Verb modal indic imperfect passive negative
VMG     Vog     Verb modal gerund
VMM     Vou----a---x    Verb modal supine active illative
VMM     Vou----a---2    Verb modal supine active inessive
VMM     Vou----a---e    Verb modal supine active elative
VMG     Vou----a---4    Verb modal supine active translative
VMASS   Vou----a---5    Verb modal supine active abessive
VMM     Vou----p        Verb modal supine passive
VMQ     Voqp---an       Verb modal quotat present active afirmative
VMQ     Voqp---pn       Verb modal quotat present passive afirmative
VMQ     Voqs---an       Verb modal quotat past active afirmative
VMQ     Voqs---pn       Verb modal quotat past passive afirmative
VMQ     Voqp---ay       Verb modal quotat present active negative
VMQ     Voqp---py       Verb modal quotat present passive negative
VMQ     Voqs---ay       Verb modal quotat past active negative
VMQ     Voqs---py       Verb modal quotat past passive negative

2.3. Examples
Corpus tags      Lexicon       Examples
-------------------------------------------
VMS     Vm------n       kuulukse
VM1     Vmip1s-an       loen
VM2     Vmip2s-an       loed
VM3     Vmip3s-an       loeb
VM1     Vmip1p-an       loeme
VM2     Vmip2p-an       loete
VM3     Vmip3p-an       loevad
VMK     Vmip---ay       pole
VMP     Vmip---pn       loetakse
VMN     Vmip---py       (ei)+loeta
VM1     Vmii1s-an       lugesin
VM2     Vmii2s-an       lugesid
VM3     Vmii3s-an       luges
VM1     Vmii1p-an       lugesime
VM2     Vmii2p-an       lugesite
VM3     Vmii3p-an       lugesid
VMAS    Vmii---ay       polnud
VMP     Vmii---pn       loeti
VMK     Vmmp2s-an       loe
VMK     Vmmp2s-ay       (ära) loe
VMK     Vmip---ay       (ei) loe
VM3     Vmmp3s-an       lugegu
VMK     Vmmp1p-an       lugegem
VMK     Vmmp2p-an       lugege
VM3     Vmmp3p-an       lugegu
VMP     Vmmp---pn       loetagu
VM3     Vmmp3s-ay       (ärgu) lugegu
VMK     Vmmp1p-ay       (ärgem) lugegem
VMK     Vmmp2p-ay       (ärge) lugege
VM3     Vmmp3p-ay       (ärgu) lugegu
VMP     Vmmp---py       (ärgu) loetagu
VMS     Vmcp1s-an       loeksin
VMS     Vmcp2s-an       loeksid
VMS     Vmcp1s-an       loeks
VMS     Vmcp2s-an       loeks
VMS     Vmcp3s-an       loeks
VMS     Vmcp1p-an       loeks
VMS     Vmcp2p-an       loeks
VMS     Vmcp3p-an       loeks
VMS     Vmcp1p-an       loeksime
VMS     Vmcp2p-an       loeksite
VMS     Vmcp3p-an       loeksid
VMS     Vmcp---ay       (ei) loeks
VMS     Vmcp---ay       poleks
VMP     Vmcp---pn       loetaks
VMS     Vmcs1s-an       lugenuksin
VMS     Vmcs2s-an       lugenuksid
VMS     Vmcs1s-an       lugenuks
VMS     Vmcs2s-an       lugenuks

VMS     Vmcs3s-an       lugenuks
VMS     Vmcs1p-an       lugenuks
VMS     Vmcs2p-an       lugenuks
VMS     Vmcs3p-an       lugenuks
VMS     Vmcs1p-an       lugenuksime
VMS     Vmcs2p-an       lugenuksite
VMS     Vmcs3p-an       lugenuksid
VMS     Vmcs---ay       (ei) lugenuks
VMS     Vmcs---ay       polnuks
VMP     Vmcs---pn       loetuks
VMP     Vmcp---py       (ei) loetaks
VMP     Vmcs---py       (ei) loetuks
VMD     Vmn     lugeda
VMAP    Vmpp---a        lugev
VMAP    Vmpp---p        loetav
VMAS    Vmps---a        lugenud
VMAS    Vmps---p        loetud
VMAS    Vmii---ay       (ei) lugenud
VMAS    Vmii---py       (ei) loetud
VMG     Vmg     lugedes
VMM     Vmu----a---x    lugema
VMM     Vmu----a---2    lugemas
VMM     Vmu----a---e    lugemast
VMG     Vmu----a---4    lugemaks
VMASS   Vmu----a---5    lugemata
VMM     Vmu----p        loetama
VMQ     Vmqp---an       lugevat
VMQ     Vmqp---ay       polevat
VMQ     Vmqp---pn       loetavat
VMQ     Vmqs---an       lugenuvat
VMQ     Vmqs---pn       loetuvat
VMQ     Vmqp---ay       (ei) lugevat
VMQ     Vmqp---py       (ei) loetavat
VMQ     Vmqs---ay       (ei) lugenuvat
VMQ     Vmqs---py       (ei) loetuvat
VM1     Vaip1s-an       olen
VM2     Vaip2s-an       oled
VM3     Vaip3s-an       on
VM1     Vaip1p-an       oleme
VM2     Vaip2p-an       olete
VM3     Vaip3p-an       on
VMK     Vaip---ay       pole
VMP     Vaip---pn       ollakse
VMN     Vaip---py       (ei)+olda
VM1     Vaii1s-an       olin
VM2     Vaii2s-an       olid
VM3     Vaii3s-an       oli
VM1     Vaii1p-an       olime
VM2     Vaii2p-an       olite
VM3     Vaii3p-an       olid
VMAS    Vaii---ay       polnud
VMP     Vaii---pn       oldi

VMK     Vamp2s-an       ole
VMK     Vamp2s-ay       (ära) ole
VMK     Vaip---ay       (ei) ole
VM3     Vamp3s-an       olgu
VMK     Vamp1p-an       olgem
VMK     Vamp2p-an       olge
VM3     Vamp3p-an       olgu
VMP     Vamp---pn       oldagu
VMS     Vamp3s-ay       (ärgu) olgu
VMS     Vamp1p-ay       (ärgem) olgem
VMS     Vamp2p-ay       (ärge) olge
VMS     Vamp3p-ay       (ärgu) olgu
VMS     Vamp---py       (ärgu) oldagu
VMP     Vamp---py       ärgu
VMK     Vamp2s-ay       ära
VMS     Vamp3s-ay       ärgu
VMS     Vamp1p-ay       ärgem
VMS     Vamp2p-ay       ärge
VMS     Vamp3p-ay       ärgu
VMS     Vacp1s-an       oleksin
VMS     Vacp2s-an       oleksid
VMS     Vacp1s-an       oleks
VMS     Vacp2s-an       oleks
VMS     Vacp3s-an       oleks
VMS     Vacp1p-an       oleks
VMS     Vacp2p-an       oleks
VMS     Vacp3p-an       oleks
VMS     Vacp1p-an       oleksime
VMS     Vacp2p-an       oleksite
VMS     Vacp3p-an       oleksid
VMS     Vacp---ay       (ei) oleks
VMS     Vacp---ay       poleks
VMP     Vacp---pn       oldaks
VMS     Vacs1s-an       olnuksin
VMS     Vacs2s-an       olnuksid
VMS     Vacs1s-an       olnuks
VMS     Vacs2s-an       olnuks
VMS     Vacs3s-an       olnuks
VMS     Vacs1p-an       olnuks
VMS     Vacs2p-an       olnuks
VMS     Vacs3p-an       olnuks
VMS     Vacs1p-an       olnuksime
VMS     Vacs2p-an       olnuksite
VMS     Vacs3p-an       olnuksid
VMS     Vacs---ay       (ei) olnuks
VMS     Vacs---ay       polnuks
VMP     Vacs---pn       olduks
VMP     Vacp---py       (ei) oldaks
VMP     Vacs---py       (ei) olduks
VMD     Van     olla
VMAP    Vapp---a        olev
VMAP    Vapp---p        oldav
VMAS    Vaps---a        olnud

VMAS    Vaps---p        oldud
VMAS    Vaii---ay       (ei) olnud
VMAS    Vaii---py       (ei) oldud
VMG     Vag     olles
VMM     Vau----a---x    olema
VMM     Vau----a---2    olemas
VMM     Vau----a---e    olemast
VMG     Vau----a---4    olemaks
VMASS   Vau----a---5    olemata
VMM     Vau----p        oldama
VMQ     Vaqp---an       olevat
VMQ     Vaqp---ay       polevat
VMQ     Vaqp---pn       oldavat
VMQ     Vaqs---an       olnuvat
VMQ     Vaqs---pn       olduvat
VMQ     Vaqp---ay       (ei) olevat
VMQ     Vaqp---py       (ei) oldavat
VMQ     Vaqs---ay       (ei) olnuvat
VMQ     Vaqs---py       (ei) olduvat
VME     Va------y       ei
VM1     Voip1s-an       saan
VM2     Voip2s-an       saad
VM3     Voip3s-an       saab
VM1     Voip1p-an       saame
VM2     Voip2p-an       saate
VM3     Voip3p-an       saavad
VMP     Voip---pn       saadakse
VMN     Voip---py       (ei)+saada
VM1     Voii1s-an       sain
VM2     Voii2s-an       said
VM3     Voii3s-an       sai
VM1     Voii1p-an       saime
VM2     Voii2p-an       saite
VM3     Voii3p-an       said
VMP     Voii---pn       saadi
VMK     Vomp2s-an       saa
VMK     Vomp2s-ay       (ära) saa
VMK     Voip---ay       (ei) saa
VM3     Vomp3s-an       saagu
VMK     Vomp1p-an       saagem
VMK     Vomp2p-an       saage
VM3     Vomp3p-an       saagu
VMP     Vomp---pn       saadagu
VM3     Vomp3s-ay       (ärgu) saagu
VMK     Vomp1p-ay       (ärgem) saagem
VMK     Vomp2p-ay       (ärge) saage
VM3     Vomp3p-ay       (ärgu) saagu
VMP     Vomp---py       (ärgu) saadagu
VMS     Vocp1s-an       saaksin
VMS     Vocp2s-an       saaksid
VMS     Vocp1s-an       saaks
VMS     Vocp2s-an       saaks

VMS     Vocp3s-an       saaks
VMS     Vocp1p-an       saaks
VMS     Vocp2p-an       saaks
VMS     Vocp3p-an       saaks
VMS     Vocp1p-an       saaksime
VMS     Vocp2p-an       saaksite
VMS     Vocp3p-an       saaksid
VMS     Vocp---ay       (ei) saaks
VMP     Vocp---pn       saadaks
VMS     Vocs1s-an       saanuksin
VMS     Vocs2s-an       saanuksid
VMS     Vocs1s-an       saanuks
VMS     Vocs2s-an       saanuks
VMS     Vocs3s-an       saanuks
VMS     Vocs1p-an       saanuks
VMS     Vocs2p-an       saanuks
VMS     Vocs3p-an       saanuks
VMS     Vocs1p-an       saanuksime
VMS     Vocs2p-an       saanuksite
VMS     Vocs3p-an       saanuksid
VMS     Vocs---ay       (ei) saanuks
VMP     Vocs---pn       saaduks
VMP     Vocp---py       (ei) saadaks
VMP     Vocs---py       (ei) saaduks
VMD     Von     saada
VMAP    Vopp---a        saav
VMAP    Vopp---p        saadav
VMAS    Vops---a        saanud
VMAS    Vops---p        saadud
VMAS    Voii---ay       (ei) saanud
VMAS    Voii---py       (ei) saadud
VMG     Vog     saades
VMM     Vou----a---x    saama
VMM     Vou----a---2    saamas
VMM     Vou----a---e    saamast
VMG     Vou----a---4    saamaks
VMASS   Vou----a---5    saamata
VMM     Vou----p        saadama
VMQ     Voqp---an       saavat
VMQ     Voqp---pn       saadavat
VMQ     Voqs---an       saanuvat
VMQ     Voqs---pn       saaduvat
VMQ     Voqp---ay       (ei) saavat
VMQ     Voqp---py       (ei) saadavat
VMQ     Voqs---ay       (ei) saanuvat
VMQ     Voqs---py       (ei) saaduvat

9.3 Adjectives (A)

3.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type                          -
- -------------- -------------- - ------------
2 Degree         positive       p oluline
                 comparative    c olulisem
                 superlative    s olulisim
- -------------- -------------- - ------------
3 Gender                        -
- -------------- -------------- - ------------
4 Number         singular       s oluline
                 plural         p olulised
- -------------- -------------- - ------------
5 Case           nominative     n oluline
                 genitive       g olulise
           l.s.  partitive      1 olulist
                 illative       x olulisesse
                 inessive       2 olulises
                 elative        e olulisest
                 allative       t olulisele
                 adessive       3 olulisel
                 ablative       b oluliselt
           l.s.  translative    4 oluliseks
                 terminative    9 oluliseni
                 essive         w olulisena
           l.s.  abessive       5 oluliseta
           l.s.  komitative     k olulisega
           l.s.  aditive        7 olulisse
* ***************************** * ************
6 Definiteness                  -
- -------------- -------------- - ------------
7 Clitic                        -
- -------------- -------------- - ------------
8 Animate                       -
- -------------- -------------- - ------------
9 Formation                     -
- -------------- -------------- - ------------
10Owner_Number                  -
- -------------- -------------- - ------------
11Owner_Person                  -
---------------- -------------- - ------------
12Owned_Number                  -
================================= ============

3.2. Definitions

Corpus tags       Lexicon       Definitions
-------------------------------------------------------------------------
ASX     A-p     Adjective positive
ASN     A-p-sn  Adjective positive singular nominative
ASG     A-p-sg  Adjective positive singular genitive
AS1     A-p-s1  Adjective positive singular partitive
ASA     A-p-sx  Adjective positive singular illative
ASA     A-p-s2  Adjective positive singular inessive
ASA     A-p-se  Adjective positive singular elative
ASA     A-p-st  Adjective positive singular allative
ASA     A-p-s3  Adjective positive singular adessive
ASA     A-p-sb  Adjective positive singular ablative
AS      A-p-s4  Adjective positive singular translative
AS      A-p-s9  Adjective positive singular terminative
AS      A-p-sw  Adjective positive singular essive
AS      A-p-s5  Adjective positive singular abessive
AS      A-p-sk  Adjective positive singular komitative
ASA     A-p-s7  Adjective positive singular aditive
ASN     A-p-pn  Adjective positive plural nominative
ASG     A-p-pg  Adjective positive plural genitive
AS1     A-p-p1  Adjective positive plural partitive
ASA     A-p-px  Adjective positive plural illative
ASA     A-p-p2  Adjective positive plural inessive
ASA     A-p-pe  Adjective positive plural elative
ASA     A-p-pt  Adjective positive plural allative
ASA     A-p-p3  Adjective positive plural adessive
ASA     A-p-pb  Adjective positive plural ablative
AS      A-p-p4  Adjective positive plural translative
AS      A-p-p9  Adjective positive plural terminative
AS      A-p-pw  Adjective positive plural essive
AS      A-p-p5  Adjective positive plural abessive
AS      A-p-pk  Adjective positive plural komitative
ASN     A-c-sn  Adjective comparative singular nominative
ASG     A-c-sg  Adjective comparative singular genitive
AS1     A-c-s1  Adjective comparative singular partitive
ASA     A-c-sx  Adjective comparative singular illative
ASA     A-c-s2  Adjective comparative singular inessive
ASA     A-c-se  Adjective comparative singular elative
ASA     A-c-st  Adjective comparative singular allative
ASA     A-c-s3  Adjective comparative singular adessive
ASA     A-c-sb  Adjective comparative singular ablative
AS      A-c-s4  Adjective comparative singular translative
AS      A-c-s9  Adjective comparative singular terminative
AS      A-c-sw  Adjective comparative singular essive
AS      A-c-s5  Adjective comparative singular abessive
AS      A-c-sk  Adjective comparative singular komitative
ASN     A-c-pn  Adjective comparative plural nominative
ASG     A-c-pg  Adjective comparative plural genitive
AS1     A-c-p1  Adjective comparative plural partitive
ASA     A-c-px  Adjective comparative plural illative
ASA     A-c-p2  Adjective comparative plural inessive

ASA     A-c-pe  Adjective comparative plural elative
ASA     A-c-pt  Adjective comparative plural allative
ASA     A-c-p3  Adjective comparative plural adessive
ASA     A-c-pb  Adjective comparative plural ablative
AS      A-c-p4  Adjective comparative plural translative
AS      A-c-p9  Adjective comparative plural terminative
AS      A-c-pw  Adjective comparative plural essive
AS      A-c-p5  Adjective comparative plural abessive
AS      A-c-pk  Adjective comparative plural komitative
ASN     A-s-sn  Adjective superlative singular nominative
ASG     A-s-sg  Adjective superlative singular genitive
AS1     A-s-s1  Adjective superlative singular partitive
ASA     A-s-sx  Adjective superlative singular illative
ASA     A-s-s2  Adjective superlative singular inessive
ASA     A-s-se  Adjective superlative singular elative
ASA     A-s-st  Adjective superlative singular allative
ASA     A-s-s3  Adjective superlative singular adessive
ASA     A-s-sb  Adjective superlative singular ablative
AS      A-s-s4  Adjective superlative singular translative
AS      A-s-s9  Adjective superlative singular terminative
AS      A-s-sw  Adjective superlative singular essive
AS      A-s-s5  Adjective superlative singular abessive
AS      A-s-sk  Adjective superlative singular komitative
ASN     A-s-pn  Adjective superlative plural nominative
ASG     A-s-pg  Adjective superlative plural genitive
AS1     A-s-p1  Adjective superlative plural partitive
ASA     A-s-px  Adjective superlative plural illative
ASA     A-s-p2  Adjective superlative plural inessive
ASA     A-s-pe  Adjective superlative plural elative
ASA     A-s-pt  Adjective superlative plural allative
ASA     A-s-p3  Adjective superlative plural adessive
ASA     A-s-pb  Adjective superlative plural ablative
AS      A-s-p4  Adjective superlative plural translative
AS      A-s-p9  Adjective superlative plural terminative
AS      A-s-pw  Adjective superlative plural essive
AS      A-s-p5  Adjective superlative plural abessive
AS      A-s-pk  Adjective superlative plural komitative

3.3. Examples
Corpus tags      Lexicon       Examples
------------------------------------------------
ASX     A-p     salkus
ASN     A-p-sn  oluline
ASG     A-p-sg  olulise
AS1     A-p-s1  olulist
ASA     A-p-sx  olulisesse
ASA     A-p-s2  olulises
ASA     A-p-se  olulisest
ASA     A-p-st  olulisele
ASA     A-p-s3  olulisel

ASA     A-p-sb  oluliselt
AS      A-p-s4  oluliseks
AS      A-p-s9  oluliseni
AS      A-p-sw  olulisena
AS      A-p-s5  oluliseta
AS      A-p-sk  olulisega
ASA     A-p-s7  olulisse
ASN     A-p-pn  olulised
ASG     A-p-pg  oluliste
AS1     A-p-p1  olulisi
ASA     A-p-px  olulistesse, olulisisse
ASA     A-p-p2  olulistes, olulisis
ASA     A-p-pe  olulistest, olulisist
ASA     A-p-pt  olulistele, olulisile
ASA     A-p-p3  olulistel, olulisil
ASA     A-p-pb  olulistelt, olulisilt
AS      A-p-p4  olulisteks, olulisiks
AS      A-p-p9  olulisteni
AS      A-p-pw  olulistena
AS      A-p-p5  olulisteta
AS      A-p-pk  olulistega
ASN     A-c-sn  olulisem
ASG     A-c-sg  olulisema
AS1     A-c-s1  olulisemat
ASA     A-c-sx  olulisemasse
ASA     A-c-s2  olulisemas
ASA     A-c-se  olulisemast
ASA     A-c-st  olulisemale
ASA     A-c-s3  olulisemal
ASA     A-c-sb  olulisemalt
AS      A-c-s4  olulisemaks
AS      A-c-s9  olulisemani
AS      A-c-sw  olulisemana
AS      A-c-s5  olulisemata
AS      A-c-sk  olulisemaga
ASN     A-c-pn  olulisemad
ASG     A-c-pg  olulisemate
AS1     A-c-p1  olulisemaid
ASA     A-c-px  olulisematesse, olulisemaisse
ASA     A-c-p2  olulisemates, olulisemais
ASA     A-c-pe  olulisematest, olulisemaist
ASA     A-c-pt  olulisematele, olulisemaile
ASA     A-c-p3  olulisematel, olulisemail
ASA     A-c-pb  olulisematelt, olulisemailt
AS      A-c-p4  olulisemateks, olulisemaiks
AS      A-c-p9  olulisemateni, olulisemaini
AS      A-c-pw  olulisematena, olulisemaina
AS      A-c-p5  olulisemateta
AS      A-c-pk  olulisematega
ASN     A-s-sn  olulisim
ASG     A-s-sg  olulisima

AS1     A-s-s1  olulisimat
ASA     A-s-sx  olulisimasse
ASA     A-s-s2  olulisimas
ASA     A-s-se  olulisimast
ASA     A-s-st  olulisimale
ASA     A-s-s3  olulisimal
ASA     A-s-sb  olulisimalt
AS      A-s-s4  olulisimaks
AS      A-s-s9  olulisimani
AS      A-s-sw  olulisimana
AS      A-s-s5  olulisimata
AS      A-s-sk  olulisimaga
ASN     A-s-pn  olulisimad
ASG     A-s-pg  olulisimate
AS1     A-s-p1  olulisimaid
ASA     A-s-px  olulisimatesse, olulisimaisse
ASA     A-s-p2  olulisimates, olulisimais
ASA     A-s-pe  olulisimatest, olulisimaist
ASA     A-s-pt  olulisimatele, olulisimaile
ASA     A-s-p3  olulisimatel, olulisimail
ASA     A-s-pb  olulisimatelt, olulisimailt
AS      A-s-p4  olulisimateks, olulisimaiks
AS      A-s-p9  olulisimateni, olulisimaini
AS      A-s-pw  olulisimatena, olulisimaina
AS      A-s-p5  olulisimateta
AS      A-s-pk  olulisimatega

9.4 Pronouns (P)

4.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           personal       p mina
                 demonstrative  d selline
                 indefinite     i mingi
                 possessive     s iseenese
                 interrogative  q milline
                 relative       r milline
                 reflexive      x iseenda
                 reciprocal     y teineteise
           l.s.  determinal     m ise
- -------------- -------------- - ------------
2 Person         first          1 mina
                 second         2 sina
                 third          3 tema
- -------------- -------------- - ------------

3 Gender                        -
- -------------- -------------- - ------------
4 Number         singular       s mina
                 plural         p meie
- -------------- -------------- - ------------
5 Case           nominative     n mina
                 genitive       g minu
           l.s.  partitive      1 mind
                 illative       x minusse
                 inessive       2 minus
                 elative        e minust
                 allative       t minule
                 adessive       3 minul
                 ablative       b minult
           l.s.  translative    4 minuks
                 terminative    9 minuni
                 essive         w minuna
           l.s.  abessive       5 minuta
           l.s.  komitative     k minuga
           l.s.  aditive        7 sellisse
- -------------- -------------- - ------------
6 Owner_Number                  -
- -------------- -------------- - ------------
7 Owner_Gender                  -
********************************* ************
8 Clitic                        -
- -------------- -------------- - ------------
9 Referent_Type                 -
- -------------- -------------- - ------------
10Syntactic_Type                -
- -------------- -------------- - ------------
11Definiteness                  -
- -------------- -------------- - ------------
12Animate                       -
- -------------- -------------- - ------------
13Clitic_s                      -
- -------------- -------------- - ------------
14Pronoun_Form                  -
- -------------- -------------- - ------------
15Owner_Person                  -
- -------------- -------------- - ------------
16Owned_Number                  -
================================= ============

4.2. Definitions
Corpus tags       Lexicon       Definitions
------------------------------------------------------------------
PP1SN   Pp1-sn  Pronoun personal first singular nominative
PP1SG   Pp1-sg  Pronoun personal first singular genitive
PP1S1   Pp1-s1  Pronoun personal first singular partitive
PP1S    Pp1-sx  Pronoun personal first singular illative

PP1SA   Pp1-s2  Pronoun personal first singular inessive
PP1SA   Pp1-se  Pronoun personal first singular elative
PP1SA   Pp1-st  Pronoun personal first singular allative
PP1SA   Pp1-s3  Pronoun personal first singular adessive
PP1SA   Pp1-sb  Pronoun personal first singular ablative
PP1S    Pp1-s4  Pronoun personal first singular translative
PP1S    Pp1-s9  Pronoun personal first singular terminative
PP1S    Pp1-sw  Pronoun personal first singular essive
PP1S    Pp1-s5  Pronoun personal first singular abessive
PP1S    Pp1-sk  Pronoun personal first singular komitative
PP1SN   Pp1-pn  Pronoun personal first plural nominative
PP1SG   Pp1-pg  Pronoun personal first plural genitive
PP1S1   Pp1-p1  Pronoun personal first plural partitive
PP1S    Pp1-px  Pronoun personal first plural illative
PP1SA   Pp1-p2  Pronoun personal first plural inessive
PP1SA   Pp1-pe  Pronoun personal first plural elative
PP1SA   Pp1-pt  Pronoun personal first plural allative
PP1SA   Pp1-p3  Pronoun personal first plural adessive
PP1SA   Pp1-pb  Pronoun personal first plural ablative
PP1S    Pp1-p4  Pronoun personal first plural translative
PP1S    Pp1-p9  Pronoun personal first plural terminative
PP1S    Pp1-pw  Pronoun personal first plural essive
PP1S    Pp1-p5  Pronoun personal first plural abessive
PP1S    Pp1-pk  Pronoun personal first plural komitative
PP2SN   Pp2-sn  Pronoun personal second singular nominative
PP2SG   Pp2-sg  Pronoun personal second singular genitive
PP2S1   Pp2-s1  Pronoun personal second singular partitive
PP2S    Pp2-sx  Pronoun personal second singular illative
PP2SA   Pp2-s2  Pronoun personal second singular inessive
PP2SA   Pp2-se  Pronoun personal second singular elative
PP2SA   Pp2-st  Pronoun personal second singular allative
PP2SA   Pp2-s3  Pronoun personal second singular adessive
PP2SA   Pp2-sb  Pronoun personal second singular ablative
PP2S    Pp2-s4  Pronoun personal second singular translative
PP2S    Pp2-s9  Pronoun personal second singular terminative
PP2S    Pp2-sw  Pronoun personal second singular essive
PP2S    Pp2-s5  Pronoun personal second singular abessive
PP2S    Pp2-sk  Pronoun personal second singular komitative
PP2SN   Pp2-pn  Pronoun personal second plural nominative
PP2SG   Pp2-pg  Pronoun personal second plural genitive
PP2S1   Pp2-p1  Pronoun personal second plural partitive
PP2S    Pp2-px  Pronoun personal second plural illative
PP2SA   Pp2-p2  Pronoun personal second plural inessive
PP2SA   Pp2-pe  Pronoun personal second plural elative
PP2SA   Pp2-pt  Pronoun personal second plural allative
PP2SA   Pp2-p3  Pronoun personal second plural adessive
PP2SA   Pp2-pb  Pronoun personal second plural ablative
PP2S    Pp2-p4  Pronoun personal second plural translative
PP2S    Pp2-p9  Pronoun personal second plural terminative
PP2S    Pp2-pw  Pronoun personal second plural essive
PP2S    Pp2-p5  Pronoun personal second plural abessive
PP2S    Pp2-pk  Pronoun personal second plural komitative
PP3SN   Pp3-sn  Pronoun personal third singular nominative

PP3SG   Pp3-sg  Pronoun personal third singular genitive
PP3S1   Pp3-s1  Pronoun personal third singular partitive
PP3S    Pp3-sx  Pronoun personal third singular illative
PP3SA   Pp3-s2  Pronoun personal third singular inessive
PP3SA   Pp3-se  Pronoun personal third singular elative
PP3SA   Pp3-st  Pronoun personal third singular allative
PP3SA   Pp3-s3  Pronoun personal third singular adessive
PP3SA   Pp3-sb  Pronoun personal third singular ablative
PP3S    Pp3-s4  Pronoun personal third singular translative
PP3S    Pp3-s9  Pronoun personal third singular terminative
PP3S    Pp3-sw  Pronoun personal third singular essive
PP3S    Pp3-s5  Pronoun personal third singular abessive
PP3S    Pp3-sk  Pronoun personal third singular komitative
PP3SN   Pp3-pn  Pronoun personal third plural nominative
PP3SG   Pp3-pg  Pronoun personal third plural genitive
PP3S1   Pp3-p1  Pronoun personal third plural partitive
PP3S    Pp3-px  Pronoun personal third plural illative
PP3SA   Pp3-p2  Pronoun personal third plural inessive
PP3SA   Pp3-pe  Pronoun personal third plural elative
PP3SA   Pp3-pt  Pronoun personal third plural allative
PP3SA   Pp3-p3  Pronoun personal third plural adessive
PP3SA   Pp3-pb  Pronoun personal third plural ablative
PP3S    Pp3-p4  Pronoun personal third plural translative
PP3S    Pp3-p9  Pronoun personal third plural terminative
PP3S    Pp3-pw  Pronoun personal third plural essive
PP3S    Pp3-p5  Pronoun personal third plural abessive
PP3S    Pp3-pk  Pronoun personal third plural komitative
PSN     Pd--sn  Pronoun demonstrative singular nominative
PSG     Pd--sg  Pronoun demonstrative singular genitive
PS1     Pd--s1  Pronoun demonstrative singular partitive
PSA     Pd--sx  Pronoun demonstrative singular illative
PSA     Pd--s2  Pronoun demonstrative singular inessive
PSA     Pd--se  Pronoun demonstrative singular elative
PSA     Pd--st  Pronoun demonstrative singular allative
PSA     Pd--s3  Pronoun demonstrative singular adessive
PSA     Pd--sb  Pronoun demonstrative singular ablative
PS      Pd--s4  Pronoun demonstrative singular translative
PS      Pd--s9  Pronoun demonstrative singular terminative
PS      Pd--sw  Pronoun demonstrative singular essive
PS      Pd--s5  Pronoun demonstrative singular abessive
PS      Pd--sk  Pronoun demonstrative singular komitative
PSA     Pd--s7  Pronoun demonstrative singular aditive
PSN     Pd--pn  Pronoun demonstrative plural nominative
PSG     Pd--pg  Pronoun demonstrative plural genitive
PS1     Pd--p1  Pronoun demonstrative plural partitive
PSA     Pd--px  Pronoun demonstrative plural illative
PSA     Pd--p2  Pronoun demonstrative plural inessive
PSA     Pd--pe  Pronoun demonstrative plural elative
PSA     Pd--pt  Pronoun demonstrative plural allative
PSA     Pd--p3  Pronoun demonstrative plural adessive
PSA     Pd--pb  Pronoun demonstrative plural ablative
PS      Pd--p4  Pronoun demonstrative plural translative

PS      Pd--p9  Pronoun demonstrative plural terminative
PS      Pd--pw  Pronoun demonstrative plural essive
PS      Pd--p5  Pronoun demonstrative plural abessive
PS      Pd--pk  Pronoun demonstrative plural komitative
PSX     Pi      Pronoun indefinite
PSN     Pi--sn  Pronoun indefinite singular nominative
PSG     Pi--sg  Pronoun indefinite singular genitive
PS1     Pi--s1  Pronoun indefinite singular partitive
PSA     Pi--sx  Pronoun indefinite singular illative
PSA     Pi--s2  Pronoun indefinite singular inessive
PSA     Pi--se  Pronoun indefinite singular elative
PSA     Pi--st  Pronoun indefinite singular allative
PSA     Pi--s3  Pronoun indefinite singular adessive
PSA     Pi--sb  Pronoun indefinite singular ablative
PS      Pi--s4  Pronoun indefinite singular translative
PS      Pi--s9  Pronoun indefinite singular terminative
PS      Pi--sw  Pronoun indefinite singular essive
PS      Pi--s5  Pronoun indefinite singular abessive
PS      Pi--sk  Pronoun indefinite singular komitative
PSA     Pi--s7  Pronoun indefinite singular aditive
PSN     Pi--pn  Pronoun indefinite plural nominative
PSG     Pi--pg  Pronoun indefinite plural genitive
PS1     Pi--p1  Pronoun indefinite plural partitive
PSA     Pi--px  Pronoun indefinite plural illative
PSA     Pi--p2  Pronoun indefinite plural inessive
PSA     Pi--pe  Pronoun indefinite plural elative
PSA     Pi--pt  Pronoun indefinite plural allative
PSA     Pi--p3  Pronoun indefinite plural adessive
PSA     Pi--pb  Pronoun indefinite plural ablative
PS      Pi--p4  Pronoun indefinite plural translative
PS      Pi--p9  Pronoun indefinite plural terminative
PS      Pi--pw  Pronoun indefinite plural essive
PS      Pi--p5  Pronoun indefinite plural abessive
PS      Pi--pk  Pronoun indefinite plural komitative
PSX     Ps      Pronoun possessive
PSN     Ps--sn  Pronoun possessive singular nominative
PSG     Ps--sg  Pronoun possessive singular genitive
PS1     Ps--s1  Pronoun possessive singular partitive
PSA     Ps--sx  Pronoun possessive singular illative
PSA     Ps--s2  Pronoun possessive singular inessive
PSA     Ps--se  Pronoun possessive singular elative
PSA     Ps--st  Pronoun possessive singular allative
PSA     Ps--s3  Pronoun possessive singular adessive
PSA     Ps--sb  Pronoun possessive singular ablative
PS      Ps--s4  Pronoun possessive singular translative
PS      Ps--s9  Pronoun possessive singular terminative
PS      Ps--sw  Pronoun possessive singular essive
PS      Ps--s5  Pronoun possessive singular abessive
PS      Ps--sk  Pronoun possessive singular komitative
PSN     Ps--pn  Pronoun possessive plural nominative
PSG     Ps--pg  Pronoun possessive plural genitive

PS1     Ps--p1  Pronoun possessive plural partitive
PSA     Ps--px  Pronoun possessive plural illative
PSA     Ps--p2  Pronoun possessive plural inessive
PSA     Ps--pe  Pronoun possessive plural elative
PSA     Ps--pt  Pronoun possessive plural allative
PSA     Ps--p3  Pronoun possessive plural adessive
PSA     Ps--pb  Pronoun possessive plural ablative
PS      Ps--p4  Pronoun possessive plural translative
PS      Ps--p9  Pronoun possessive plural terminative
PS      Ps--pw  Pronoun possessive plural essive
PS      Ps--p5  Pronoun possessive plural abessive
PS      Ps--pk  Pronoun possessive plural komitative
PSN     Pq--sn  Pronoun interrogative singular nominative
PSG     Pq--sg  Pronoun interrogative singular genitive
PS1     Pq--s1  Pronoun interrogative singular partitive
PSA     Pq--sx  Pronoun interrogative singular illative
PSA     Pq--s2  Pronoun interrogative singular inessive
PSA     Pq--se  Pronoun interrogative singular elative
PSA     Pq--st  Pronoun interrogative singular allative
PSA     Pq--s3  Pronoun interrogative singular adessive
PSA     Pq--sb  Pronoun interrogative singular ablative
PS      Pq--s4  Pronoun interrogative singular translative
PS      Pq--s9  Pronoun interrogative singular terminative
PS      Pq--sw  Pronoun interrogative singular essive
PS      Pq--s5  Pronoun interrogative singular abessive
PS      Pq--sk  Pronoun interrogative singular komitative
PSA     Pq--s7  Pronoun interrogative singular aditive
PSN     Pq--pn  Pronoun interrogative plural nominative
PSG     Pq--pg  Pronoun interrogative plural genitive
PS1     Pq--p1  Pronoun interrogative plural partitive
PSA     Pq--px  Pronoun interrogative plural illative
PSA     Pq--p2  Pronoun interrogative plural inessive
PSA     Pq--pe  Pronoun interrogative plural elative
PSA     Pq--pt  Pronoun interrogative plural allative
PSA     Pq--p3  Pronoun interrogative plural adessive
PSA     Pq--pb  Pronoun interrogative plural ablative
PS      Pq--p4  Pronoun interrogative plural translative
PS      Pq--p9  Pronoun interrogative plural terminative
PS      Pq--pw  Pronoun interrogative plural essive
PS      Pq--p5  Pronoun interrogative plural abessive
PS      Pq--pk  Pronoun interrogative plural komitative
PSN     Pr--sn  Pronoun relative singular nominative
PSG     Pr--sg  Pronoun relative singular genitive
PS1     Pr--s1  Pronoun relative singular partitive
PSA     Pr--sx  Pronoun relative singular illative
PSA     Pr--s2  Pronoun relative singular inessive
PSA     Pr--se  Pronoun relative singular elative
PSA     Pr--st  Pronoun relative singular allative
PSA     Pr--s3  Pronoun relative singular adessive
PSA     Pr--sb  Pronoun relative singular ablative
PS      Pr--s4  Pronoun relative singular translative

PS      Pr--s9  Pronoun relative singular terminative
PS      Pr--sw  Pronoun relative singular essive
PS      Pr--s5  Pronoun relative singular abessive
PS      Pr--sk  Pronoun relative singular komitative
PSA     Pr--s7  Pronoun relative singular aditive
PSN     Pr--pn  Pronoun relative plural nominative
PSG     Pr--pg  Pronoun relative plural genitive
PS1     Pr--p1  Pronoun relative plural partitive
PSA     Pr--px  Pronoun relative plural illative
PSA     Pr--p2  Pronoun relative plural inessive
PSA     Pr--pe  Pronoun relative plural elative
PSA     Pr--pt  Pronoun relative plural allative
PSA     Pr--p3  Pronoun relative plural adessive
PSA     Pr--pb  Pronoun relative plural ablative
PS      Pr--p4  Pronoun relative plural translative
PS      Pr--p9  Pronoun relative plural terminative
PS      Pr--pw  Pronoun relative plural essive
PS      Pr--p5  Pronoun relative plural abessive
PS      Pr--pk  Pronoun relative plural komitative
PSX     Px      Pronoun reflexive
PSN     Px--sn  Pronoun reflexive singular nominative
PSG     Px--sg  Pronoun reflexive singular genitive
PS1     Px--s1  Pronoun reflexive singular partitive
PSA     Px--sx  Pronoun reflexive singular illative
PSA     Px--s2  Pronoun reflexive singular inessive
PSA     Px--se  Pronoun reflexive singular elative
PSA     Px--st  Pronoun reflexive singular allative
PSA     Px--s3  Pronoun reflexive singular adessive
PSA     Px--sb  Pronoun reflexive singular ablative
PS      Px--s4  Pronoun reflexive singular translative
PS      Px--s9  Pronoun reflexive singular terminative
PS      Px--sw  Pronoun reflexive singular essive
PS      Px--s5  Pronoun reflexive singular abessive
PS      Px--sk  Pronoun reflexive singular komitative
PSN     Px--pn  Pronoun reflexive plural nominative
PSG     Px--pg  Pronoun reflexive plural genitive
PS1     Px--p1  Pronoun reflexive plural partitive
PSA     Px--px  Pronoun reflexive plural illative
PSA     Px--p2  Pronoun reflexive plural inessive
PSA     Px--pe  Pronoun reflexive plural elative
PSA     Px--pt  Pronoun reflexive plural allative
PSA     Px--p3  Pronoun reflexive plural adessive
PSA     Px--pb  Pronoun reflexive plural ablative
PS      Px--p4  Pronoun reflexive plural translative
PS      Px--p9  Pronoun reflexive plural terminative
PS      Px--pw  Pronoun reflexive plural essive
PS      Px--p5  Pronoun reflexive plural abessive
PS      Px--pk  Pronoun reflexive plural komitative
PSG     Py--sg  Pronoun reciprocal singular genitive
PS1     Py--s1  Pronoun reciprocal singular partitive
PSA     Py--sx  Pronoun reciprocal singular illative

PSA     Py--s2  Pronoun reciprocal singular inessive
PSA     Py--se  Pronoun reciprocal singular elative
PSA     Py--st  Pronoun reciprocal singular allative
PSA     Py--s3  Pronoun reciprocal singular adessive
PSA     Py--sb  Pronoun reciprocal singular ablative
PS      Py--s4  Pronoun reciprocal singular translative
PS      Py--s9  Pronoun reciprocal singular terminative
PS      Py--sw  Pronoun reciprocal singular essive
PS      Py--s5  Pronoun reciprocal singular abessive
PS      Py--sk  Pronoun reciprocal singular komitative
PSA     Py--s7  Pronoun reciprocal singular aditive
PSG     Py--pg  Pronoun reciprocal plural genitive
PS1     Py--p1  Pronoun reciprocal plural partitive
PSA     Py--px  Pronoun reciprocal plural illative
PSA     Py--p2  Pronoun reciprocal plural inessive
PSA     Py--pe  Pronoun reciprocal plural elative
PSA     Py--pt  Pronoun reciprocal plural allative
PSA     Py--p3  Pronoun reciprocal plural adessive
PSA     Py--pb  Pronoun reciprocal plural ablative
PS      Py--p4  Pronoun reciprocal plural translative
PS      Py--p9  Pronoun reciprocal plural terminative
PS      Py--pw  Pronoun reciprocal plural essive
PS      Py--p5  Pronoun reciprocal plural abessive
PS      Py--pk  Pronoun reciprocal plural komitative
PSX     Pm      Pronoun determinal
PSN     Pm--sn  Pronoun determinal singular nominative
PSG     Pm--sg  Pronoun determinal singular genitive
PS1     Pm--s1  Pronoun determinal singular partitive
PSA     Pm--sx  Pronoun determinal singular illative
PSA     Pm--s2  Pronoun determinal singular inessive
PSA     Pm--se  Pronoun determinal singular elative
PSA     Pm--st  Pronoun determinal singular allative
PSA     Pm--s3  Pronoun determinal singular adessive
PSA     Pm--sb  Pronoun determinal singular ablative
PS      Pm--s4  Pronoun determinal singular translative
PS      Pm--s9  Pronoun determinal singular terminative
PS      Pm--sw  Pronoun determinal singular essive
PS      Pm--s5  Pronoun determinal singular abessive
PS      Pm--sk  Pronoun determinal singular komitative
PSA     Pm--s7  Pronoun determinal singular aditive
PSN     Pm--pn  Pronoun determinal plural nominative
PSG     Pm--pg  Pronoun determinal plural genitive
PS1     Pm--p1  Pronoun determinal plural partitive
PSA     Pm--px  Pronoun determinal plural illative
PSA     Pm--p2  Pronoun determinal plural inessive
PSA     Pm--pe  Pronoun determinal plural elative
PSA     Pm--pt  Pronoun determinal plural allative
PSA     Pm--p3  Pronoun determinal plural adessive
PSA     Pm--pb  Pronoun determinal plural ablative
PS      Pm--p4  Pronoun determinal plural translative
PS      Pm--p9  Pronoun determinal plural terminative
PS      Pm--pw  Pronoun determinal plural essive

PS      Pm--p5  Pronoun determinal plural abessive
PS      Pm--pk  Pronoun determinal plural komitative

4.3. Examples
Corpus tags      Lexicon       Examples
---------------------------------------------
PP1SN   Pp1-sn  mina
PP1SG   Pp1-sg  minu
PP1S1   Pp1-s1  mind
PP1S    Pp1-sx  minusse
PP1SA   Pp1-s2  minus
PP1SA   Pp1-se  minust
PP1SA   Pp1-st  minule
PP1SA   Pp1-s3  minul
PP1SA   Pp1-sb  minult
PP1S    Pp1-s4  minuks
PP1S    Pp1-s9  minuni
PP1S    Pp1-sw  minuna
PP1S    Pp1-s5  minuta
PP1S    Pp1-sk  minuga
PP1SN   Pp1-pn  meie
PP1SG   Pp1-pg  meie
PP1S1   Pp1-p1  meid
PP1S    Pp1-px  meisse
PP1SA   Pp1-p2  meis
PP1SA   Pp1-pe  meist
PP1SA   Pp1-pt  meile
PP1SA   Pp1-p3  meil
PP1SA   Pp1-pb  meilt
PP1S    Pp1-p4  meieks
PP1S    Pp1-p9  meieni
PP1S    Pp1-pw  meiena
PP1S    Pp1-p5  meieta
PP1S    Pp1-pk  meiega
PP2SN   Pp2-sn  sina
PP2SG   Pp2-sg  sinu
PP2S1   Pp2-s1  sind
PP2S    Pp2-sx  sinusse
PP2SA   Pp2-s2  sinus
PP2SA   Pp2-se  sinust
PP2SA   Pp2-st  sinule
PP2SA   Pp2-s3  sinul
PP2SA   Pp2-sb  sinult
PP2S    Pp2-s4  sinuks
PP2S    Pp2-s9  sinuni
PP2S    Pp2-sw  sinuna
PP2S    Pp2-s5  sinuta
PP2S    Pp2-sk  sinuga
PP2SN   Pp2-pn  teie
PP2SG   Pp2-pg  teie

PP2S1   Pp2-p1  teid
PP2S    Pp2-px  teisse
PP2SA   Pp2-p2  teis
PP2SA   Pp2-pe  teist
PP2SA   Pp2-pt  teile
PP2SA   Pp2-p3  teil
PP2SA   Pp2-pb  teilt
PP2S    Pp2-p4  teieks
PP2S    Pp2-p9  teieni
PP2S    Pp2-pw  teiena
PP2S    Pp2-p5  teieta
PP2S    Pp2-pk  teiega
PP3SN   Pp3-sn  tema
PP3SG   Pp3-sg  tema
PP3S1   Pp3-s1  teda
PP3S    Pp3-sx  temasse
PP3SA   Pp3-s2  temas
PP3SA   Pp3-se  temast
PP3SA   Pp3-st  temale
PP3SA   Pp3-s3   temal
PP3SA   Pp3-sb  temalt
PP3S    Pp3-s4  temaks
PP3S    Pp3-s9  temani
PP3S    Pp3-sw  temana
PP3S    Pp3-s5  temata
PP3S    Pp3-sk  temaga
PP3SN   Pp3-pn  nemad
PP3SG   Pp3-pg  nende
PP3S1   Pp3-p1  neid
PP3S    Pp3-px  nendesse, neisse
PP3SA   Pp3-p2  nendes, neis
PP3SA   Pp3-pe  nendest, neist
PP3SA   Pp3-pt  nendele, neile
PP3SA   Pp3-p3  nendel, neil
PP3SA   Pp3-pb  nendelt, neilt
PP3S    Pp3-p4  nendeks, neiks
PP3S    Pp3-p9  nendeni
PP3S    Pp3-pw  nendena
PP3S    Pp3-p5  nendeta
PP3S    Pp3-pk  nendega
PSN     Pd--sn  selline
PSG     Pd--sg  sellise
PS1     Pd--s1  sellist
PSA     Pd--sx  sellisesse
PSA     Pd--s2  sellises
PSA     Pd--se  sellisest
PSA     Pd--st  sellisele
PSA     Pd--s3  sellisel
PSA     Pd--sb  selliselt
PS      Pd--s4  selliseks
PS      Pd--s9  selliseni
PS      Pd--sw  sellisena

PS      Pd--s5  selliseta
PS      Pd--sk  sellisega
PSA     Pd--s7  sellisse
PSN     Pd--pn  sellised
PSG     Pd--pg  selliste
PS1     Pd--p1  selliseid
PSA     Pd--px  sellistesse, selliseisse
PSA     Pd--p2  sellistes, selliseis
PSA     Pd--pe  sellistest, selliseist
PSA     Pd--pt  sellistele, selliseile
PSA     Pd--p3  sellistel, selliseil
PSA     Pd--pb  sellistelt, selliseilt
PS      Pd--p4  sellisteks, selliseiks
PS      Pd--p9  sellisteni, selliseini
PS      Pd--pw  sellistena, selliseina
PS      Pd--p5  sellisteta, selliseita
PS      Pd--pk  sellistega selliseiga
PSX     Pi      muist
PSN     Pi--sn  mingi
PSG     Pi--sg  mingi
PS1     Pi--s1  mingit
PSA     Pi--sx  mingisse
PSA     Pi--s2  mingis
PSA     Pi--se  mingist
PSA     Pi--st  mingile
PSA     Pi--s3  mingil
PSA     Pi--sb  mingilt
PS      Pi--s4  mingiks
PS      Pi--s9  mingini
PS      Pi--sw  mingina
PS      Pi--s5  mingita
PS      Pi--sk  mingiga
PSA     Pi--s7  mitmetaolisse
PSN     Pi--pn  mingid
PSG     Pi--pg  mingite
PS1     Pi--p1  mingeid
PSA     Pi--px  mingitesse, mingeisse
PSA     Pi--p2  mingites, mingeis
PSA     Pi--pe  mingitest, mingeist
PSA     Pi--pt  mingitele, mingeile
PSA     Pi--p3  mingitel, mingeil
PSA     Pi--pb  mingitelt, mingeilt
PS      Pi--p4  mingiteks, mingeiks
PS      Pi--p9  mingiteni, mingeini
PS      Pi--pw  mingitena, mingeina
PS      Pi--p5  mingiteta, mingeita
PS      Pi--pk  mingitega, mingeiga
PSX     Ps      oma
PSN     Ps--sn  iseenese
PSG     Ps--sg  iseenese
PS1     Ps--s1  iseennast

PSA     Ps--sx  iseenesesse
PSA     Ps--s2  iseeneses
PSA     Ps--se  iseenesest
PSA     Ps--st  iseenesele
PSA     Ps--s3  iseenesel
PSA     Ps--sb  iseeneselt
PS      Ps--s4  iseeneseks
PS      Ps--s9  iseeneseni
PS      Ps--sw  iseenesena
PS      Ps--s5  iseeneseta
PS      Ps--sk  iseenesega
PSN     Ps--pn  iseenese?
PSG     Ps--pg  iseeneste
PS1     Ps--p1  iseendid
PSA     Ps--px  iseenestesse
PSA     Ps--p2  iseenestes
PSA     Ps--pe  iseenestest
PSA     Ps--pt  iseenestele
PSA     Ps--p3  iseenestel
PSA     Ps--pb  iseenestelt
PS      Ps--p4  iseenesteks
PS      Ps--p9  iseenesteni
PS      Ps--pw  iseenestena
PS      Ps--p5  iseenesteta
PS      Ps--pk  iseenestega
PSN     Pq--sn  milline
PSG     Pq--sg  millise
PS1     Pq--s1  millist
PSA     Pq--sx  millisesse
PSA     Pq--s2  millises
PSA     Pq--se  millisest
PSA     Pq--st  millisele
PSA     Pq--s3  millisel
PSA     Pq--sb  milliselt
PS      Pq--s4  milliseks
PS      Pq--s9  milliseni
PS      Pq--sw  millisena
PS      Pq--s5  milliseta
PS      Pq--sk  millisega
PSA     Pq--s7  millisse?
PSN     Pq--pn  millised
PSG     Pq--pg  milliste
PS1     Pq--p1  milliseid
PSA     Pq--px  millistesse, milliseisse
PSA     Pq--p2  millistes, milliseis
PSA     Pq--pe  millistest, milliseist
PSA     Pq--pt  millistele, milliseile
PSA     Pq--p3  millistel, milliseil
PSA     Pq--pb  millistelt, milliseilt
PS      Pq--p4  millisteks, milliseiks
PS      Pq--p9  millisteni, milliseini
PS      Pq--pw  millistena, milliseina

PS      Pq--p5  millisteta, milliseita
PS      Pq--pk  millistega, milliseiga
PSN     Pr--sn  milline
PSG     Pr--sg  millise
PS1     Pr--s1  millist
PSA     Pr--sx  millisesse
PSA     Pr--s2  millises
PSA     Pr--se  millisest
PSA     Pr--st  millisele
PSA     Pr--s3  millisel
PSA     Pr--sb  milliselt
PS      Pr--s4  milliseks
PS      Pr--s9  milliseni
PS      Pr--sw  millisena
PS      Pr--s5  milliseta
PS      Pr--sk  millisega
PSA     Pr--s7  millisse?
PSN     Pr--pn  millised
PSG     Pr--pg  milliste
PS1     Pr--p1  milliseid
PSA     Pr--px  millistesse, milliseisse
PSA     Pr--p2  millistes, milliseis
PSA     Pr--pe  millistest, milliseist
PSA     Pr--pt  millistele, milliseile
PSA     Pr--p3  millistel, milliseil
PSA     Pr--pb  millistelt, milliseilt
PS      Pr--p4  millisteks, milliseiks
PS      Pr--p9  millisteni, milliseini
PS      Pr--pw  millistena, milliseina
PS      Pr--p5  millisteta, milliseita
PS      Pr--pk  millistega, milliseiga
PSX     Px      oma
PSN     Px--sn  ise
PSG     Px--sg  iseenda
PS1     Px--s1  iseend
PSA     Px--sx  iseendasse
PSA     Px--s2  iseendas
PSA     Px--se  iseendast
PSA     Px--st  iseendale
PSA     Px--s3  iseendal
PSA     Px--sb  iseendalt
PS      Px--s4  iseendaks
PS      Px--s9  iseendani
PS      Px--sw  iseendana
PS      Px--s5  iseendata
PS      Px--sk  iseendaga
PSN     Px--pn  ise
PSG     Px--pg  iseendi
PS1     Px--p1  iseendid
PSA     Px--px  iseendisse
PSA     Px--p2  iseendis

PSA     Px--pe  iseendist
PSA     Px--pt  iseendile
PSA     Px--p3  iseendil
PSA     Px--pb  iseendilt
PS      Px--p4  iseendiks
PS      Px--p9  iseendini
PS      Px--pw  iseendina
PS      Px--p5  iseendita
PS      Px--pk  iseendiga
PSG     Py--sg  teineteise
PS1     Py--s1  teineteist
PSA     Py--sx  teineteisesse
PSA     Py--s2  teineteises
PSA     Py--se  teineteisest
PSA     Py--st  teineteisele
PSA     Py--s3  teineteisel
PSA     Py--sb  teineteiselt
PS      Py--s4  teineteiseks
PS      Py--s9  teineteiseni
PS      Py--sw  teineteisena
PS      Py--s5  teineteiseta
PS      Py--sk  teineteisega
PSA     Py--s7  teineteise
PSG     Py--pg  teineteiste
PS1     Py--p1  teineteisi
PSA     Py--px  teineteistesse
PSA     Py--p2  teineteistes
PSA     Py--pe  teineteistest
PSA     Py--pt  teineteistele
PSA     Py--p3  teineteistel
PSA     Py--pb  teineteistelt
PS      Py--p4  teineteisteks
PS      Py--p9  teineteisteni
PS      Py--pw  teineteistena
PS      Py--p5  teineteisteta
PS      Py--pk  teineteistega
PSX     Pm      ise
PSN     Pm--sn  ise
PSG     Pm--sg  enda
PS1     Pm--s1  end
PSA     Pm--sx  endasse
PSA     Pm--s2  endas
PSA     Pm--se  endast
PSA     Pm--st  endale
PSA     Pm--s3  endal
PSA     Pm--sb  endalt
PS      Pm--s4  endaks
PS      Pm--s9  endani
PS      Pm--sw  endana
PS      Pm--s5  endata
PS      Pm--sk  endaga

PSA     Pm--s7  emba-kumba
PSNPm--pn       ise
PSG     Pm--pg  endi
PS1     Pm--p1  endid
PSA     Pm--px  endisse
PSA     Pm--p2  endis
PSA     Pm--pe  endist
PSA     Pm--pt  endile
PSA     Pm--p3  endil
PSA     Pm--pb  endilt
PS      Pm--p4  endiks
PS      Pm--p9  endini
PS      Pm--pw  endina
PS      Pm--p5  endita
PS      Pm--pk  endiga

9.5 Determiners (D)

Not applicable.

9.6 Atricles (T)

Not applicable.

9.7 Adverbs (R)

7.1 Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type                          - asjatult
- -------------- -------------- - ------------
2 Degree                        -
**********************************************
3 Clitic                        -
- -------------- -------------- - ------------
4 Number                        -
- -------------- -------------- - ------------

5 Person                        -
================================= ============

7.2. Definitions
Corpus tags       Lexicon       Definitions
-----------------------------------------------
RR      R       Adverb

7.3. Examples
Corpus tags      Lexicon       Examples
----------------------------------------------
RR      R       asjatult

9.8 Adpositions (S)

8.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           preposition    p enne
                 postposition   t kaudu
- -------------- -------------- - ------------
2 Formation                     -
**********************************************
3 Case                          -
- -------------- -------------- - ------------
4 Clitic                        -
================================= ============

8.2. Definitions
Corpus tags       Lexicon       Definitions
----------------------------------------------
SP      Sp      Adposition preposition
ST      St      Adposition postposition

8.3. Examples
Corpus tags      Lexicon       Examples
----------------------------------------------

SP      Sp      enne
ST      St      kaudu

9.9 Conjunctions (C)

9.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           coordinating   c ja
                 subordinating  s et
**********************************************
2 Formation                     -
- -------------- -------------- - ------------
3 Coord_Type                    -
- -------------- -------------- - ------------
4 Sub_Type                      -
- -------------- -------------- - ------------
5 Clitic                        -
- -------------- -------------- - ------------
6 Number                        -
- -------------- -------------- - ------------
7 Person                        -
================================= ============

9.2. Definitions
Corpus tags       Lexicon       Definitions
----------------------------------------------
CC      Cc      Conjuction coordinating
CS      Cs      Conjuction subordinating

9.3. Examples
Corpus tags      Lexicon       Examples
--------------------------------------------
CC      Cc      ja
CS      Cs      et

9.10 Numerals (M)

10.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           cardinal       c kaheksa
                 ordinal        o kaheksas
- -------------- -------------- - ------------
2 Gender                        -
- -------------- -------------- - ------------
3 Number         singular       s kaheksa
                 plural         p kaheksad
- -------------- -------------- - ------------
4 Case           nominative     n kaheksa
                 genitive       g kaheksa
           l.s.  partitive      1 kaheksat
                 illative       x kaheksasse
                 inessive       2 kaheksas
                 elative        e kaheksast
                 allative       t kaheksale
                 adessive       3 kaheksal
                 ablative       b kaheksalt
           l.s.  translative    4 kaheksaks
                 terminative    9 kaheksani
                 essive         w kaheksana
           l.s.  abessive       5 kaheksata
           l.s.  komitative     k kaheksaga
           l.s.  aditive        7 viite
********************************* ------------
5 Form           digit          d 1984
                 roman          r IX
                 letter         l kaheksa
- -------------- -------------- - ------------
6 Definiteness                  -
- -------------- -------------- - ------------
7 Clitic                        -
- -------------- -------------- - ------------
8 Class                         -
- -------------- -------------- - ------------
9 Animate                       -
- -------------- -------------- - ------------
10Owner_Number                  -
- -------------- -------------- - ------------
11Owner_Person                  -
- -------------- -------------- - ------------
12Owned_Number                  -
================================= ============

10.2. Definitions
Corpus tags       Lexicon       Definitions
------------------------------------------------

MCSX    Mc---d  Numeral cardinal digit
MCSN    Mc-snl  Numeral cardinal singular nominative letter
MCSG    Mc-sgl  Numeral cardinal singular genitive letter
MCS1    Mc-s1l  Numeral cardinal singular partitive letter
MCSA    Mc-sxl  Numeral cardinal singular illative letter
MCSA    Mc-s2l  Numeral cardinal singular inessive letter
MCSA    Mc-sel  Numeral cardinal singular elative letter
MCSA    Mc-stl  Numeral cardinal singular allative letter
MCSA    Mc-s3l  Numeral cardinal singular adessive letter
MCSA    Mc-sbl  Numeral cardinal singular ablative letter
MCS     Mc-s4l  Numeral cardinal singular translative letter
MCS     Mc-s9l  Numeral cardinal singular terminative letter
MCS     Mc-swl  Numeral cardinal singular essive letter
MCS     Mc-s5l  Numeral cardinal singular abessive letter
MCS     Mc-skl  Numeral cardinal singular komitative letter
MCSA    Mc-s7l  Numeral cardinal singular aditive letter
MCS     Mc-pnl  Numeral cardinal plural nominative letter
MCS     Mc-pgl  Numeral cardinal plural genitive letter
MCS     Mc-p1l  Numeral cardinal plural partitive letter
MCSA    Mc-pxl  Numeral cardinal plural illative letter
MCSA    Mc-p2l  Numeral cardinal plural inessive letter
MCSA    Mc-pel  Numeral cardinal plural elative letter
MCSA    Mc-ptl  Numeral cardinal plural allative letter
MCSA    Mc-p3l  Numeral cardinal plural adessive letter
MCSA    Mc-pbl  Numeral cardinal plural ablative letter
MCS     Mc-p4l  Numeral cardinal plural translative letter
MCS     Mc-p9l  Numeral cardinal plural terminative letter
MCS     Mc-pwl  Numeral cardinal plural essive letter
MCS     Mc-p5l  Numeral cardinal plural abessive letter
MCS     Mc-pkl  Numeral cardinal plural komitative letter
MOSX    Mo---d  Numeral ordinal digit
MOSX    Mo---r  Numeral ordinal roman
MOSN    Mo-snl  Numeral ordinal singular nominative letter
MOSG    Mo-sgl  Numeral ordinal singular genitive letter
MOS1    Mo-s1l  Numeral ordinal singular partitive letter
MOSA    Mo-sxl  Numeral ordinal singular illative letter
MOSA    Mo-s2l  Numeral ordinal singular inessive letter
MOSA    Mo-sel  Numeral ordinal singular elative letter
MOSA    Mo-stl  Numeral ordinal singular allative letter
MOSA    Mo-s3l  Numeral ordinal singular adessive letter
MOSA    Mo-sbl  Numeral ordinal singular ablative letter
MOS     Mo-s4l  Numeral ordinal singular translative letter
MOS     Mo-s9l  Numeral ordinal singular terminative letter
MOS     Mo-swl  Numeral ordinal singular essive letter
MOS     Mo-s5l  Numeral ordinal singular abessive letter
MOS     Mo-skl  Numeral ordinal singular komitative letter
MOSA    Mo-s7l  Numeral ordinal singular aditive letter
MOS     Mo-pnl  Numeral ordinal plural nominative letter
MOS     Mo-pgl  Numeral ordinal plural genitive letter
MOS     Mo-p1l  Numeral ordinal plural partitive letter
MOSA    Mo-pxl  Numeral ordinal plural illative letter
MOSA    Mo-p2l  Numeral ordinal plural inessive letter

MOSA    Mo-pel  Numeral ordinal plural elative letter
MOSA    Mo-ptl  Numeral ordinal plural allative letter
MOSA    Mo-p3l  Numeral ordinal plural adessive letter
MOSA    Mo-pbl  Numeral ordinal plural ablative letter
MOS     Mo-p4l  Numeral ordinal plural translative letter
MOS     Mo-p9l  Numeral ordinal plural terminative letter
MOS     Mo-pwl  Numeral ordinal plural essive letter
MOS     Mo-p5l  Numeral ordinal plural abessive letter
MOS     Mo-pkl  Numeral ordinal plural komitative letter

10.3. Examples
Corpus tags      Lexicon       Examples
-----------------------------------------
MCSX    Mc---d  8
MCSN    Mc-snl  kaheksa
MCSG    Mc-sgl  kaheksa
MCS1    Mc-s1l  kaheksat
MCSA    Mc-sxl  kaheksasse
MCSA    Mc-s2l  kaheksas
MCSA    Mc-sel  kaheksast
MCSA    Mc-stl  kaheksale
MCSA    Mc-s3l  kaheksal
MCSA    Mc-sbl  kaheksalt
MCS     Mc-s4l  kaheksaks
MCS     Mc-s9l  kaheksani
MCS     Mc-swl  kaheksana
MCS     Mc-s5l  kaheksata
MCS     Mc-skl  kaheksaga
MCSA    Mc-s7l  viide
MCS     Mc-pnl  kaheksad
MCS     Mc-pgl  kaheksate
MCS     Mc-p1l  kaheksaid
MCSA    Mc-pxl  kaheksatesse, kaheksaisse
MCSA    Mc-p2l  kaheksates, kaheksais
MCSA    Mc-pel  kaheksatest, kaheksaist
MCSA    Mc-ptl  kaheksatele, kaheksaile
MCSA    Mc-p3l  kaheksatel, kaheksail
MCSA    Mc-pbl  kaheksatelt, kaheksailt
MCS     Mc-p4l  kaheksateks, kaheksaiks
MCS     Mc-p9l  kaheksateni, kaheksaini
MCS     Mc-pwl  kaheksatena, kaheksaina
MCS     Mc-p5l  kaheksateta
MCS     Mc-pkl  kaheksatega
MOSX    Mo---d  8.
MOSX    Mo---r  IX
MOSN    Mo-snl  kaheksas
MOSG    Mo-sgl  kaheksanda
MOS1    Mo-s1l  kaheksandat
MOSA    Mo-sxl  kaheksandasse
MOSA    Mo-s2l  kaheksandas

MOSA    Mo-sel  kaheksandast
MOSA    Mo-stl  kaheksandale
MOSA    Mo-s3l  kaheksandal
MOSA    Mo-sbl  kaheksandalt
MOS     Mo-s4l  kaheksandaks
MOS     Mo-s9l  kaheksandani
MOS     Mo-swl  kaheksandana
MOS     Mo-s5l  kaheksandata
MOS     Mo-skl  kaheksandaga
MOSA    Mo-s7l  teise
MOS     Mo-pnl  kaheksandad
MOS     Mo-pgl  kaheksandate
MOS     Mo-p1l  kaheksandaid
MOSA    Mo-pxl  kaheksandatesse, kaheksandaisse
MOSA    Mo-p2l  kaheksandates, kaheksandais
MOSA    Mo-pel  kaheksandatest, kaheksandaist
MOSA    Mo-ptl  kaheksandatele, kaheksandaile
MOSA    Mo-p3l  kaheksandatel, kaheksandail
MOSA    Mo-pbl  kaheksandatelt, kaheksandailt
MOS     Mo-p4l  kaheksandateks, kaheksandaiks
MOS     Mo-p9l  kaheksandateni, kaheksandaini
MOS     Mo-pwl  kaheksandatena, kaheksandaina
MOS     Mo-p5l  kaheksandateta
MOS     Mo-pkl  kaheksandatega

9.11 Interjections (I)

11.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type                          - oi
- -------------- -------------- - ------------
2 Formation                     -
================================= ============

11.2. Definitions
Corpus tags       Lexicon       Definitions
---------------------------------------
II      I       Interjection

11.3. Examples
Corpus tags      Lexicon       Examples

---------------------------------------
II      I       oi

9.12 Residual (X)

12.1. Lexicon
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
                                -
=================================

13.2. Definitions
Corpus tags       Lexicon       Definitions
------------------------------------------------
X       X       foreign word

13.2. Examples
Corpus tags      Lexicon       Examples
---------------------------------------
X       X       qwerty

9.13 Abreviations (Y)

13.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C
= ============== ============== = ============
1 Syntactic_Type nominal        n CD
                 verbal         v vt
                 adjectival     a hisp
                 adverbial      r jne
- ------------------------------- ------------
2 Gender                        -
- -------------- -------------- - ------------
3 Number         singular       s CD
                 plural         p CDd

- -------------- -------------- - ------------
4 Case     l.s.  nominative     n CD
           l.s.  genitive       g CD
           l.s.  partitive      1 CDd
           l.s.  illative       x CDsse
           l.s.  inessive       2 CDs
           l.s.  elative        e CDst
           l.s.  allative       t CDle
           l.s.  adessive       3 CDl
           l.s.  ablative       b CDlt
           l.s.  translative    4 CDks
           l.s.  terminative    9 CDni
           l.s.  essive         w CDna
           l.s.  abessive       5 CDta
           l.s.  komitative     k CDga
           l.s.  aditive        7
- -------------- -------------- - ------------
5 Definiteness                  -
================================= ============

13.2. Definitions
Corpus tags       Lexicon       Definitions
------------------------------------------------
YSX     Yn      Abbreviation nominal
YSN     Yn-sn   Abbreviation nominal singular nominative
YSG     Yn-sg   Abbreviation nominal singular genitive
YS1     Yn-s1   Abbreviation nominal singular partitive
YSA     Yn-sx   Abbreviation nominal singular illative
YSA     Yn-s2   Abbreviation nominal singular inessive
YSA     Yn-se   Abbreviation nominal singular elative
YSA     Yn-st   Abbreviation nominal singular allative
YSA     Yn-s3   Abbreviation nominal singular adessive
YSA     Yn-sb   Abbreviation nominal singular ablative
YS      Yn-s4   Abbreviation nominal singular translative
YS      Yn-s9   Abbreviation nominal singular terminative
YS      Yn-sw   Abbreviation nominal singular essive
YS      Yn-s5   Abbreviation nominal singular abessive
YS      Yn-sk   Abbreviation nominal singular komitative
YS      Yn-s7   Abbreviation nominal singular aditive
YS      Yn-pn   Abbreviation nominal plural nominative
YS      Yn-pg   Abbreviation nominal plural genitive
YS      Yn-p1   Abbreviation nominal plural partitive
YSA     Yn-px   Abbreviation nominal plural illative
YSA     Yn-p2   Abbreviation nominal plural inessive
YSA     Yn-pe   Abbreviation nominal plural elative
YSA     Yn-pt   Abbreviation nominal plural allative
YSA     Yn-p3   Abbreviation nominal plural adessive
YSA     Yn-pb   Abbreviation nominal plural ablative
YS      Yn-p4   Abbreviation nominal plural translative
YS      Yn-p9   Abbreviation nominal plural terminative
YS      Yn-pw   Abbreviation nominal plural essive

YS      Yn-p5   Abbreviation nominal plural abessive
YS      Yn-pk   Abbreviation nominal plural komitative
YSX     Ya      Abbreviation adjectival
YSX     Yr      Abbreviation adverbial
YSX     Yv      Abbreviation verbal

13.3. Examples
Corpus tags      Lexicon       Examples
----------------------------------
YSX     Yn      A
YSN     Yn-sn   CD
YSG     Yn-sg   CD
YS1     Yn-s1   CDd
YSA     Yn-sx   CDsse
YSA     Yn-s2   CDs
YSA     Yn-se   CDst
YSA     Yn-st   CDle
YSA     Yn-s3   CDl
YSA     Yn-sb   CDlt
YS      Yn-s4   CDks
YS      Yn-s9   CDni
YS      Yn-sw   CDna
YS      Yn-s5   CDta
YS      Yn-sk   CDga
YS      Yn-s7   RASi
YS      Yn-pn   CDd
YS      Yn-pg   CDde
YS      Yn-p1   CDsid
YSA     Yn-px   CDdesse
YSA     Yn-p2   CDdes
YSA     Yn-pe   CDdest
YSA     Yn-pt   CDdele
YSA     Yn-p3   CDdel
YSA     Yn-pb   CDdelt
YS      Yn-p4   CDdeks
YS      Yn-p9   CDdeni
YS      Yn-pw   CDdena
YS      Yn-p5   CDdeta
YS      Yn-pk   CDdega
YSX     Ya      hisp
YSX     Yr      jne
YSX     Yv      vt

Multext-East / Concede  Morphosyntactic specifications V.2— Estonian  181

9.14 Particles (Q)

Not applicable.

Application to Hungarian

Authors: László Tihanyi, Tamás Váradi, Csaba Oravecz.

10.1 Noun (N)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           common         c
proper         p
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number         singular       s
plural         p
- -------------- -------------- -
4 Case           nominative     n
genitive       g
dative         d
accusative     a
instrumental   i
illative       x
inessive       2
elative        e
allative       t
adessive       3
ablative       b
terminative    9
essive         w
l.s.  temporalis     m
l.s.  causalis       c
l.s.  sublative      s
l.s.  delative       h
l.s.  sociative      q
l.s.  factive        y

Multext-East / Concede  Morphosyntactic specifications V.2— Hungarian  183

           l.s.  superessive    p
           l.s.  distributive   u
* ***************************** *
5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -
- -------------- -------------- -
7 Animate        no             n
                 yes            y
- -------------- -------------- -
8 Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
9 Owner_Person   first          1
                 second         2
                 third          3
---------------- -------------- -
10Owned_Number   singular       s
                 plural         p
=================================

Notes:
1. Gender: Not applicable to Hungarian.
2. Numbers:
   Hungarian has three types of number in the nominal inflection:
   1. The number of the noun.
   2. The number of owners that own the noun.
   3. The number of the context given referent, which is some possession
      of the noun, i.e. belongs to the noun (anaphoric possessive).
3. Genitive is rarely marked in Hungarian. If marked then with the same
   suffix as that of dative case.
   Nouns with zero suffix can be nominative or genitive, so they are
   ambigious.
4. Common Hungarian nouns can have
   2(Number)*7(OwnerPerson,OwnerNumber)*3(OwnedNumber)*22(Case)=1324
   different forms.
5. The usage of -szerte[SZERTE] meaning 'many places around' is restricted
   so we do not take it as a regular case ending. The same stands for
   '-lag/-leg' marked with [LAG], e.g. 'házilag' means (made) at home.

Combinations:
0   1    3    4                        7     8        9         10
--- ---- ---- ------------------------ ----  -------- --------- --------
Pos Type Num  Case                     Anim  OwnerNum OwnerPers OwnedNum
--- ---- ---- ------------------------ ----  -------- --------- --------

 N  [cp] [sp] [ngdaix2et3b9wmcshqypu] [yn]  [sp]     [123]     [sp]

Examples on combinations:
a) Combinations on Number, OwnerNumber, OwnerPerson and OwnedNumber:
01234567890
Nc-sn--n---  asztal[N]                          table
Nc-sn--n--s  asztal[N]+é[POS]                   thing of the table
Nc-sn--n--p  asztal[N]+éi[POSi]                 things of the table
Nc-sn--ns1-  asztal[N]+om[PSe1]                 my table
Nc-sn--ns1s  asztal[N]+om[PSe1]+é[POS]          thing of my table
Nc-sn--ns1p  asztal[N]+om[PSe1]+éi[POSi]        things of my table
Nc-sn--ns2-  asztal[N]+od[PSe2]                 your table
Nc-sn--ns2s  asztal[N]+od[PSe2]+é[POS]          thing of your table
Nc-sn--ns2p  asztal[N]+od[PSe2]+éi[POSi]        things of your table
Nc-sn--ns3-  asztal[N]+a[PSe3]                  his/her table
Nc-sn--ns3s  asztal[N]+á[PSe3]+é[POS]           thing of his/her table
Nc-sn--ns3p  asztal[N]+á[PSe3]+éi[POSi]         things of his/her table
Nc-sn--np1-  asztal[N]+unk[PSt1]                our table
Nc-sn--np1s  asztal[N]+unk[PSt1]+é[POS]         thing of our table
Nc-sn--np1p  asztal[N]+unk[PSt1]+éi[POSi]       things of our table
Nc-sn--np2-  asztal[N]+otok[PSt2]               your table
Nc-sn--np2s  asztal[N]+otok[PSt2]+é[POS]        thing of your table
Nc-sn--np2p  asztal[N]+otok[PSt2]+éi[POSi]      things of your table
Nc-sn--np3-  asztal[N]+uk[PSt2]                 their table
Nc-sn--np3s  asztal[N]+uk[PSt2]+é[POS]          thing of their table
Nc-sn--np3p  asztal[N]+uk[PSt2]+éi[POSi]        things of their table
Nc-pn--n---  asztal[N]+ok[PL]                   tables
Nc-pn--n--s  asztal[N]+ok[PL]+é[POS]            thing of the tables
Nc-pn--n--p  asztal[N]+ok[PL]+éi[POSi]          things of the tables
Nc-pn--ns1-  asztal[N]+aim[PSe1i]               my tables
Nc-pn--ns1s  asztal[N]+aim[PSe1i]+é[POS]        thing of my tables
Nc-pn--ns1p  asztal[N]+aim[PSe1i]+éi[POSi]      things of my tables
Nc-pn--ns2-  asztal[N]+aid[PSe2i]               your tables
Nc-pn--ns2s  asztal[N]+aid[PSe2i]+é[POS]        thing of your tables
Nc-pn--ns2p  asztal[N]+aid[PSe2i]+éi[POSi]      things of your tables
Nc-pn--ns3-  asztal[N]+ai[PSe3i]                his/her tables
Nc-pn--ns3s  asztal[N]+ai[PSe3i]+é[POS]         thing of his/her tables
Nc-pn--ns3p  asztal[N]+ai[PSe3i]+éi[POSi]       things of his/her tables
Nc-pn--np1-  asztal[N]+aink[PSt1i]              our tables
Nc-pn--np1s  asztal[N]+aink[PSt1i]+é[POS]       thing of our tables
Nc-pn--np1p  asztal[N]+aink[PSt1i]+éi[POSi]     things of our tables
Nc-pn--np2-  asztal[N]+aitok[PSt2i]             your tables
Nc-pn--np2s  asztal[N]+aitok[PSt2i]+é[POS]      thing of your tables
Nc-pn--np2p  asztal[N]+aitok[PSt2i]+éi[POSi]    things of your tables
Nc-pn--np3-  asztal[N]+aik[PSt3i]               their tables
Nc-pn--np3s  asztal[N]+aik[PSt3i]+é[POS]        thing of their tables
Nc-pn--np3p  asztal[N]+aik[PSt3i]+éi[POSi]      things of their tables

b) Combinations on Case:

Category:   Case:          Example analysis:
01234567890
Nc-sn------ nominative     asztal[N]              The table is new.
Nc-sg------ genitive       asztal[N]+nak[GEN]     The leg of the table is red.
Nc-sd------ dative         asztal[N]+nak[DAT]     It gives a new look to the table.
Nc-sa------ accusative     asztal[N]+t[ACC]       I see the table.
Nc-si------ instrumental   asztal[N]+lal[INS]     Do something with the table.
Nc-sx------ illative       asztal[N]+ba[ILL]      She put it into the table.
Nc-s2------ inessive       asztal[N]+ban[INE]     He keeps it in this table.
Nc-se------ elative        asztal[N]+ból[ELA]     She took it out from this table.
Nc-st------ allative       asztal[N]+hoz[ALL]     Go to the table.
Nc-s3------ adessive       asztal[N]+nál[ADE]     At the table there is a chair.
Nc-sb------ ablative       asztal[N]+tól[ABL]     She came from the table.
Nc-s9------ terminative    asztal[N]+ig[TER]      Go as far as the table.
Nc-sw------ essive         asztal[N]+ul[ESS]      The plate served us as a table.
Nc-sm------ temporalis     hat[N]+kor[TEM]        We arrived at six.
Nc-sc------ causalis       asztal[N]+ért[CAU]     We do it for the table.
Nc-ss------ sublative      asztal[N]+ra[SUB]      Put it onto the table.
Nc-sh------ delative       asztal[N]+ról[DEL]     Talk about the table.
Nc-sq------ sociative      asztal[N]+ostul[SOC]   It is sold together with the table.
Nc-sy------ factive        asztal[N]+lá[FAC]      Soon it became a table.
Nc-sp------ superessive    asztal[N]+on[SUP]      The cat sits on the table.
Nc-su------ distributive   asztal[N]+onként[DIS]  We collected 10 forints in each table.
-------------------------------------------------------------------------------

10.2 Verb (V)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           main           m
                 auxiliary      a
- -------------- -------------- -
2 VForm          indicative     i
                 imperative     m
                 conditional    c
                 infinitive     n
- -------------- -------------- -
3 Tense          present        p
                 past           s
- -------------- -------------- -
4 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
5 Number         singular       s
                 plural         p
- -------------- -------------- -
6 Gender                        -

*********************************
7 Voice                         -
- -------------- -------------- -
8 Negative                      -
- -------------- -------------- -
9 Definiteness   no             n
                 yes            y
           l.s.  1s2s           2
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
=================================

Notes:
1. 1s2s is a special form for definitness, in which the speaker's person is
   first singular (I) and the target of the transitivity is second singular
   (you).
2. A Hungarian verb has
   3(Person)*2(Number)*4(Present indicative, Present imperativ,
   Present conditional,Past indicative)*2(Transitivity)
   +4(1s2s)+7(Infinitive)=59 different forms.

Combinations:
--- ---- ----- ----- ------ ------ -----------
Pos Type VForm Tense Person Number Definitness
--- ---- ----- ----- ------ ------ -----------
 V   m   i     [ps] [123]   [sp]   [yn]
 V   m   c     p    [123]   [sp]   [yn]
 V   m   m     p    [123]   [sp]   [yn]
 V   m   i     [ps] 1       s      2
 V   m   [cm]  p    1       s      2
 V   m   n     -    -       -      -
 V   m   n     p    [123]   [sp]   -
 V   a   i     p    [123]   [sp]   n
--- ---- ----- ----- ------ ------ -----------

Examples on combinations:
01234567890123
Vmip1s------n-   tanul[V]+ok[e1]                 I learn (a)
Vmip2s------n-   tanul[V]+sz[e2]                 you learn (a)
Vmip3s------n-   tanul[V]                        she learns (a)

Vmip1p------n-   tanul[V]+unk[t1]                we learn (a)
Vmip2p------n-   tanul[V]+tok[t2]                you learn  (a)
Vmip3p------n-   tanul[V]+nak[t3]                they learn (a)
Vmip1s------y-   tanul[V]+om[Te1]                I learn (the)
Vmip2s------y-   tanul[V]+od[Te2]                you learn (the)
Vmip3s------y-   tanul[V]+ja[Te3]                she learns (the)
Vmip1p------y-   tanul[V]+juk[Tt1]               we learn (the)
Vmip2p------y-   tanul[V]+játok[Tt2]             you learn  (the)
Vmip3p------y-   tanul[V]+ják[Tt3]               they learn (the)
Vmis1s------n-   tanul[V]+tam[Me1]               I learnt (a)
Vmis2s------n-   tanul[V]+tál[Me2]               you learnt (a)
Vmis3s------n-   tanul[V]+t[Me3]                 she learnt (a)
Vmis1p------n-   tanul[V]+tunk[Mt1]              we learnt (a)
Vmis2p------n-   tanul[V]+tatok[Mt2]             you learnt  (a)
Vmis3p------n-   tanul[V]+tak[Mt3]               they learnt (a)
Vmis1s------y-   tanul[V]+tam[TMe1]              I learnt (the)
Vmis2s------y-   tanul[V]+tad[TMe2]              you learnt (the)
Vmis3s------y-   tanul[V]+ta[TMe3]               she learnt (the)
Vmis1p------y-   tanul[V]+tuk[TMt1]              we learnt (the)
Vmis2p------y-   tanul[V]+tátok[TMt2]            you learnt  (the)
Vmis3p------y-   tanul[V]+ták[TMt3]              they learnt (the)
Vmcp1s------n-   tanul[V]+nék[Fe1]               (if) I learnt (a)
Vmcp2s------n-   tanul[V]+nál[Fe2]               (if) you learnt (a)
Vmcp3s------n-   tanul[V]+na[Fe3]                (if) she learnt (a)
Vmcp1p------n-   tanul[V]+nánk[Ft1]              (if) we learnt (a)
Vmcp2p------n-   tanul[V]+nátok[Ft2]             (if) you learnt  (a)
Vmcp3p------n-   tanul[V]+nának[Ft3]             (if) they learnt (a)
Vmcp1s------y-   tanul[V]+nám[TFe1]              (if) I learnt (the)
Vmcp2s------y-   tanul[V]+nád[TFe2]              (if) you learnt (the)
Vmcp3s------y-   tanul[V]+ná[TFe3]               (if) she learnt (the)
Vmcp1p------y-   tanul[V]+nánk[TFt1]             (if) we learnt (the)
Vmcp2p------y-   tanul[V]+nátok[TFt2]            (if) you learnt  (the)
Vmcp3p------y-   tanul[V]+nák[TFt3]              (if) they learnt (the)
Vmmp1s------n-   tanul[V]+jak[Pe1]               (let) me learn (a)
Vmmp2s------n-   tanul[V]+jál[Pe2]               (let) you learn (a)
Vmmp3s------n-   tanul[V]+jon[Pe3]               (let) she learn (a)
Vmmp1p------n-   tanul[V]+junk[Pt1]              (let) we learn (a)
Vmmp2p------n-   tanul[V]+jatok[Pt2]             (let) you learn  (a)
Vmmp3p------n-   tanul[V]+janak[Pt3]             (let) they learn (a)
Vmmp1s------y-   tanul[V]+jam[TPe1]              (let) me learn (the)
Vmmp2s------y-   tanul[V]+jad[TPe2]              (let) you learn (the)
Vmmp3s------y-   tanul[V]+ja[TPe3]               (let) she learn (the)
Vmmp1p------y-   tanul[V]+juk[TPt1]              (let) we learn (the)
Vmmp2p------y-   tanul[V]+játok[TPt2]            (let) you learn  (the)
Vmmp3p------y-   tanul[V]+ják[TPt3]              (let) they learn (the)
Vmip1s------2-   tanul[V]+lak[Ie1]               I learn you
Vmis1s------2-   tanul[V]+talak[IMe1]            I learnt you
Vmcp1s------2-   tanul[V]+nálak[IFe1]            (if) I learnt you
Vmmp1s------2-   tanul[V]+jalak[IPe1]            (let) me learn you

Vmn-----------   tanul[V]+ni[INF]                to learn
Vmnp1s--------   tanul[V]+nom[INRe1]             for me to learn
Vmnp2s--------   tanul[V]+nod[INRe2]             for you to learn
Vmnp3s--------   tanul[V]+nia[INRe3]             for her to learn
Vmnp1p--------   tanul[V]+nunk[INRt1]            for us to learn
Vmnp2p--------   tanul[V]+notok[INRt2]           for you to learn
Vmnp3p--------   tanul[V]+niuk[INRt3]            for them to learn
Vaip1s--------   fog[V]+ok[e1]                   I shall

10.3 Adjective (A)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           qualificative  f
- -------------- -------------- -
2 Degree         positive       p
                 comparative    c
                 superlative    s
- -------------- -------------- -
3 Gender                        -
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 instrumental   i
                 illative       x
                 inessive       2
                 elative        e
                 allative       t
                 adessive       3
                 ablative       b
                 terminative    9
                 essive         w
           l.s.  temporalis     m
           l.s.  causalis       c
           l.s.  sublative      s
           l.s.  delative       h
           l.s.  sociative      q
           l.s.  factive        y
           l.s.  superessive    p
           l.s.  distributive   u
           l.s.  essive_formal  f
* ***************************** *
6 Definiteness                  -

- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Animate        no             n
                 yes            y
- -------------- -------------- -
9 Formation                     -
- -------------- -------------- -
10Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
11Owner_Person   first          1
                 second         2
                 third          3
---------------- -------------- -
12Owned_Number   singular       s
                 plural         p
=================================

Notes:
1. Adjectives have the same suffixes that Nouns have plus Degree.
   Thus the number of possibble forms is 3*924=2772.
2. Types like ordinal and cardinal are grouped as numerals.

Combinations:
--- ---- ----- ---- ------------------------ ----  -------- --------- --------
Pos Type Degr  Num  Case                     Anim  OwnerNum OwnerPers OwnedNum
--- ---- ----- ---- ------------------------ ----  -------- --------- --------
 A  f    [pcs] [sp] [ngdaix2et3b9wmcshqypuf] [yn]  [sp]     [123]     [sp]
--- ---- ----- ---- ------------------------ ----  -------- --------- --------

Examples on combinations:
0123456789012
Afp-sn-------    nagy[A]                           big, great
Afc-sn-------    nagy[A]+obb[FOK]                  bigger
Afs-sn-------    leg[FF]+nagy[A]+obb[FOK]          biggest
Afs-sn----p1s    leg[FF]+nagy[A]+obb[FOK]+jaink[PSt1i]+é[POS]
                                                   that of our greatest (heroes)
Afp-sw-------    magyar[A]+ul[ESSMOD]              in Hungarian

10.4 Pronoun (P)

= ============== ============== =

P ATT            VAL            C
= ============== ============== =
1 Type           personal       p
demonstrative  d
indefinite     i
possessive     s
interrogative  q
relative       r
reflexive      x
reciprocal     y
- -------------- -------------- -
2 Person         first          1
second         2
third          3
- -------------- -------------- -
3 Gender                        -
- -------------- -------------- -
4 Number         singular       s
plural         p
- -------------- -------------- -
5 Case           nominative     n
genitive       g
dative         d
accusative     a
instrumental   i
illative       x
inessive       2
elative        e
allative       t
adessive       3
ablative       b
terminative    9
essive         w
l.s.  temporalis     m
l.s.  causalis       c
l.s.  sublative      s
l.s.  delative       h
l.s.  sociative      q
l.s.  factive        y
l.s.  superessive    p
l.s.  distributive   u
l.s.  essive_formal  f
- -------------- -------------- -
6 Owner_Number   singular       s
plural         p
- -------------- -------------- -
7 Owner_Gender                  -
*********************************
8 Clitic                        -
- -------------- -------------- -
9 Referent_Type                 -
- -------------- -------------- -
10Syntactic_Type                -

- -------------- -------------- -
11Definiteness                  -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
- -------------- -------------- -
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person   first          1
                 second         2
                 third          3
- -------------- -------------- -
16Owned_Number   singular       s
                 plural         p
=================================

Combinations:
--- ---- ----- ---- ------------------------  -----------  ------------ ------------
Pos Type Pers  Num  Case                      Owner_Number Owner_Person Owned_Number
--- ---- ----- ---- ------------------------  ------------ ------------ ------------
 P  [px] [123] [sp] [ngdaix2et3b9wmcshqypuf]  [sp]         [123]        [sp]
 P    y     3   p   [ngdaix2et3b9wmcshqypuf]  [sp]         [123]        [sp]
 P    s  [123] [sp] [ngdaix2et3b9wmcshqypuf]  -             -
 P [diqr]  3   [sp] [ngdaix2et3b9wmcshqypuf]  [sp]         [123]        [sp]

Examples on combinations:
01234567890123456
Pp1-s------------         én[P]                    I
Pp1-p-s--------1-         enyéim[P]                my things (plural)
Pd3-s------------         ez[P]                    this
Pi3-s------------         minden[P]                every
Ps1-s-s--------1-         én[P]=enyém+[POS]        my
Pq3-s------------         ki[P]                    who (in questions)
Pr3-s------------         aki[P]                   who (relative)
Px3-s------------         maga[P]                  himself
Pp3-p------------         ti[P]                    you
Pp3-s-p--------2-         tietek[P]                your something
Pp3-p-p--------2-         tieitek[P]               your somethings
Pp3-s-p--------2s         tieteké[P]               your something
Py3-p------------         egymás[P]                each other

10.5 Determiner (D)

Not applicable.

10.6 Article (T)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           definite       f
                 indefinite     i
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
*********************************
5 Clitic                        -
=================================

Notes:
We have tree articles: a, az and egy.  These may not have number and
case. The word 'az' may have but that is a pronoun in those cases.

Combinations:
--- ----
Pos Type
--- ----
T   [fi]
01
012345
Tf----              a[DET]              the
Tf----              az[DET]             the
Ti----              egy[DET]            a

10.7 Adverb (R)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           general        g
                 particle       p
           l.s.  causal         o
           l.s.  verbal         v
                 modifier       m
           l.s.  interrogative  q
- -------------- -------------- -
2 Degree                        -

*********************************
3 Clitic         no             n
                 yes            y
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Person         first          1
                 second         2
                 third          3
=================================

Combinations:
----  ------ ------  ------  ------ ------
Pos   Type   Degree  Clitic  Number Person
----  ------ ------  ------  ------ ------
 R    [gmp]   -        -        -     -
 R       q    -        y        -     -
 R       o    -        -        [sp]  [123]
----  ------ ------  ------  ------ ------

Examples on combinations:
012345
Rg----         most[R]                               now
Rg----         hogyan[R]                             how
Rgp---         belül[R]                              inside
Rgc---         belül[R]=belj+ebb[FOK]                deeper inside
Rgs---         leg[FF]+belül[R]=belj+ebb[FOK]        deepest inside
Rv----         olvas[V]+va[HIN]                      being read
Rm----         nem[R]                                no
Rp----         be[IK]                                into
Rq-y--         kutya[N]+-e[KSZE] (kutya-e?)          dog?
Ro--s1         áll[V]+tomban[HIVe1]                  as I stand
Ro--s2         áll[V]+todban[HIVe2]                  as you stand
Ro--s3         áll[V]+tában[HIVe3]                   as she stands
Ro--p1         áll[V]+tunkban[HIVp1]                 as we stand
Ro--p2         áll[V]+totokban[HIVp2]                as you stand
Ro--sp         áll[V]+tukban[HIVep]                  as they stand
Ro----         áll[V]+ván[HINN]                      as sy stands

Notes:
1. The modifier -e question word (the only Hungarian clitic) is attached
   to the preceding word with a hyphen.
2. Some adverbs may have degree, but these do not formulate a special
   class. Presently we cannot give a criteria for this.

10.8 Adposition (S)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           postposition   t
- -------------- -------------- -
2 Formation                     -
*********************************
3 Case                          -
- -------------- -------------- -
4 Clitic                        -
=================================

Notes:
1. Some of them require a suffix on the preceding noun (e.g. asztal[N]+
   on[SUP] túl[S]) while others do not (e.g. asztal[N] mellett[S]).
   This feature is not represented here.

Combinations:
---  ----
Pos  Type
---  ----
 S     t
---  ----
Examples on combinations:
01234
St---      miatt[S]                      because of
St---      keresztül[S]                  through

10.9 Conjunction (C)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           coordinating   c
                 subordinating  s
*********************************
2 Formation      simple         s
                 compound       c
- -------------- -------------- -
3 Coord_Type     sentence       p
                 words          w
- -------------- -------------- -

4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- -
6 Number                        -
- -------------- -------------- -
7 Person                        -
=================================

Notes:
1. 'form': Some conjuctions in Hungarian might occur alone (1) while
    others in pairs (2).

Combinations:
---  ---- ---------- ---------
Pos  Type Coord-Type Formation
---  ---- ---------- ---------
 C    [cs]   [pw]       [12]
---  ---- ---------- ---------

Examples on combinations:              Meaning:
01234567
Csp-1---          hogy[C]                that
Ccp-1---          és[C]                  and
Ccw-2---          is[C]                  also
Ccw-2---          mind[C]                both

10.10 Numeral (M)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           cardinal       c
                 ordinal        o
                 fractal        f
                 collect        l
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case           nominative     n
                 genitive       g

                 dative         d
                 accusative     a
                 instrumental   i
                 illative       x
                 inessive       2
                 elative        e
                 allative       t
                 adessive       3
                 ablative       b
                 terminative    9
                 essive         w
           l.s.  temporalis     m
           l.s.  causalis       c
           l.s.  sublative      s
           l.s.  delative       h
           l.s.  sociative      q
           l.s.  factive        y
           l.s.  superessive    p
           l.s.  distributive   u
           l.s.  essive_formal  f
           l.s.  multiplicative 6
*********************************
5 Form           digit          d
                 roman          r
                 letter         l
- -------------- -------------- -
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Class                         -
- -------------- -------------- -
9 Animate                       -
- -------------- -------------- -
10Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
11Owner_Person   first          1
                 second         2
                 third          3
- -------------- -------------- -
12Owned_Number   singular       s
                 plural         p
=================================

Combinations:
--- ---- ---- ------------------------ ----  -------- --------- --------
Pos Type Num  Case                     Form  OwnerNum OwnerPers OwnedNum
--- ---- ---- ------------------------ ----  -------- --------- --------
 N  [cp] [sp] [ngdaix2et3b9wmcshqypuf] [dlr] [sp]    [123]     [sp]
--- ---- ---- ------------------------ ----  -------- --------- --------

Examples on combinations:
0123456789012
Mc-sn--------        egy[M]                  one
Mc-pn--------        nyolc[M]+ak[PL]         those eight (men)
Mc-pn-----p3s        nyolc[M]+ak[PL]+é[POS]  this belongs to those eight (men)
Mc-sa--------        sok[M]+at[ACC]          many (in accusative)
Mo-sn--------        hat[M]+odik[SORSZ]      6th
Mf-sn--------        hat[M]+od[TORT]         1/6
Mc-s6--------        egy[M]+szer[MUL]        one time
Mc-sw--------        hat[M]+an[ESSMOD]       six of them

10.11 Interjection (I)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           mood           m
                 other          o
- -------------- -------------- -
2 Formation                     -
=================================

10.12 Residual (X)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
                                -
=================================

10.13 Abbreviation (Y)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Syntactic_Type                -
- -------------------------------
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -

- -------------- -------------- -
5 Definiteness                  -
=================================

10.14 Particle (Q)

Not applicable.

10.15 Derivation

Without handling derivation a satisfactory morphological analysis is not possible for Hungarian. The HUMOR system, a general purpose morphological analyzer, handles it and the results can be converted to Multext format since on syntactic level the morphological origin of stems are generally irrelevant. The resulting word class is defined by the rightmost derivational suffix. The suffix characters are literally attached to the word.

Here are the derivations that the analyser recognizes but instead of the origin we place only the resulting class to the output. (Suffix tags used in HUMOR are in upper case, actual suffixes are in lower case.)

Noun   -> Adjective
-------------------
-----   -------        --------------    ------------------------
Type    Form           Example           Comment
-----   -------        --------------    ------------------------
 BELI  : beli           ház+beli         property of living in the house
 FAJTA : fajta          más+fajta        of some other kind
 FELE  : féle           bútor+féle       similar to furniture
 FORMA : forma          tojás+forma      egg shaped
 SZERU : szerű          tej+szerű        milk+y
 IKEP  : i              ház+i            home (e.g. made)
 SKEP  : s, as, os,
         es, ös         gyerek+es        child+ish
 UKEP  : ú,
         ű,
         jú,
         jű             arc+ú            (red)-face+d
 FFOSZ : tlan,tlen,
         atlan,etlen,
         talan,telen                     'devoid of', '-less'
 MER   : nyi            kanál+nyi        spoon+full

Noun,A -> Noun
--------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------

 COL   : ság,          barát+ság          friend+ship
         ség

Noun   -> Noun
--------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 DIM   : cska,acska,
         ecske,öcske,
         ocska         utcá+cska          little street
 FEM   : né            Kovács+né          Mrs. Kovács

Noun   -> Verb
--------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 FI    : z,az,oz,ez,
         öz            autó+z             go by car

Verb   -> Adjective
-------------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 IFOSZT: atlan, etlen  felel+etlen        sg not being answered
 MIF   : ó,
         ő             felel+ő            sy who answers
 MIB   : t,ott,ett,
         ött           felel+t            the answered (question)
 MIA   : andó,
         endő          felel+endő         sg that should be answered
 NIVALÖ  anivaló,
         enivaló,
         nivaló        néz+nivaló         sg that should be seen

Verb   -> Adverb
----------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 HIN   : va,ve         olvas+va           (while ) reading (the book)

Numeral-> Adjective
-------------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment

-----   -------        --------------     ------------------------
 KIEM  : ik            hatod+ik           six+th
 LAGOS : lagos,leges   másod+lagos        second+ary

Verb   -> Noun
--------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 IF    : ás,
         és            olvas+ás           read+ing (gerund)
 DES   : hatnék,
         hetnék        olvas+hatnék       the intention of reading

Adj    -> Verb
--------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 FAK   : ít            szép+ít            make it pretty (in compounds only)
 MI    : od,ed,öd      vállas+od+ik       becomes strong
 MIGY  : kod,ked,
         köd           okos+kod+ik        plays the smart (frequently)

Verb   -> Verb
--------------
-----   -------        --------------     ------------------------
Type    Form           Example            Comment
-----   -------        --------------     ------------------------
 MUV   : at,et,tat,tet olvas+tat          makes him read
 GYAK  : gat,get,ogat,
         eget,öget     olvas+gat          he reads frequently
 HAT   : hat,het       olvas+hat          he may read
 VISSZ : ód,
         őd            old+ódik           dissolves
 SZENV : tatik, tetik  olvas+tatik        makes the book being read

Examples:

If ’szemtelés’ (littering, action of throwing away litter) is not in the dictionary we derive it from the verb ’litter’ szemetel[V] + és[IF] (where IF=Verb2Noun).

Instead of giving the extra attribute to the verb expressing that it has a derivational suffix we simply give the result of the analysis+conversion: szemetelés[N].

In Hungarian some derivation may follow the inflectional suffix. For these derivations the suffix+derivation together forms a compound derivation. The a new stem is generated from the stem + inflection + derivation segments and the resulting part of speech is determined by the derivation.

------------------     -----   --------------     ------------------------
Type                   Form    Example            Comment
------------------     -----   --------------     ------------------------
Nc-sn--ns1-   +FAM  :   ék     apá+m+ék           some people with my father
Nc-su--u---   +IKEP :   i      asztal+onként+i    sg. done by each every table
Afc-sn--n---- +KIEM :   ik     nagy+obb+ik        the bigger one

There are adverbs that may get case endings. Since case inflections derivate adverb from nouns these constructions can be handled as derivations. That means the stem is the stem + inflection combination and the part of speech is adverb.

------------------     -----   --------------     ------------------------
Type                    Form   Example            Comment
------------------     -----   --------------     ------------------------
Ag----     + ablativ    tól    akkor+tól          since then

10.16 Compounding

Compounding is handled in a very similar way to derivation. The rightmost word class is always the resulting one. If it contains some derivation as well then the result is the word class that the derivation determines.

Examples:

1. Hungarian is similar to German but we have a competence limit: we do not put together more then 2 nouns. These two words might be compounds as well but they must be lexicalized forms. So e.g.: rendőregyenruha (police uniform) is valid since it is put together from rendőr (police) and egyenruha (uniform) where rendőr=rend (order) + őr (guard) and egyenruha=egyen (uni) + ruha (clothes).

2. Of course the situation is more complicated since nouns can be derived from verbs, adjectives or other nouns, and these derivations can be parts of compoundns: e.g. autóbuszmegálló (bus stop) where meg+áll+ó is derived from the verb áll (stand). Whereas megálló is lexicalized so dictionaries might contain it, the morphological rule is productive so compounding works well for constructions like tandíjbefizetés (paying the tution fee)= tan(tution)+ díj(fee)+ be(in) +fizet(pay) +és(ment).

3. Compounding, such as derivation, will not be mirrorred in the Multext format morphological analysis* a converter will produce an acceptable segmentation in the corpus. The word class (POS) for these compounds will be the word class of the output compound (e.g. tandíjbefizetés/N ’the action of paying in the tuition fee’).

Application to Croatian

Authors: Marko Tadić, Zavod za lingvistiku Filozofskoga fakulteta sveučilišta u Zagrebu.

This specification is implemented in the lemmatisation module used for the Croatian National Corpus.

Zagreb, 10. July 1998.

11.1 Noun (N)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           common         c
                 proper         p
- -------------- -------------- -
2 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 vocative       v
                 locative       l
                 instrumental   i
* ***************************** *
5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -
- -------------- -------------- -
7 Animate        no             n

TELRI  Morphosyntactic specifications V.2— Croatian  203

                 yes            y
- -------------- -------------- -
8 Owner_Number                  -
- -------------- -------------- -
9 Owner_Person                  -
- -------------- -------------- -
10Owned_Number                  -
=================================

11.2 Verb (V)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           main           m
                 auxiliary      a
                 modal          o
                 copula         c
- -------------- -------------- -
2 VForm          indicative     i
                 imperative     m
                 conditional    c
                 infinitive     n
                 participle     p
- -------------- -------------- -
3 Tense          present        p
                 imperfect      i
                 future         f
                 past           s
           l.s.  pluperfect     l
           l.s.  aorist         a
- -------------- -------------- -
4 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
5 Number         singular       s
                 plural         p
- -------------- -------------- -
6 Gender         masculine      m
                 feminine       f
                 neuter         n
*********************************
7 Voice          active         a
                 passive        p
- -------------- -------------- -
8 Negative       no             n
                 yes            y
- -------------- -------------- -
9 Definiteness                  -

- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
=================================

11.3 Adjective (A)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           qualificative  f
                 possessive     s
- -------------- -------------- -
2 Degree         positive       p
                 comparative    c
                 superlative    s
- -------------- -------------- -
3 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 vocative       v
                 locative       l
                 instrumental   i
* ***************************** *
6 Definiteness   no             n
                 yes            y
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Animate        no             n
                 yes            y
- -------------- -------------- -
9 Formation                     -
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -

---------------- -------------- -
12Owned_Number                  -
=================================

11.4 Pronoun (P)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           personal       p
                 demonstrative  d
                 indefinite     i
                 possessive     s
                 interrogative  q
                 relative       r
                 reflexive      x
- -------------- -------------- -
2 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
3 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 vocative       v
                 locative       l
                 instrumental   i
- -------------- -------------- -
6 Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
7 Owner_Gender   masculine      m
                 feminine       f
                 neuter         n
*********************************
8 Clitic         no             n
                 yes            y
- -------------- -------------- -
9 Referent_Type  personal       p
                 possessive     s
- -------------- -------------- -
10Syntactic_Type nominal        n

                 adjectival     a
- -------------- -------------- -
11Definiteness                  -
- -------------- -------------- -
12Animate        no             n
                 yes            y
- -------------- -------------- -
13Clitic_s                      -
- -------------- -------------- -
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person                  -
- -------------- -------------- -
16Owned_Number                  -
- -------------- -------------- -
17Wh_Type                       -
=================================

11.5 Determiner (D)

Not applicable.

11.6 Article (T)

Not applicable.

11.7 Adverb (R)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           general        g
- -------------- -------------- -
2 Degree         positive       p
                 comparative    c
                 superlative    s
*********************************
3 Clitic                        -
- -------------- -------------- -
4 Number                        -
- -------------- -------------- -
5 Person                        -
- -------------- -------------- -
6 Wh_Type                       -
=================================

11.8 Adposition (S)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           preposition    p
- -------------- -------------- -
2 Formation      simple         s
                 compound       c
*********************************
3 Case
  (req.by prep.) genitive       g
                 dative         d
                 accusative     a
                 locative       l
                 instrumental   i
- -------------- -------------- -
4 Clitic                        -
=================================

11.9 Conjunction (C)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           coordinating   c
                 subordinating  s
*********************************
2 Formation      simple         s
                 compound       c
- -------------- -------------- -
3 Coord_Type                    -
- -------------- -------------- -
4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- -
6 Number                        -
- -------------- -------------- -
7 Person                        -
=================================

11.10 Numeral (M)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           cardinal       c

                 ordinal        o
                 multiple       m
                 special        s
- -------------- -------------- -
2 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
           l.s.  vocative       v
                 locative       l
                 instrumental   i
*********************************
5 Form           digit          d
                 roman          r
                 letter         l
- -------------- -------------- -
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Class                         -
- -------------- -------------- -
9 Animate        no             n
                 yes            y
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================

11.11 Interjection (I)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type                          -
- -------------- -------------- -
2 Formation      simple         s
                 compound       c
=================================

11.12 Residual (X)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
                                -
=================================

11.13 Abbreviation (Y)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Syntactic_Type nominal        n
                 adverbial      r
- -------------------------------
2 Gender         masculine      m
                 feminine       f
                 neuter         n
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case           nominative     n
                 genitive       g
           l.s.  dative         d
           l.s.  accusative     a
           l.s.  locative       l
           l.s.  instrumental   i
- -------------- -------------- -
5 Definiteness                  -
=================================

11.14 Particle (Q)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           negative       z
                 interrogative  q
                 modal          o
                 affirmative    r
- -------------- -------------- -
2 Formation                     -
- -------------- -------------- -
3 Clitic                        -
=================================

TELRI  Morphosyntactic specifications V.2— Croatian  210

Notes:
Type=negative used for negation with 'ne', =interrogative for 'li' and
'zar, =affirative for 'da'.